%% %% This is file ‘.tex’, %% generated with the docstrip utility. %% %% The original source files were: %% %% fileerr.dtx (with options: ‘return’) %% %% This is a generated file. %% %% Copyright 1993 1994 1995 1996 1997 1998 1999 %% The LaTeX3 Project and any individual authors listed elsewhere %% in this file. %% %% This file is part of the Standard LaTeX ‘Tools Bundle’. %% ——————————————————- %% %% This file, and all the other files in this bundle as listed below, %% may be distributed under the terms of the LaTeX Project Public %% License, as described in lppl.txt in the base LaTeX distribution. %% Either version 1.0 or, at your option, any later version. %% \messageFile ignored \endinput %% %% End of file ‘.tex’.

Debugging with [No value for “GDBN”]

The gnu Source-Level Debugger

Ninth Edition, for [No value for “GDBN”] version [No value for “GDBVN”] December 2001

Richard Stallman, Roland Pesch, Stan Shebs, et al.

(Send bugs and comments on [No value for “GDBN”] to bug-gdb@gnu.org.)

Debugging with [No value for “GDBN”] TEXinfo 1999-09-25.10

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1988,1989,1990,1991,1992,1993,1994,1995,1996,1998,1999,2000,2001 Free Software Foundation, Inc.

Published by the Free Software Foundation 59 Temple Place - Suite 330, Boston, MA 02111-1307 USA ISBN 1-882114-77-9

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with the Invariant Sections being “Free Software” and “Free Software Needs Free Documentation”, with the Front-Cover Texts being “A GNU Manual,” and with the Back-Cover Texts as in (a) below.

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.”

Table of Contents

Summary of [No value for “GDBN”] . . . . . . . . . . . 1

Free software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Free Software Needs Free Documentation . . . . . . . . . . . . . . . . . . . . . . 1 Contributors to [No value for “GDBN”] . . . . . . . . . . . . . . . . . . . . . . . . 3

1 A Sample [No value for “GDBN”] Session … 7

2 Getting In and Out of [No value for “GDBN”] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2.1 Invoking [No value for “GDBN”] . . . . . . . . . . . . . . . . . . . . . . . . 11 2.1.1 Choosing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.1.2 Choosing modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.2 Quitting [No value for “GDBN”] . . . . . . . . . . . . . . . . . . . . . . . . 15 2.3 Shell commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3 [No value for “GDBN”] Commands . . . . . . . . 17

3.1 Command syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2 Command completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.3 Getting help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

4 Running Programs Under [No value for “GDBN”] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

4.1 Compiling for debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.2 Starting your program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.3 Your program’s arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.4 Your program’s environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.5 Your program’s working directory . . . . . . . . . . . . . . . . . . . . . . . 26 4.6 Your program’s input and output. . . . . . . . . . . . . . . . . . . . . . . . 26 4.7 Debugging an already-running process . . . . . . . . . . . . . . . . . . . 27 4.8 Killing the child process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 4.9 Debugging programs with multiple threads . . . . . . . . . . . . . . . 28 4.10 Debugging programs with multiple processes . . . . . . . . . . . . 31

5 Stopping and Continuing . . . . . . . . . . . . . . . . . . 33

5.1 Breakpoints, watchpoints, and catchpoints . . . . . . . . . . . . . . . 33 5.1.1 Setting breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 5.1.2 Setting watchpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 5.1.3 Setting catchpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 5.1.4 Deleting breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.1.5 Disabling breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.1.6 Break conditions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.1.7 Breakpoint command lists . . . . . . . . . . . . . . . . . . . . . . 43 5.1.8 Breakpoint menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.1.9 “Cannot insert breakpoints” . . . . . . . . . . . . . . . . . . . . 45

5.2 Continuing and stepping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.3 Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 5.4 Stopping and starting multi-thread programs . . . . . . . . . . . . . 50

6 Examining the Stack . . . . . . . . . . . . . . . . . . . . . . 53

6.1 Stack frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 6.2 Backtraces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.3 Selecting a frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 6.4 Information about a frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

7 Examining Source Files . . . . . . . . . . . . . . . . . . . 59

7.1 Printing source lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.2 Searching source files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.3 Specifying source directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.4 Source and machine code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

8 Examining Data . . . . . . . . . . . . . . . . . . . . . . . . . . 65

8.1 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 8.2 Program variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 8.3 Artificial arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 8.4 Output formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 8.5 Examining memory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 8.6 Automatic display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 8.7 Print settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.8 Value history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 8.9 Convenience variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 8.10 Registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 8.11 Floating point hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.12 Memory Region Attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

8.12.1 Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 8.12.1.1 Memory Access Mode . . . . . . . . . . . . . . . . 81 8.12.1.2 Memory Access Size. . . . . . . . . . . . . . . . . . 81 8.12.1.3 Data Cache. . . . . . . . . . . . . . . . . . . . . . . . . . 82

9 Tracepoints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

9.1 Commands to Set Tracepoints. . . . . . . . . . . . . . . . . . . . . . . . . . . 83 9.1.1 Create and Delete Tracepoints . . . . . . . . . . . . . . . . . . 83 9.1.2 Enable and Disable Tracepoints. . . . . . . . . . . . . . . . . 84 9.1.3 Tracepoint Passcounts . . . . . . . . . . . . . . . . . . . . . . . . . 84 9.1.4 Tracepoint Action Lists . . . . . . . . . . . . . . . . . . . . . . . . 85 9.1.5 Listing Tracepoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 9.1.6 Starting and Stopping Trace Experiment . . . . . . . . 87

9.2 Using the collected data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.2.1 tfind n. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.2.2 tdump. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 9.2.3 save-tracepoints filename . . . . . . . . . . . . . . . . . . . . 90

9.3 Convenience Variables for Tracepoints . . . . . . . . . . . . . . . . . . . 90

10 Using [No value for “GDBN”] with Different Languages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

10.1 Switching between source languages . . . . . . . . . . . . . . . . . . . . 93 10.1.1 List of filename extensions and languages . . . . . . . 93 10.1.2 Setting the working language . . . . . . . . . . . . . . . . . . 94 10.1.3 Having [No value for “GDBN”] infer the source

language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

10.2 Displaying the language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 10.3 Type and range checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

10.3.1 An overview of type checking . . . . . . . . . . . . . . . . . . 95 10.3.2 An overview of range checking . . . . . . . . . . . . . . . . . 96

10.4 Supported languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

10.4.1 C and C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 10.4.1.1 C and C++ operators . . . . . . . . . . . . . . . . . 98 10.4.1.2 C and C++ constants . . . . . . . . . . . . . . . . 100 10.4.1.3 C++ expressions. . . . . . . . . . . . . . . . . . . . . 100 10.4.1.4 C and C++ defaults . . . . . . . . . . . . . . . . . 101 10.4.1.5 C and C++ type and range checks …. 102 10.4.1.6 [No value for “GDBN”] and C . . . . . . . 102 10.4.1.7 [No value for “GDBN”] features for C++

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

10.4.2 Modula-2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 10.4.2.1 Operators . . . . . . . . . . . . . . . . . . . . . . . . . . 104 10.4.2.2 Built-in functions and procedures …. 105 10.4.2.3 Constants . . . . . . . . . . . . . . . . . . . . . . . . . . 106 10.4.2.4 Modula-2 defaults. . . . . . . . . . . . . . . . . . . 107 10.4.2.5 Deviations from standard Modula-2 . . 107 10.4.2.6 Modula-2 type and range checks . . . . . 107 10.4.2.7 The scope operators :: and . . . . . . . . . 107 10.4.2.8 [No value for “GDBN”] and Modula-2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

10.4.3 Chill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 10.4.3.1 How modes are displayed . . . . . . . . . . . . 108 10.4.3.2 Locations and their accesses . . . . . . . . . 110 10.4.3.3 Values and their Operations . . . . . . . . . 111 10.4.3.4 Chill type and range checks . . . . . . . . . . 113 10.4.3.5 Chill defaults . . . . . . . . . . . . . . . . . . . . . . . 113

11 Examining the Symbol Table . . . . . . . . . . . . 115

12 Altering Execution . . . . . . . . . . . . . . . . . . . . . 119

12.1 Assignment to variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 12.2 Continuing at a different address . . . . . . . . . . . . . . . . . . . . . . 120 12.3 Giving your program a signal . . . . . . . . . . . . . . . . . . . . . . . . . 121 12.4 Returning from a function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 12.5 Calling program functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 12.6 Patching programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

13 [No value for “GDBN”] Files . . . . . . . . . . . . 123

13.1 Commands to specify files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 13.2 Errors reading symbol files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

14 Specifying a Debugging Target . . . . . . . . . . 129

14.1 Active targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 14.2 Commands for managing targets . . . . . . . . . . . . . . . . . . . . . . 129 14.3 Choosing target byte order. . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 14.4 Remote debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

14.4.1 The [No value for “GDBN”] remote serial protocol

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 14.4.1.1 What the stub can do for you . . . . . . . . 133 14.4.1.2 What you must do for the stub . . . . . . 134 14.4.1.3 Putting it all together . . . . . . . . . . . . . . . 135 14.4.1.4 Communication protocol . . . . . . . . . . . . 136 14.4.1.5 Using the gdbserver program . . . . . . . 148 14.4.1.6 Using the gdbserve.nlm program …. 149

14.5 Kernel Object Display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150

15 Configuration-Specific Information. . . . . . . 151

15.1 Native . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 15.1.1 HP-UX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 15.1.2 SVR4 process information . . . . . . . . . . . . . . . . . . . . 151 15.1.3 Features for Debugging djgpp Programs . . . . . . 152

15.2 Embedded Operating Systems. . . . . . . . . . . . . . . . . . . . . . . . . 153

15.2.1 Using [No value for “GDBN”] with VxWorks . . 154 15.2.1.1 Connecting to VxWorks . . . . . . . . . . . . . 154 15.2.1.2 VxWorks download . . . . . . . . . . . . . . . . . 155 15.2.1.3 Running tasks . . . . . . . . . . . . . . . . . . . . . . 155

15.3 Embedded Processors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

15.3.1 AMD A29K Embedded . . . . . . . . . . . . . . . . . . . . . . 156 15.3.1.1 A29K UDI . . . . . . . . . . . . . . . . . . . . . . . . . 156 15.3.1.2 EBMON protocol for AMD29K . . . . . . 156 15.3.1.3 Communications setup . . . . . . . . . . . . . . 156 15.3.1.4 EB29K cross-debugging . . . . . . . . . . . . . 158 15.3.1.5 Remote log . . . . . . . . . . . . . . . . . . . . . . . . . 158

15.3.2 ARM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 15.3.3 Hitachi H8/300. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

15.3.3.1 Connecting to Hitachi boards . . . . . . . . 159 15.3.3.2 Using the E7000 in-circuit emulator . . 161 15.3.3.3 Special [No value for “GDBN”] commands for Hitachi micros. . . . . . . . . . . . . . . . . . . . . . . . 161

15.3.4 H8/500 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 15.3.5 Intel i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

15.3.5.1 Startup with Nindy . . . . . . . . . . . . . . . . . 162 15.3.5.2 Options for Nindy . . . . . . . . . . . . . . . . . . 162 15.3.5.3 Nindy reset command . . . . . . . . . . . . . . . 163

15.3.6 Mitsubishi M32R/D . . . . . . . . . . . . . . . . . . . . . . . . . 163 15.3.7 M68k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 15.3.8 M88K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 15.3.9 MIPS Embedded . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 15.3.10 PowerPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 15.3.11 HP PA Embedded . . . . . . . . . . . . . . . . . . . . . . . . . . 166 15.3.12 Hitachi SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 15.3.13 Tsqware Sparclet . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

15.3.13.1 Setting file to debug. . . . . . . . . . . . . . . . 167 15.3.13.2 Connecting to Sparclet . . . . . . . . . . . . . 167 15.3.13.3 Sparclet download . . . . . . . . . . . . . . . . . 167 15.3.13.4 Running and debugging . . . . . . . . . . . . 167

15.3.14 Fujitsu Sparclite. . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 15.3.15 Tandem ST2000 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 15.3.16 Zilog Z8000 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

15.4 Architectures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 15.4.1 A29K. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 15.4.2 Alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 15.4.3 MIPS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170

16 Controlling [No value for “GDBN”]. . . . . . 171

16.1 Prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 16.2 Command editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 16.3 Command history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 16.4 Screen size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 16.5 Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 16.6 Optional warnings and messages . . . . . . . . . . . . . . . . . . . . . . 174 16.7 Optional messages about internal happenings. . . . . . . . . . . 175

17 Canned Sequences of Commands . . . . . . . . 177

17.1 User-defined commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 17.2 User-defined command hooks . . . . . . . . . . . . . . . . . . . . . . . . . 178 17.3 Command files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 17.4 Commands for controlled output . . . . . . . . . . . . . . . . . . . . . . 180

18 [No value for “GDBN”] Text User Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183

18.1 TUI overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 18.2 TUI Key Bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 18.3 TUI specific commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 18.4 TUI configuration variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

19 Using [No value for “GDBN”] under gnu Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

20 [No value for “GDBN”] Annotations. . . . . 191

20.1 What is an Annotation? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 20.2 The Server Prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 20.3 Values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 20.4 Frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 20.5 Displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 20.6 Annotation for [No value for “GDBN”] Input. . . . . . . . . . . 195 20.7 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 20.8 Information on Breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 20.9 Invalidation Notices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 20.10 Running the Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 20.11 Displaying Source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 20.12 Annotations We Might Want in the Future. . . . . . . . . . . . 198

21 Reporting Bugs in [No value for “GDBN”] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

21.1 Have you found a bug? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 21.2 How to report bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

Appendix A Formatting Documentation …. 205

Appendix B Installing [No value for “GDBN”] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

B.1 Compiling [No value for “GDBN”] in another directory . . 208 B.2 Specifying names for hosts and targets . . . . . . . . . . . . . . . . . 209 B.3 configure options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

Summary of [No value for “GDBN”]

The purpose of a debugger such as [No value for “GDBN”] is to allow you to see what is going on “inside” another program while it executes—or what another program was doing at the moment it crashed.

[No value for “GDBN”] can do four main kinds of things (plus other things in support of these) to help you catch bugs in the act:

• Start your program, specifying anything that might affect its behavior.

• Make your program stop on specified conditions.

• Examine what has happened, when your program has stopped.

• Change things in your program, so you can experiment with correcting the effects of one bug and go on to learn about another.

You can use [No value for “GDBN”] to debug programs written in C and C++. For more information, see Section 10.4 [Supported languages], page 97. For more information, see Section 10.4.1 [C and C++], page 98.

Support for Modula-2 and Chill is partial. For information on Modula-2, see Section 10.4.2 [Modula-2], page 103. For information on Chill, see Section 10.4.3 [Chill], page 108.

Debugging Pascal programs which use sets, subranges, file variables, or nested functions does not currently work. [No value for “GDBN”] does not support entering expressions, printing values, or similar features using Pascal syntax.

[No value for “GDBN”] can be used to debug programs written in Fortran, although it may be necessary to refer to some variables with a trailing underscore.

Free software

[No value for “GDBN”] is free software, protected by the gnu General Public License (GPL). The GPL gives you the freedom to copy or adapt a licensed program—but every person getting a copy also gets with it the freedom to modify that copy (which means that they must get access to the source code), and the freedom to distribute further copies. Typical software companies use copyrights to limit your freedoms; the Free Software Foundation uses the GPL to preserve these freedoms.

Fundamentally, the General Public License is a license which says that you have these freedoms and that you cannot take these freedoms away from anyone else.

Free Software Needs Free Documentation

The biggest deficiency in the free software community today is not in the software—it is the lack of good free documentation that we can include with the free software. Many of our most important programs do not come with free reference manuals and free introductory texts. Documentation is an essential part of any software package; when an important free software package does not come with a free manual and a free tutorial, that is a major gap. We have many such gaps today.

Consider Perl, for instance. The tutorial manuals that people normally use are non-free. How did this come about? Because the authors of those manuals published them with restrictive terms—no copying, no modification, source files not available—which exclude them from the free software world.

That wasn’t the first time this sort of thing happened, and it was far from the last. Many times we have heard a GNU user eagerly describe a manual that he is writing, his intended contribution to the community, only to learn that he had ruined everything by signing a publication contract to make it non-free.

Free documentation, like free software, is a matter of freedom, not price. The problem with the non-free manual is not that publishers charge a price for printed copies—that in itself is fine. (The Free Software Foundation sells printed copies of manuals, too.) The problem is the restrictions on the use of the manual. Free manuals are available in source code form, and give you permission to copy and modify. Non-free manuals do not allow this.

The criteria of freedom for a free manual are roughly the same as for free software. Redistribution (including the normal kinds of commercial redistribution) must be permitted, so that the manual can accompany every copy of the program, both on-line and on paper.

Permission for modification of the technical content is crucial too. When people modify the software, adding or changing features, if they are conscientious they will change the manual too—so they can provide accurate and clear documentation for the modified program. A manual that leaves you no choice but to write a new manual to document a changed version of the program is not really available to our community.

Some kinds of limits on the way modification is handled are acceptable. For example, requirements to preserve the original author’s copyright notice, the distribution terms, or the list of authors, are ok. It is also no problem to require modified versions to include notice that they were modified. Even entire sections that may not be deleted or changed are acceptable, as long as they deal with nontechnical topics (like this one). These kinds of restrictions are acceptable because they don’t obstruct the community’s normal use of the manual.

However, it must be possible to modify all the technical content of the manual, and then distribute the result in all the usual media, through all the usual channels. Otherwise, the restrictions obstruct the use of the manual, it is not free, and we need another manual to replace it.

Please spread the word about this issue. Our community continues to lose manuals to proprietary publishing. If we spread the word that free software needs free reference manuals and free tutorials, perhaps the next person who wants to contribute by writing documentation will realize, before it is too late, that only free manuals contribute to the free software community.

If you are writing documentation, please insist on publishing it under the GNU Free Documentation License or another free documentation license. Remember that this decision requires your approval—you don’t have to let the publisher decide. Some commercial publishers will use a free license if you insist, but they will not propose the option; it is up to you to raise the issue and say firmly that this is what you want. If the publisher you are dealing with refuses, please try other publishers. If you’re not sure whether a proposed license is free, write to licensing@gnu.org.

You can encourage commercial publishers to sell more free, copylefted manuals and tutorials by buying them, and particularly by buying copies from the publishers that paid for their writing or for major improvements. Meanwhile, try to avoid buying non-free documentation at all. Check the distribution terms of a manual before you buy it, and insist that whoever seeks your business must respect your freedom. Check the history of the book, and try to reward the publishers that have paid or pay the authors to work on it.

The Free Software Foundation maintains a list of free documentation published by other

publishers, at http://www.fsf.org/doc/other-free-books.html.

Contributors to [No value for “GDBN”]

Richard Stallman was the original author of [No value for “GDBN”], and of many other gnu programs. Many others have contributed to its development. This section attempts to credit major contributors. One of the virtues of free software is that everyone is free to contribute to it; with regret, we cannot actually acknowledge everyone here. The file ‘ChangeLog’ in the [No value for “GDBN”] distribution approximates a blow-by-blow account.

Changes much prior to version 2.0 are lost in the mists of time.

Plea: Additions to this section are particularly welcome. If you or your friends (or enemies, to be evenhanded) have been unfairly omitted from this list, we would like to add your names!

So that they may not regard their many labors as thankless, we particularly thank those who shepherded [No value for “GDBN”] through major releases: Andrew Cagney (releases 5.0 and 5.1); Jim Blandy (release 4.18); Jason Molenda (release 4.17); Stan Shebs (release 4.14); Fred Fish (releases 4.16, 4.15, 4.13, 4.12, 4.11, 4.10, and 4.9); Stu Grossman and John Gilmore (releases 4.8, 4.7, 4.6, 4.5, and 4.4); John Gilmore (releases 4.3, 4.2, 4.1, 4.0, and 3.9); Jim Kingdon (releases 3.5, 3.4, and 3.3); and Randy Smith (releases 3.2, 3.1, and 3.0).

Richard Stallman, assisted at various times by Peter TerMaat, Chris Hanson, and Richard Mlynarik, handled releases through 2.8.

Michael Tiemann is the author of most of the gnu C++ support in [No value for “GDBN”], with significant additional contributions from Per Bothner and Daniel Berlin. James Clark wrote the gnu C++ demangler. Early work on C++ was by Peter TerMaat (who also did much general update work leading to release 3.0).

[No value for “GDBN”] uses the BFD subroutine library to examine multiple object-file formats; BFD was a joint project of David V. Henkel-Wallace, Rich Pixley, Steve Chamberlain, and John Gilmore.

David Johnson wrote the original COFF support; Pace Willison did the original support for encapsulated COFF.

Brent Benson of Harris Computer Systems contributed DWARF2 support.

Adam de Boor and Bradley Davis contributed the ISI Optimum V support. Per Bothner, Noboyuki Hikichi, and Alessandro Forin contributed MIPS support. Jean-Daniel Fekete contributed Sun 386i support. Chris Hanson improved the HP9000 support. Noboyuki Hikichi and Tomoyuki Hasei contributed Sony/News OS 3 support. David Johnson contributed Encore Umax support. Jyrki Kuoppala contributed Altos 3068 support. Jeff Law contributed HP PA and SOM support. Keith Packard contributed NS32K support.

Doug Rabson contributed Acorn Risc Machine support. Bob Rusk contributed Harris Nighthawk CX-UX support. Chris Smith contributed Convex support (and Fortran debugging). Jonathan Stone contributed Pyramid support. Michael Tiemann contributed SPARC support. Tim Tucker contributed support for the Gould NP1 and Gould Powernode. Pace Willison contributed Intel 386 support. Jay Vosburgh contributed Symmetry support.

Andreas Schwab contributed M68K Linux support.

Rich Schaefer and Peter Schauer helped with support of SunOS shared libraries.

Jay Fenlason and Roland McGrath ensured that [No value for “GDBN”] and GAS agree about several machine instruction sets.

Patrick Duval, Ted Goldstein, Vikram Koka and Glenn Engel helped develop remote debugging. Intel Corporation, Wind River Systems, AMD, and ARM contributed remote debugging modules for the i960, VxWorks, A29K UDI, and RDI targets, respectively.

Brian Fox is the author of the readline libraries providing command-line editing and command history.

Andrew Beers of SUNY Buffalo wrote the language-switching code, the Modula-2 support, and contributed the Languages chapter of this manual.

Fred Fish wrote most of the support for Unix System Vr4. He also enhanced the command-completion support to cover C++ overloaded symbols.

Hitachi America, Ltd. sponsored the support for H8/300, H8/500, and Super-H processors.

NEC sponsored the support for the v850, Vr4xxx, and Vr5xxx processors.

Mitsubishi sponsored the support for D10V, D30V, and M32R/D processors.

Toshiba sponsored the support for the TX39 Mips processor.

Matsushita sponsored the support for the MN10200 and MN10300 processors.

Fujitsu sponsored the support for SPARClite and FR30 processors.

Kung Hsu, Jeff Law, and Rick Sladkey added support for hardware watchpoints.

Michael Snyder added support for tracepoints.

Stu Grossman wrote gdbserver.

Jim Kingdon, Peter Schauer, Ian Taylor, and Stu Grossman made nearly innumerable bug fixes and cleanups throughout [No value for “GDBN”].

The following people at the Hewlett-Packard Company contributed support for the PA-RISC 2.0 architecture, HP-UX 10.20, 10.30, and 11.0 (narrow mode), HP’s implementation of kernel threads, HP’s aC++ compiler, and the terminal user interface: Ben Krepp, Richard Title, John Bishop, Susan Macchia, Kathy Mann, Satish Pai, India Paul, Steve Rehrauer, and Elena Zannoni. Kim Haase provided HP-specific information in this manual.

DJ Delorie ported [No value for “GDBN”] to MS-DOS, for the DJGPP project. Robert Hoehne made significant contributions to the DJGPP port.

Cygnus Solutions has sponsored [No value for “GDBN”] maintenance and much of its development since 1991. Cygnus engineers who have worked on [No value for “GDBN”] fulltime include Mark Alexander, Jim Blandy, Per Bothner, Kevin Buettner, Edith Epstein, Chris Faylor, Fred Fish, Martin Hunt, Jim Ingham, John Gilmore, Stu Grossman, Kung Hsu, Jim Kingdon, John Metzler, Fernando Nasser, Geoffrey Noer, Dawn Perchik, Rich Pixley, Zdenek Radouch, Keith Seitz, Stan Shebs, David Taylor, and Elena Zannoni. In addition, Dave Brolley, Ian Carmichael, Steve Chamberlain, Nick Clifton, JT Conklin, Stan Cox, DJ Delorie, Ulrich Drepper, Frank Eigler, Doug Evans, Sean Fagan, David Henkel-Wallace, Richard Henderson, Jeff Holcomb, Jeff Law, Jim Lemke, Tom Lord, Bob Manson, Michael Meissner, Jason Merrill, Catherine Moore, Drew Moseley, Ken Raeburn, Gavin Romig-Koch, Rob Savoye, Jamie Smith, Mike Stump, Ian Taylor, Angela Thomas, Michael Tiemann, Tom Tromey, Ron Unrau, Jim Wilson, and David Zuhn have made contributions both large and small.

6 Debugging with [No value for “GDBN”]

1 A Sample [No value for “GDBN”] Session

You can use this manual at your leisure to read all about [No value for “GDBN”]. However, a handful of commands are enough to get started using the debugger. This chapter illustrates those commands.

In this sample session, we emphasize user input like this: input, to make it easier to pick out from the surrounding output.

One of the preliminary versions of gnu m4 (a generic macro processor) exhibits the following bug: sometimes, when we change its quote strings from the default, the commands used to capture one macro definition within another stop working. In the following short m4 session, we define a macro foo which expands to 0000; we then use the m4 built-in defn to define bar as the same thing. However, when we change the open quote string to <QUOTE> and the close quote string to <UNQUOTE>, the same procedure fails to define a new synonym

baz:

$ cd gnu/m4 $ ./m4 define(foo,0000)

foo

0000

define(bar,defn(‘foo’))

bar

0000

changequote(<QUOTE>,<UNQUOTE>)

define(baz,defn(<QUOTE>foo<UNQUOTE>)) baz C-d

m4: End of input: 0: fatal error: EOF in string

Let us use [No value for “GDBN”] to try to see what is going on.

$ [No value for “GDBP”] m4

[No value for ``GDBN''] is free software and you are welcome to distribute copies
 of it under certain conditions; type "show copying" to see
 the conditions.
There is absolutely no warranty for [No value for ``GDBN'']; type "show warranty"
 for details.
[No value for ``GDBN''] [No value for ``GDBVN''], Copyright 1999 Free Software Foundation, Inc...
([No value for ``GDBP''])

[No value for “GDBN”] reads only enough symbol data to know where to find the rest when needed; as a result, the first prompt comes up very quickly. We now tell [No value for “GDBN”] to use a narrower display width than usual, so that examples fit in this manual.

([No value for ``GDBP'']) set width 70

We need to see how the m4 built-in changequote works. Having looked at the source, we know the relevant subroutine is m4_changequote, so we set a breakpoint there with the [No value for “GDBN”] break command.

([No value for ``GDBP'']) break m4 changequote
Breakpoint 1 at 0x62f4: file builtin.c, line 879.

Using the run command, we start m4 running under [No value for “GDBN”] control; as long as control does not reach the m4_changequote subroutine, the program runs as usual:

([No value for ``GDBP'']) run
Starting program: /work/Editorial/gdb/gnu/m4/m4

define(foo,0000)

foo

0000

To trigger the breakpoint, we call changequote. [No value for “GDBN”] suspends execution of m4, displaying information about the context where it stops.

changequote(<QUOTE>,<UNQUOTE>)

Breakpoint 1, m4_changequote (argc=3, argv=0x33c70)
    at builtin.c:879
879         if (bad_argc(TOKEN_DATA_TEXT(argv[0]),argc,1,3))

Now we use the command n (next) to advance execution to the next line of the current function.

([No value for ``GDBP'']) n
882         set_quotes((argc >= 2) ? TOKEN_DATA_TEXT(argv[1])\
 : nil,

set_quotes looks like a promising subroutine. We can go into it by using the command s (step) instead of next. step goes to the next line to be executed in any subroutine, so it

steps into set_quotes.
([No value for ``GDBP'']) s
set_quotes (lq=0x34c78 "<QUOTE>", rq=0x34c88 "<UNQUOTE>")
    at input.c:530
530         if (lquote != def_lquote)

The display that shows the subroutine where m4 is now suspended (and its arguments) is called a stack frame display. It shows a summary of the stack. We can use the backtrace command (which can also be spelled bt), to see where we are in the stack as a whole: the backtrace command displays a stack frame for each active subroutine.

([No value for ``GDBP'']) bt
#0  set_quotes (lq=0x34c78 "<QUOTE>", rq=0x34c88 "<UNQUOTE>")
    at input.c:530
#1  0x6344 in m4_changequote (argc=3, argv=0x33c70)
    at builtin.c:882
#2  0x8174 in expand_macro (sym=0x33320) at macro.c:242
#3  0x7a88 in expand_token (obs=0x0, t=209696, td=0xf7fffa30)
    at macro.c:71
#4  0x79dc in expand_input () at macro.c:40
#5  0x2930 in main (argc=0, argv=0xf7fffb20) at m4.c:195

We step through a few more lines to see what happens. The first two times, we can use ‘s’; the next two times we use n to avoid falling into the xstrdup subroutine.

([No value for ``GDBP'']) s
0x3b5c  532         if (rquote != def_rquote)
([No value for ``GDBP'']) s
0x3b80  535         lquote = (lq == nil || *lq == '\0') ?  \
def_lquote : xstrdup(lq);
([No value for ``GDBP'']) n
536         rquote = (rq == nil || *rq == '\0') ? def_rquote\
 : xstrdup(rq);
([No value for ``GDBP'']) n
538         len_lquote = strlen(rquote);

The last line displayed looks a little odd; we can examine the variables lquote and rquote to see if they are in fact the new left and right quotes we specified. We use the command p (print) to see their values.

([No value for ``GDBP'']) p lquote
$1 = 0x35d40 "<QUOTE>"
([No value for ``GDBP'']) p rquote
$2 = 0x35d50 "<UNQUOTE>"

lquote and rquote are indeed the new left and right quotes. To look at some context, we can display ten lines of source surrounding the current line with the l (list) command.

([No value for ``GDBP'']) l
533             xfree(rquote);
534
535         lquote = (lq == nil || *lq == '\0') ? def_lquote\
 : xstrdup (lq);
536         rquote = (rq == nil || *rq == '\0') ? def_rquote\
 : xstrdup (rq);
537
538         len_lquote = strlen(rquote);
539         len_rquote = strlen(lquote);
540     }
541
542     void

Let us step past the two lines that set len_lquote and len_rquote, and then examine the values of those variables.

([No value for ``GDBP'']) n
539         len_rquote = strlen(lquote);
([No value for ``GDBP'']) n
540     }
([No value for ``GDBP'']) p len lquote
$3 = 9
([No value for ``GDBP'']) p len rquote
$4 = 7

That certainly looks wrong, assuming len_lquote and len_rquote are meant to be the lengths of lquote and rquote respectively. We can set them to better values using the p command, since it can print the value of any expression—and that expression can include subroutine calls and assignments.

([No value for ``GDBP'']) p len lquote=strlen(lquote)

$5 = 7

([No value for ``GDBP'']) p len rquote=strlen(rquote)

$6 = 9

Is that enough to fix the problem of using the new quotes with the m4 built-in defn? We can allow m4 to continue executing with the c (continue) command, and then try the example that caused trouble initially:

([No value for ``GDBP'']) c
Continuing.

define(baz,defn(<QUOTE>foo<UNQUOTE>))

baz
0000

Success! The new quotes now work just as well as the default ones. The problem seems to have been just the two typos defining the wrong lengths. We allow m4 exit by giving it an EOF as input:

C-d

Program exited normally.

The message ‘Program exited normally.’ is from [No value for “GDBN”]; it indicates m4 has finished executing. We can end our [No value for “GDBN”] session with the [No value for “GDBN”] quit command.

([No value for ``GDBP'']) quit

2 Getting In and Out of [No value for “GDBN”]

This chapter discusses how to start [No value for “GDBN”], and how to get out of it. The essentials are:

• type ‘[No value for ``GDBP'']’ to start [No value for “GDBN”].

• type quit or C-d to exit.

2.1 Invoking [No value for “GDBN”]

Invoke [No value for “GDBN”] by running the program [No value for ``GDBP'']. Once started, [No value for “GDBN”] reads commands from the terminal until you tell it to exit.

You can also run [No value for ``GDBP''] with a variety of arguments and options, to specify more of your debugging environment at the outset.

The command-line options described here are designed to cover a variety of situations; in some environments, some of these options may effectively be unavailable.

The most usual way to start [No value for “GDBN”] is with one argument, specifying an executable program:

     [No value for ``GDBP''] program

You can also start with both an executable program and a core file specified:

     [No value for ``GDBP''] program core

You can, instead, specify a process ID as a second argument, if you want to debug a running process:

     [No value for ``GDBP''] program 1234

would attach [No value for “GDBN”] to process 1234 (unless you also have a file named ‘1234’; [No value for “GDBN”] does check for a core file first).

Taking advantage of the second command-line argument requires a fairly complete operating system; when you use [No value for “GDBN”] as a remote debugger attached to a bare board, there may not be any notion of “process”, and there is often no way to get a core dump. [No value for “GDBN”] will warn you if it is unable to attach or to read core dumps.

You can run [No value for ``GDBP''] without printing the front material, which describes [No value for “GDBN”]’s non-warranty, by specifying -silent:

     [No value for ``GDBP''] -silent

You can further control how [No value for “GDBN”] starts up by using command-line options. [No value for “GDBN”] itself can remind you of the options available. Type

     [No value for ``GDBP''] -help

to display all available options and briefly describe their use (‘[No value for ``GDBP''] -h’ is a shorter equivalent).

All options and command line arguments you give are processed in sequential order. The order makes a difference when the ‘-x’ option is used.

2.1.1 Choosing files

When [No value for “GDBN”] starts, it reads any arguments other than options as specifying an executable file and core file (or process ID). This is the same as if the arguments were specified by the ‘-se’ and ‘-c’ options respectively. ([No value for “GDBN”] reads the first argument that does not have an associated option flag as equivalent to the ‘-se’ option followed by that argument; and the second argument that does not have an associated option flag, if any, as equivalent to the ‘-c’ option followed by that argument.)

If [No value for “GDBN”] has not been configured to included core file support, such as for most embedded targets, then it will complain about a second argument and ignore it.

Many options have both long and short forms; both are shown in the following list. [No value for “GDBN”] also recognizes the long forms if you truncate them, so long as enough of the option is present to be unambiguous. (If you prefer, you can flag option arguments with ‘--’ rather than ‘-’, though we illustrate the more usual convention.)

-symbols file

-s file Read symbol table from file file.

-exec file

-e file Use file file as the executable file to execute when appropriate, and for examining pure data in conjunction with a core dump.

-se file Read symbol table from file file and use it as the executable file.

-core file

-c file Use file file as a core dump to examine.

-c number Connect to process ID number, as with the attach command (unless there is a file in core-dump format named number, in which case ‘-c’ specifies that file as a core dump to read).

-command file

-x file Execute [No value for “GDBN”] commands from file file. See Section 17.3 [Command files], page 179.

-directory directory

-d directory Add directory to the path to search for source files.

-m

-mapped Warning: this option depends on operating system facilities that are not supported on all systems. If memory-mapped files are available on your system through the mmap system call, you can use this option to have [No value for “GDBN”] write the symbols from your program into a reusable file in the current directory. If the program you are debugging is called ‘/tmp/fred’, the mapped symbol file is ‘/tmp/fred.syms’. Future [No value for “GDBN”] debugging sessions notice the presence of this file, and can quickly map in symbol information from it, rather than reading the symbol table from the executable program.

The ‘.syms’ file is specific to the host machine where [No value for “GDBN”] is run. It holds an exact image of the internal [No value for “GDBN”] symbol table. It cannot be shared across multiple host platforms.

-r

-readnow Read each symbol file’s entire symbol table immediately, rather than the default, which is to read it incrementally as it is needed. This makes startup slower, but makes future operations faster.

You typically combine the -mapped and -readnow options in order to build a ‘.syms’ file that contains complete symbol information. (See Section 13.1 [Commands to specify files], page 123, for information on ‘.syms’ files.) A simple [No value for “GDBN”] invocation to do nothing but build a ‘.syms’ file for future use is:

gdb -batch -nx -mapped -readnow programname

2.1.2 Choosing modes

You can run [No value for “GDBN”] in various alternative modes—for example, in batch mode or quiet mode.

-nx

-n Do not execute commands found in any initialization files (normally called ‘.gdbinit’, or ‘gdb.ini’ on PCs). Normally, [No value for “GDBN”] executes the commands in these files after all the command options and arguments have been processed. See Section 17.3 [Command files], page 179.

-quiet
-silent

-q “Quiet”. Do not print the introductory and copyright messages. These messages are also suppressed in batch mode.

-batch Run in batch mode. Exit with status 0 after processing all the command files specified with ‘-x’ (and all commands from initialization files, if not inhibited with ‘-n’). Exit with nonzero status if an error occurs in executing the [No value for “GDBN”] commands in the command files. Batch mode may be useful for running [No value for “GDBN”] as a filter, for example to download and run a program on another computer; in order to make this more useful, the message

     Program exited normally.

(which is ordinarily issued whenever a program running under [No value for “GDBN”] control terminates) is not issued when running in batch mode.

-nowindows

-nw “No windows”. If [No value for “GDBN”] comes with a graphical user interface (GUI) built in, then this option tells [No value for “GDBN”] to only use the command-line interface. If no GUI is available, this option has no effect.

-windows

-w If [No value for “GDBN”] includes a GUI, then this option requires it to be used if possible.

-cd directory Run [No value for “GDBN”] using directory as its working directory, instead of the current directory.

-fullname

-f gnu Emacs sets this option when it runs [No value for “GDBN”] as a subprocess. It tells [No value for “GDBN”] to output the full file name and line number in a standard, recognizable fashion each time a stack frame is displayed (which includes each time your program stops). This recognizable format looks like two ‘\032’ characters, followed by the file name, line number and character position separated by colons, and a newline. The Emacs-to-[No value for “GDBN”] interface program uses the two ‘\032’ characters as a signal to display the source code for the frame.

-epoch The Epoch Emacs-[No value for “GDBN”] interface sets this option when it runs [No value for “GDBN”] as a subprocess. It tells [No value for “GDBN”] to modify its print routines so as to allow Epoch to display values of expressions in a separate window.

-annotate level

This option sets the annotation level inside [No value for “GDBN”]. Its effect is identical to using ‘set annotate level’ (see Chapter 20 [Annotations], page 191). Annotation level controls how much information does [No value for “GDBN”] print together with its prompt, values of expressions, source lines, and other types of output. Level 0 is the normal, level 1 is for use when [No value for “GDBN”] is run as a subprocess of gnu Emacs, level 2 is the maximum annotation suitable for programs that control [No value for “GDBN”].

-async Use the asynchronous event loop for the command-line interface. [No value for “GDBN”] processes all events, such as user keyboard input, via a special event loop. This allows [No value for “GDBN”] to accept and process user commands in parallel with the debugged process being run^{1}, so you don’t need to wait for control to return to [No value for “GDBN”] before you type the next command. (Note: as of version 5.1, the target side of the asynchronous operation is not yet in place, so ‘-async’ does not work fully yet.) When the standard input is connected to a terminal device, [No value for “GDBN”] uses the asynchronous event loop by default, unless disabled by the ‘-noasync’ option.

-noasync Disable the asynchronous event loop for the command-line interface.

-baud bps

-b bps Set the line speed (baud rate or bits per second) of any serial interface used by [No value for “GDBN”] for remote debugging.

-tty device -t device Run using device for your program’s standard input and output.

1 [No value for “GDBN”] built with djgpp tools for MS-DOS/MS-Windows supports this mode of opera-

tion, but the event loop is suspended when the debuggee runs.

-tui Activate the Terminal User Interface when starting. The Terminal User Interface manages several text windows on the terminal, showing source, assembly, registers and [No value for “GDBN”] command outputs (see Chapter 18 [[No value for “GDBN”] Text User Interface], page 183). Do not use this option if you run [No value for “GDBN”] from Emacs (see Chapter 19 [Using [No value for “GDBN”] under gnu Emacs], page 187).

-interpreter interp

Use the interpreter interp for interface with the controlling program or device. This option is meant to be set by programs which communicate with [No value for “GDBN”] using it as a back end.

‘--interpreter=mi’ (or ‘--interpreter=mi1’) causes [No value for “GDBN”]

to use the gdb/mi interface (see ⟨undefined⟩ [The gdb/mi Interface], page ⟨undefined⟩). The older gdb/mi interface, included in [No value for “GDBN”] version 5.0 can be selected with ‘--interpreter=mi0’.

-write Open the executable and core files for both reading and writing. This is equivalent to the ‘set write on’ command inside [No value for “GDBN”] (see Section 12.6 [Patching], page 122).

-statistics

This option causes [No value for “GDBN”] to print statistics about time and memory usage after it completes each command and returns to the prompt.

-version This option causes [No value for “GDBN”] to print its version number and no-warranty blurb, and exit.

2.2 Quitting [No value for “GDBN”]

quit [expression] q To exit [No value for “GDBN”], use the quit command (abbreviated q), or

type an end-of-file character (usually C-d). If you do not supply expression, [No value for “GDBN”] will terminate normally; otherwise it will terminate using the result of expression as the error code.

An interrupt (often C-c) does not exit from [No value for “GDBN”], but rather terminates the action of any [No value for “GDBN”] command that is in progress and returns to [No value for “GDBN”] command level. It is safe to type the interrupt character at any time because [No value for “GDBN”] does not allow it to take effect until a time when it is safe.

If you have been using [No value for “GDBN”] to control an attached process or device, you can release it with the detach command (see Section 4.7 [Debugging an already-running process], page 27).

2.3 Shell commands

If you need to execute occasional shell commands during your debugging session, there is no need to leave or suspend [No value for “GDBN”]; you can just use the shell command. shell command string

Invoke a standard shell to execute command string. If it exists, the environment variable SHELL determines which shell to run. Otherwise [No value for “GDBN”] uses the default shell (‘/bin/sh’ on Unix systems, ‘COMMAND.COM’ on MS-DOS, etc.).

The utility make is often needed in development environments. You do not have to use the shell command for this purpose in [No value for “GDBN”]:

make make-args Execute the make program with the specified arguments. This is equivalent to ‘shell make make-args’.

3 [No value for “GDBN”] Commands

You can abbreviate a [No value for “GDBN”] command to the first few letters of the command name, if that abbreviation is unambiguous; and you can repeat certain [No value for “GDBN”] commands by typing just^{⟨}\underline{\overline{RET}}^{⟩}. You can also use the^{⟨}\underline{\overline{TAB}}^{⟩} key to get [No value for “GDBN”] to fill out the rest of a word in a command (or to show you the alternatives available, if there is more than one possibility).

3.1 Command syntax

A [No value for “GDBN”] command is a single line of input. There is no limit on how long it can be. It starts with a command name, which is followed by arguments whose meaning depends on the command name. For example, the command step accepts an argument which is the number of times to step, as in ‘step 5’. You can also use the step command with no arguments. Some commands do not allow any arguments.

[No value for “GDBN”] command names may always be truncated if that abbreviation is unambiguous. Other possible command abbreviations are listed in the documentation for individual commands. In some cases, even ambiguous abbreviations are allowed; for example, s is specially defined as equivalent to step even though there are other commands whose names start with s. You can test abbreviations by using them as arguments to the help command.

A blank line as input to [No value for “GDBN”] (typing just ^{⟨}\underline{\overline{RET}}^{⟩}) means to repeat the previous command. Certain commands (for example, run) will not repeat this way; these are commands whose unintentional repetition might cause trouble and which you are unlikely to want to repeat.

The list and x commands, when you repeat them with ^{⟨}\underline{\overline{RET}}^{⟩}, construct new arguments rather than repeating exactly as typed. This permits easy scanning of source or memory.

[No value for “GDBN”] can also use ^{⟨}\underline{\overline{RET}}^{⟩} in another way: to partition lengthy output, in a way similar to the common utility more (see Section 16.4 [Screen size], page 173). Since it is easy to press one ^{⟨}\underline{\overline{RET}}^{⟩} too many in this situation, [No value for “GDBN”] disables command repetition after any command that generates this sort of display.

Any text from a # to the end of the line is a comment; it does nothing. This is useful mainly in command files (see Section 17.3 [Command files], page 179).

3.2 Command completion

[No value for “GDBN”] can fill in the rest of a word in a command for you, if there is only one possibility; it can also show you what the valid possibilities are for the next word in a command, at any time. This works for [No value for “GDBN”] commands, [No value for “GDBN”] subcommands, and the names of symbols in your program.

Press the ^{⟨}\underline{\overline{TAB}}^{⟩} key whenever you want [No value for “GDBN”] to fill out the rest of a word. If there is only one possibility, [No value for “GDBN”] fills in the word, and waits for you to finish the command (or press ^{⟨}\underline{\overline{RET}}^{⟩} to enter it). For example, if you type

     ([No value for ``GDBP'']) info bre ^{⟨}\underline{\overline{TAB}}^{⟩}

[No value for “GDBN”] fills in the rest of the word ‘breakpoints’, since that is the only info subcommand beginning with ‘bre’:

     ([No value for ``GDBP'']) info breakpoints

You can either press ^{⟨}\underline{\overline{RET}}^{⟩} at this point, to run the info breakpoints command, or backspace and enter something else, if ‘breakpoints’ does not look like the command you expected. (If you were sure you wanted info breakpoints in the first place, you might as well just type ^{⟨}\underline{\overline{RET}}^{⟩} immediately after ‘info bre’, to exploit command abbreviations rather than command completion).

If there is more than one possibility for the next word when you press ^{⟨}\underline{\overline{TAB}}^{⟩}, [No value for “GDBN”] sounds a bell. You can either supply more characters and try again, or just press ^{⟨}\underline{\overline{TAB}}^{⟩} a second time; [No value for “GDBN”] displays all the possible completions for that word. For example, you might want to set a breakpoint on a subroutine whose name begins with ‘make_’, but when you type b make_^{⟨}\underline{\overline{TAB}}^{⟩} [No value for “GDBN”] just sounds the bell. Typing ^{⟨}\underline{\overline{TAB}}^{⟩} again displays all the function names in your program that begin with those characters, for example:

     ([No value for ``GDBP'']) b make_^{⟨}\underline{\overline{TAB}}^{⟩}

[No value for “GDBN”] sounds bell; press ^{⟨}\underline{\overline{TAB}}^{⟩} again, to see:

     make_a_section_from_file     make_environ
     make_abs_section             make_function_type
     make_blockvector             make_pointer_type
     make_cleanup                 make_reference_type
     make_command                 make_symbol_completion_list
     ([No value for ``GDBP'']) b make_

After displaying the available possibilities, [No value for “GDBN”] copies your partial input (‘b make_’ in the example) so you can finish the command.

If you just want to see the list of alternatives in the first place, you can press M-? rather than pressing ^{⟨}\underline{\overline{TAB}}^{⟩} twice. M-? means ^{⟨}\underline{\overline{META}}^{⟩} ?. You can type this either by holding down a key designated as the ^{⟨}\underline{\overline{META}}^{⟩} shift on your keyboard (if there is one) while typing ?, or as ⟨\underline{\overline{ESC}}^{⟩} followed by ?.

Sometimes the string you need, while logically a “word”, may contain parentheses or other characters that [No value for “GDBN”] normally excludes from its notion of a word. To permit word completion to work in this situation, you may enclose words in ' (single quote marks) in [No value for “GDBN”] commands.

The most likely situation where you might need this is in typing the name of a C++ function. This is because C++ allows function overloading (multiple definitions of the same function, distinguished by argument type). For example, when you want to set a breakpoint you may need to distinguish whether you mean the version of name that takes an int parameter, name(int), or the version that takes a float parameter, name(float). To use the word-completion facilities in this situation, type a single quote ' at the beginning of the function name. This alerts [No value for “GDBN”] that it may need to consider more information than usual when you press ^{⟨}\underline{\overline{TAB}}^{⟩} or M-? to request word completion:

     ([No value for ``GDBP'']) b 'bubble( M-?
     bubble(double,double)    bubble(int,int)
     ([No value for ``GDBP'']) b 'bubble(

In some cases, [No value for “GDBN”] can tell that completing a name requires using quotes. When this happens, [No value for “GDBN”] inserts the quote for you (while completing as much as it can) if you do not type the quote in the first place:

([No value for ``GDBP'']) b bub ^{⟨}\underline{\overline{TAB}}^{⟩}

[No value for “GDBN”] alters your input line to the following, and rings a bell:

([No value for ``GDBP'']) b 'bubble(

In general, [No value for “GDBN”] can tell that a quote is needed (and inserts it) if you have not yet started typing the argument list when you ask for completion on an overloaded symbol.

For more information about overloaded functions, see Section 10.4.1.3 [C++ expressions], page 100. You can use the command set overload-resolution off to disable overload resolution; see Section 10.4.1.7 [[No value for “GDBN”] features for C++], page 102.

3.3 Getting help

You can always ask [No value for “GDBN”] itself for information on its commands, using the command help.

help

h You can use help (abbreviated h) with no arguments to display a short list of

named classes of commands:

([No value for ``GDBP'']) help
List of classes of commands:
aliases -- Aliases of other commands
breakpoints -- Making program stop at certain points
data -- Examining data
files -- Specifying and examining files
internals -- Maintenance commands
obscure -- Obscure features
running -- Running the program
stack -- Examining the stack
status -- Status inquiries
support -- Support facilities
tracepoints -- Tracing of program execution without
               stopping the program
user-defined -- User-defined commands
Type "help" followed by a class name for a list of
commands in that class.
Type "help" followed by command name for full
documentation.
Command name abbreviations are allowed if unambiguous.
([No value for ``GDBP''])

help class Using one of the general help classes as an argument, you can get a list of the individual commands in that class. For example, here is the help display for the class status:

([No value for ``GDBP'']) help status
Status inquiries.

     List of commands:
     info -- Generic command for showing things
      about the program being debugged
     show -- Generic command for showing things
      about the debugger
     Type "help" followed by command name for full
     documentation.
     Command name abbreviations are allowed if unambiguous.
     ([No value for ``GDBP''])

help command With a command name as help argument, [No value for “GDBN”] displays a short paragraph on how to use that command.

apropos args

The apropos args command searches through all of the [No value for “GDBN”] commands, and their documentation, for the regular expression specified in args. It prints out all matches found. For example:

     apropos reload

results in:

     set symbol-reloading -- Set dynamic symbol table reloading
                                      multiple times in one run
     show symbol-reloading -- Show dynamic symbol table reloading
                                      multiple times in one run
complete args

The complete args command lists all the possible completions for the beginning of a command. Use args to specify the beginning of the command you want completed. For example:

     complete i

results in:

     if ignore info inspect

This is intended for use by gnu Emacs.

In addition to help, you can use the [No value for “GDBN”] commands info and show to inquire about the state of your program, or the state of [No value for “GDBN”] itself. Each command supports many topics of inquiry; this manual introduces each of them in the appropriate context. The listings under info and under show in the Index point to all the sub-commands. See [Index], page 211.

info This command (abbreviated i) is for describing the state of your program. For example, you can list the arguments given to your program with info args, list the registers currently in use with info registers, or list the breakpoints you have set with info breakpoints. You can get a complete list of the info sub-commands with help info.

set You can assign the result of an expression to an environment variable with set. For example, you can set the [No value for “GDBN”] prompt to a $-sign with

set prompt $.

show In contrast to info, show is for describing the state of [No value for “GDBN”] itself. You can change most of the things you can show, by using the related command set; for example, you can control what number system is used for displays with set radix, or simply inquire which is currently in use with show

radix.

To display all the settable parameters and their current values, you can use show with no arguments; you may also use info set. Both commands produce the same display.

Here are three miscellaneous show subcommands, all of which are exceptional in lacking corresponding set commands:

show version

Show what version of [No value for “GDBN”] is running. You should include this information in [No value for “GDBN”] bug-reports. If multiple versions of [No value for “GDBN”] are in use at your site, you may need to determine which version of [No value for “GDBN”] you are running; as [No value for “GDBN”] evolves, new commands are introduced, and old ones may wither away. Also, many system vendors ship variant versions of [No value for “GDBN”], and there are variant versions of [No value for “GDBN”] in gnu/Linux distributions as well. The version number is the same as the one announced when you start [No value for “GDBN”].

show copying

Display information about permission for copying [No value for “GDBN”].

show warranty

Display the gnu “NO WARRANTY” statement, or a warranty, if your version of [No value for “GDBN”] comes with one.

22 Debugging with [No value for “GDBN”]

4 Running Programs Under [No value for “GDBN”]

When you run a program under [No value for “GDBN”], you must first generate debugging information when you compile it.

You may start [No value for “GDBN”] with its arguments, if any, in an environment of your choice. If you are doing native debugging, you may redirect your program’s input and output, debug an already running process, or kill a child process.

4.1 Compiling for debugging

In order to debug a program effectively, you need to generate debugging information when you compile it. This debugging information is stored in the object file; it describes the data type of each variable or function and the correspondence between source line numbers and addresses in the executable code.

To request debugging information, specify the ‘-g’ option when you run the compiler. Many C compilers are unable to handle the ‘-g’ and ‘-O’ options together. Using those

compilers, you cannot generate optimized executables containing debugging information.

[No value for “NGCC”], the gnu C compiler, supports ‘-g’ with or without ‘-O’, making it possible to debug optimized code. We recommend that you always use ‘-g’ whenever you compile a program. You may think your program is correct, but there is no sense in pushing your luck.

When you debug a program compiled with ‘-g -O’, remember that the optimizer is rearranging your code; the debugger shows you what is really there. Do not be too surprised when the execution path does not exactly match your source file! An extreme example: if you define a variable, but never use it, [No value for “GDBN”] never sees that variable— because the compiler optimizes it out of existence.

Some things do not work as well with ‘-g -O’ as with just ‘-g’, particularly on machines with instruction scheduling. If in doubt, recompile with ‘-g’ alone, and if this fixes the problem, please report it to us as a bug (including a test case!).

Older versions of the gnu C compiler permitted a variant option ‘-gg’ for debugging information. [No value for “GDBN”] no longer supports this format; if your gnu C compiler has this option, do not use it.

4.2 Starting your program

run

r Use the run command to start your program under [No value for “GDBN”]. You must first specify the program name (except on VxWorks) with an argument to [No value for “GDBN”] (see Chapter 2 [Getting In and Out of [No value for “GDBN”]], page 11), or by using the file or exec-file command (see Section 13.1 [Commands to specify files], page 123).

If you are running your program in an execution environment that supports processes, run creates an inferior process and makes that process run your program. (In environments without processes, run jumps to the start of your program.)

The execution of a program is affected by certain information it receives from its superior. [No value for “GDBN”] provides ways to specify this information, which you must do before starting your program. (You can change it after starting your program, but such changes only affect your program the next time you start it.) This information may be divided into four categories:

The arguments.

Specify the arguments to give your program as the arguments of the run command. If a shell is available on your target, the shell is used to pass the arguments, so that you may use normal conventions (such as wildcard expansion or variable substitution) in describing the arguments. In Unix systems, you can control which shell is used with the SHELL environment variable. See Section 4.3 [Your program’s arguments], page 24.

The environment.

Your program normally inherits its environment from [No value for “GDBN”], but you can use the [No value for “GDBN”] commands set environment and unset environment to change parts of the environment that affect your program. See Section 4.4 [Your program’s environment], page 25.

The working directory.

Your program inherits its working directory from [No value for “GDBN”]. You can set the [No value for “GDBN”] working directory with the cd command in [No value for “GDBN”]. See Section 4.5 [Your program’s working directory], page 26.

The standard input and output.

Your program normally uses the same device for standard input and standard output as [No value for “GDBN”] is using. You can redirect input and output in the run command line, or you can use the tty command to set a different device for your program. See Section 4.6 [Your program’s input and output], page 26. Warning: While input and output redirection work, you cannot use pipes to pass the output of the program you are debugging to another program; if you attempt this, [No value for “GDBN”] is likely to wind up debugging the wrong program.

When you issue the run command, your program begins to execute immediately. See Chapter 5 [Stopping and continuing], page 33, for discussion of how to arrange for your program to stop. Once your program has stopped, you may call functions in your program, using the print or call commands. See Chapter 8 [Examining Data], page 65.

If the modification time of your symbol file has changed since the last time [No value for “GDBN”] read its symbols, [No value for “GDBN”] discards its symbol table, and reads it again. When it does this, [No value for “GDBN”] tries to retain your current breakpoints.

4.3 Your program’s arguments

The arguments to your program can be specified by the arguments of the run command. They are passed to a shell, which expands wildcard characters and performs redirection of I/O, and thence to your program. Your SHELL environment variable (if it exists) specifies what shell [No value for “GDBN”] uses. If you do not define SHELL, [No value for “GDBN”] uses the default shell (‘/bin/sh’ on Unix).

On non-Unix systems, the program is usually invoked directly by [No value for “GDBN”], which emulates I/O redirection via the appropriate system calls, and the wildcard characters are expanded by the startup code of the program, not by the shell.

run with no arguments uses the same arguments used by the previous run, or those set by the set args command.

set args Specify the arguments to be used the next time your program is run. If set args has no arguments, run executes your program with no arguments. Once you have run your program with arguments, using set args before the next run is the only way to run it again without arguments.

show args Show the arguments to give your program when it is started.

4.4 Your program’s environment

The environment consists of a set of environment variables and their values. Environment variables conventionally record such things as your user name, your home directory, your terminal type, and your search path for programs to run. Usually you set up environment variables with the shell and they are inherited by all the other programs you run. When debugging, it can be useful to try running your program with a modified environment without having to start [No value for “GDBN”] over again.

path directory Add directory to the front of the PATH environment variable (the search path for executables) that will be passed to your program. The value of PATH used by [No value for “GDBN”] does not change. You may specify several directory names, separated by whitespace or by a system-dependent separator character (‘:’ on Unix, ‘;’ on MS-DOS and MS-Windows). If directory is already in the path, it is moved to the front, so it is searched sooner. You can use the string ‘$cwd’ to refer to whatever is the current working directory at the time [No value for “GDBN”] searches the path. If you use ‘.’ instead, it refers to the directory where you executed the path command. [No value for “GDBN”] replaces ‘.’ in the directory argument (with the current path) before adding directory to the search path.

show paths

Display the list of search paths for executables (the PATH environment variable).

show environment [varname]

Print the value of environment variable varname to be given to your program when it starts. If you do not supply varname, print the names and values of all environment variables to be given to your program. You can abbreviate

environment as env.

set environment varname [=value]

Set environment variable varname to value. The value changes for your program only, not for [No value for “GDBN”] itself. value may be any string; the values of environment variables are just strings, and any interpretation is supplied by your program itself. The value parameter is optional; if it is eliminated, the variable is set to a null value. For example, this command:

               set env USER = foo

tells the debugged program, when subsequently run, that its user is named ‘foo’. (The spaces around ‘=’ are used for clarity here; they are not actually required.)

unset environment varname

Remove variable varname from the environment to be passed to your program. This is different from ‘set env varname =’; unset environment removes the variable from the environment, rather than assigning it an empty value.

Warning: On Unix systems, [No value for “GDBN”] runs your program using the shell indicated by your SHELL environment variable if it exists (or /bin/sh if not). If your SHELL variable names a shell that runs an initialization file—such as ‘.cshrc’ for C-shell, or ‘.bashrc’ for BASH—any variables you set in that file affect your program. You may wish to move setting of environment variables to files that are only run when you sign on,

such as ‘.login’ or ‘.profile’.

4.5 Your program’s working directory

Each time you start your program with run, it inherits its working directory from the current working directory of [No value for “GDBN”]. The [No value for “GDBN”] working directory is initially whatever it inherited from its parent process (typically the shell), but you can specify a new working directory in [No value for “GDBN”] with the cd command.

The [No value for “GDBN”] working directory also serves as a default for the commands that specify files for [No value for “GDBN”] to operate on. See Section 13.1 [Commands to specify files], page 123.

cd directory Set the [No value for “GDBN”] working directory to directory.

pwd Print the [No value for “GDBN”] working directory.

4.6 Your program’s input and output

By default, the program you run under [No value for “GDBN”] does input and output to the same terminal that [No value for “GDBN”] uses. [No value for “GDBN”] switches the terminal to its own terminal modes to interact with you, but it records the terminal modes your program was using and switches back to them when you continue running your program.

info terminal

Displays information recorded by [No value for “GDBN”] about the terminal modes your program is using.

You can redirect your program’s input and/or output using shell redirection with the run command. For example,

     run > outfile

starts your program, diverting its output to the file ‘outfile’.

Another way to specify where your program should do input and output is with the tty command. This command accepts a file name as argument, and causes this file to be the default for future run commands. It also resets the controlling terminal for the child process, for future run commands. For example,

     tty /dev/ttyb

directs that processes started with subsequent run commands default to do input and output on the terminal ‘/dev/ttyb’ and have that as their controlling terminal.

An explicit redirection in run overrides the tty command’s effect on the input/output device, but not its effect on the controlling terminal.

When you use the tty command or redirect input in the run command, only the input for your program is affected. The input for [No value for “GDBN”] still comes from your terminal.

4.7 Debugging an already-running process

attach process-id This command attaches to a running process—one that was started outside [No value for “GDBN”]. (info files shows your active targets.) The command takes as argument a process ID. The usual way to find out the process-id of a Unix process is with the ps utility, or with the ‘jobs -l’ shell command. attach does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} a second time after executing the command.

To use attach, your program must be running in an environment which supports processes; for example, attach does not work for programs on bare-board targets that lack an operating system. You must also have permission to send the process a signal.

When you use attach, the debugger finds the program running in the process first by looking in the current working directory, then (if the program is not found) by using the source file search path (see Section 7.3 [Specifying source directories], page 61). You can also use the file command to load the program. See Section 13.1 [Commands to Specify Files], page 123.

The first thing [No value for “GDBN”] does after arranging to debug the specified process is to stop it. You can examine and modify an attached process with all the [No value for “GDBN”] commands that are ordinarily available when you start processes with run. You can insert breakpoints; you can step and continue; you can modify storage. If you would rather the process continue running, you may use the continue command after attaching [No value for “GDBN”] to the process. detach When you have finished debugging the attached process, you can use the detach

command to release it from [No value for “GDBN”] control. Detaching the process continues its execution. After the detach command, that process and [No value for “GDBN”] become completely independent once more, and you are ready to attach another process or start one with run. detach does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} again after executing the command.

If you exit [No value for “GDBN”] or use the run command while you have an attached process, you kill that process. By default, [No value for “GDBN”] asks for confirmation if you try to do either of these things; you can control whether or not you need to confirm by using the set confirm command (see Section 16.6 [Optional warnings and messages], page 174).

4.8 Killing the child process

kill Kill the child process in which your program is running under [No value for “GDBN”].

This command is useful if you wish to debug a core dump instead of a running process. [No value for “GDBN”] ignores any core dump file while your program is running.

On some operating systems, a program cannot be executed outside [No value for “GDBN”] while you have breakpoints set on it inside [No value for “GDBN”]. You can use the kill command in this situation to permit running your program outside the debugger.

The kill command is also useful if you wish to recompile and relink your program, since on many systems it is impossible to modify an executable file while it is running in a process. In this case, when you next type run, [No value for “GDBN”] notices that the file has changed, and reads the symbol table again (while trying to preserve your current breakpoint settings).

4.9 Debugging programs with multiple threads

In some operating systems, such as HP-UX and Solaris, a single program may have more than one thread of execution. The precise semantics of threads differ from one operating system to another, but in general the threads of a single program are akin to multiple processes—except that they share one address space (that is, they can all examine and modify the same variables). On the other hand, each thread has its own registers and execution stack, and perhaps private memory.

[No value for “GDBN”] provides these facilities for debugging multi-thread programs:

• automatic notification of new threads

• ‘thread threadno’, a command to switch among threads

• ‘info threads’, a command to inquire about existing threads

• ‘thread apply [threadno] [all] args’, a command to apply a command to a list of threads

• thread-specific breakpoints Warning: These facilities are not yet available on every [No value for “GDBN”] configuration where the operating system supports threads. If your [No value for “GDBN”] does not support threads, these commands have no effect. For example, a system without thread support shows no output from ‘info threads’, and always rejects the thread command, like this:

          ([No value for ``GDBP'']) info threads
          ([No value for ``GDBP'']) thread 1
          Thread ID 1 not known.  Use the "info threads" command to
          see the IDs of currently known threads.

The [No value for “GDBN”] thread debugging facility allows you to observe all threads while your program runs—but whenever [No value for “GDBN”] takes control, one thread in particular is always the focus of debugging. This thread is called the current thread. Debugging commands show program information from the perspective of the current thread.

Whenever [No value for “GDBN”] detects a new thread in your program, it displays the target system’s identification for the thread with a message in the form ‘[New systag]’. systag is a thread identifier whose form varies depending on the particular system. For example, on LynxOS, you might see

     [New process 35 thread 27]

when [No value for “GDBN”] notices a new thread. In contrast, on an SGI system, the systag is simply something like ‘process 368’, with no further qualifier.

For debugging purposes, [No value for “GDBN”] associates its own thread number— always a single integer—with each thread in your program.

info threads

Display a summary of all threads currently in your program. [No value for “GDBN”] displays for each thread (in this order):

1. the thread number assigned by [No value for “GDBN”]

2. the target system’s thread identifier (systag)

3. the current stack frame summary for that thread

An asterisk ‘*’ to the left of the [No value for “GDBN”] thread number indicates the current thread. For example,

     ([No value for ``GDBP'']) info threads
       3 process 35 thread 27  0x34e5 in sigpause ()
       2 process 35 thread 23  0x34e5 in sigpause ()
     * 1 process 35 thread 13  main (argc=1, argv=0x7ffffff8)
         at threadtest.c:68

On HP-UX systems:

For debugging purposes, [No value for “GDBN”] associates its own thread number—a small integer assigned in thread-creation order—with each thread in your program.

Whenever [No value for “GDBN”] detects a new thread in your program, it displays both [No value for “GDBN”]’s thread number and the target system’s identification for the thread with a message in the form ‘[New systag]’. systag is a thread identifier whose form varies depending on the particular system. For example, on HP-UX, you see

[New thread 2 (system thread 26594)]

when [No value for “GDBN”] notices a new thread.

info threads

Display a summary of all threads currently in your program. [No value for “GDBN”] displays for each thread (in this order):

1. the thread number assigned by [No value for “GDBN”]

2. the target system’s thread identifier (systag)

3. the current stack frame summary for that thread

An asterisk ‘*’ to the left of the [No value for “GDBN”] thread number indicates the current thread. For example,

([No value for ``GDBP'']) info threads
* 3 system thread 26607  worker (wptr=0x7b09c318 "@") \
                          at quicksort.c:137
 2 system thread 26606  0x7b0030d8 in __ksleep () \
                          from /usr/lib/libc.2
 1 system thread 27905  0x7b003498 in _brk () \
                          from /usr/lib/libc.2

thread threadno Make thread number threadno the current thread. The command argument threadno is the internal [No value for “GDBN”] thread number, as shown in the first field of the ‘info threads’ display. [No value for “GDBN”] responds by displaying the system identifier of the thread you selected, and its current stack frame summary:

     ([No value for ``GDBP'']) thread 2
     [Switching to process 35 thread 23]
     0x34e5 in sigpause ()

As with the ‘[New ...]’ message, the form of the text after ‘Switching to’ depends on your system’s conventions for identifying threads.

thread apply [threadno] [all] args

The thread apply command allows you to apply a command to one or more threads. Specify the numbers of the threads that you want affected with the command argument threadno. threadno is the internal [No value for “GDBN”] thread number, as shown in the first field of the ‘info threads’ display. To apply a command to all threads, use thread apply all args.

Whenever [No value for “GDBN”] stops your program, due to a breakpoint or a signal, it automatically selects the thread where that breakpoint or signal happened. [No value for “GDBN”] alerts you to the context switch with a message of the form ‘[Switching to systag]’ to identify the thread.

See Section 5.4 [Stopping and starting multi-thread programs], page 50, for more information about how [No value for “GDBN”] behaves when you stop and start programs with multiple threads.

See Section 5.1.2 [Setting watchpoints], page 37, for information about watchpoints in programs with multiple threads.

4.10 Debugging programs with multiple processes

On most systems, [No value for “GDBN”] has no special support for debugging programs which create additional processes using the fork function. When a program forks, [No value for “GDBN”] will continue to debug the parent process and the child process will run unimpeded. If you have set a breakpoint in any code which the child then executes, the child will get a SIGTRAP signal which (unless it catches the signal) will cause it to terminate.

However, if you want to debug the child process there is a workaround which isn’t too painful. Put a call to sleep in the code which the child process executes after the fork. It may be useful to sleep only if a certain environment variable is set, or a certain file exists, so that the delay need not occur when you don’t want to run [No value for “GDBN”] on the child. While the child is sleeping, use the ps program to get its process ID. Then tell [No value for “GDBN”] (a new invocation of [No value for “GDBN”] if you are also debugging the parent process) to attach to the child process (see Section 4.7 [Attach], page 27). From that point on you can debug the child process just like any other process which you attached to.

On HP-UX (11.x and later only?), [No value for “GDBN”] provides support for debugging programs that create additional processes using the fork or vfork function.

By default, when a program forks, [No value for “GDBN”] will continue to debug the parent process and the child process will run unimpeded.

If you want to follow the child process instead of the parent process, use the command

set follow-fork-mode.
set follow-fork-mode mode

Set the debugger response to a program call of fork or vfork. A call to fork or vfork creates a new process. The mode can be:

parent The original process is debugged after a fork. The child process runs unimpeded. This is the default.

child The new process is debugged after a fork. The parent process runs unimpeded.

ask The debugger will ask for one of the above choices.

show follow-fork-mode

Display the current debugger response to a fork or vfork call.

If you ask to debug a child process and a vfork is followed by an exec, [No value for “GDBN”] executes the new target up to the first breakpoint in the new target. If you have a breakpoint set on main in your original program, the breakpoint will also be set on the child process’s main.

When a child process is spawned by vfork, you cannot debug the child or parent until an exec call completes.

If you issue a run command to [No value for “GDBN”] after an exec call executes, the new target restarts. To restart the parent process, use the file command with the parent executable name as its argument.

You can use the catch command to make [No value for “GDBN”] stop whenever a fork, vfork, or exec call is made. See Section 5.1.3 [Setting catchpoints], page 39.

5 Stopping and Continuing

The principal purposes of using a debugger are so that you can stop your program before it terminates; or so that, if your program runs into trouble, you can investigate and find out why.

Inside [No value for “GDBN”], your program may stop for any of several reasons, such as a signal, a breakpoint, or reaching a new line after a [No value for “GDBN”] command such as step. You may then examine and change variables, set new breakpoints or remove old ones, and then continue execution. Usually, the messages shown by [No value for “GDBN”] provide ample explanation of the status of your program—but you can also explicitly request this information at any time.

info program

Display information about the status of your program: whether it is running or not, what process it is, and why it stopped.

5.1 Breakpoints, watchpoints, and catchpoints

A breakpoint makes your program stop whenever a certain point in the program is reached. For each breakpoint, you can add conditions to control in finer detail whether your program stops. You can set breakpoints with the break command and its variants (see Section 5.1.1 [Setting breakpoints], page 34), to specify the place where your program should stop by line number, function name or exact address in the program.

In HP-UX, SunOS 4.x, SVR4, and Alpha OSF/1 configurations, you can set breakpoints in shared libraries before the executable is run. There is a minor limitation on HP-UX systems: you must wait until the executable is run in order to set breakpoints in shared library routines that are not called directly by the program (for example, routines that are arguments in a pthread_create call).

A watchpoint is a special breakpoint that stops your program when the value of an expression changes. You must use a different command to set watchpoints (see Section 5.1.2 [Setting watchpoints], page 37), but aside from that, you can manage a watchpoint like any other breakpoint: you enable, disable, and delete both breakpoints and watchpoints using the same commands.

You can arrange to have values from your program displayed automatically whenever [No value for “GDBN”] stops at a breakpoint. See Section 8.6 [Automatic display], page 70.

A catchpoint is another special breakpoint that stops your program when a certain kind of event occurs, such as the throwing of a C++ exception or the loading of a library. As with watchpoints, you use a different command to set a catchpoint (see Section 5.1.3 [Setting catchpoints], page 39), but aside from that, you can manage a catchpoint like any other breakpoint. (To stop when your program receives a signal, use the handle command; see Section 5.3 [Signals], page 49.)

[No value for “GDBN”] assigns a number to each breakpoint, watchpoint, or catchpoint when you create it; these numbers are successive integers starting with one. In many of the commands for controlling various features of breakpoints you use the breakpoint number to say which breakpoint you want to change. Each breakpoint may be enabled or disabled; if disabled, it has no effect on your program until you enable it again.

Some [No value for “GDBN”] commands accept a range of breakpoints on which to operate. A breakpoint range is either a single breakpoint number, like ‘5’, or two such numbers, in increasing order, separated by a hyphen, like ‘5-7’. When a breakpoint range is given to a command, all breakpoint in that range are operated on.

5.1.1 Setting breakpoints

Breakpoints are set with the break command (abbreviated b). The debugger convenience variable ‘$bpnum’ records the number of the breakpoint you’ve set most recently; see Section 8.9 [Convenience variables], page 78, for a discussion of what you can do with convenience variables.

You have several ways to say where the breakpoint should go.

break function Set a breakpoint at entry to function function. When using source languages that permit overloading of symbols, such as C++, function may refer to more than one possible place to break. See Section 5.1.8 [Breakpoint menus], page 45, for a discussion of that situation.

break +offset
break -offset

Set a breakpoint some number of lines forward or back from the position at which execution stopped in the currently selected stack frame. (See Section 6.1 [Frames], page 53, for a description of stack frames.)

break linenum Set a breakpoint at line linenum in the current source file. The current source file is the last file whose source text was printed. The breakpoint will stop your program just before it executes any of the code on that line.

break filename:linenum Set a breakpoint at line linenum in source file filename.

break filename:function Set a breakpoint at entry to function function found in file filename. Specifying a file name as well as a function name is superfluous except when multiple files contain similarly named functions.

break *address Set a breakpoint at address address. You can use this to set breakpoints in parts of your program which do not have debugging information or source files.

break When called without any arguments, break sets a breakpoint at the next instruction to be executed in the selected stack frame (see Chapter 6 [Examining the Stack], page 53). In any selected frame but the innermost, this makes your program stop as soon as control returns to that frame. This is similar to the effect of a finish command in the frame inside the selected frame—except that finish does not leave an active breakpoint. If you use break without an argument in the innermost frame, [No value for “GDBN”] stops the next time it reaches the current location; this may be useful inside loops.

[No value for “GDBN”] normally ignores breakpoints when it resumes execution, until at least one instruction has been executed. If it did not do this, you would be unable to proceed past a breakpoint without first disabling the breakpoint. This rule applies whether or not the breakpoint already existed when your program stopped.

break ... if cond

Set a breakpoint with condition cond; evaluate the expression cond each time the breakpoint is reached, and stop only if the value is nonzero—that is, if cond evaluates as true. ‘...’ stands for one of the possible arguments described above (or no argument) specifying where to break. See Section 5.1.6 [Break conditions], page 42, for more information on breakpoint conditions.

tbreak args

Set a breakpoint enabled only for one stop. args are the same as for the break command, and the breakpoint is set in the same way, but the breakpoint is automatically deleted after the first time your program stops there. See Section 5.1.5 [Disabling breakpoints], page 41.

hbreak args

Set a hardware-assisted breakpoint. args are the same as for the break command and the breakpoint is set in the same way, but the breakpoint requires hardware support and some target hardware may not have this support. The main purpose of this is EPROM/ROM code debugging, so you can set a breakpoint at an instruction without changing the instruction. This can be used with the new trap-generation provided by SPARClite DSU and some x86-based targets. These targets will generate traps when a program accesses some data or instruction address that is assigned to the debug registers. However the hardware breakpoint registers can take a limited number of breakpoints. For example, on the DSU, only two data breakpoints can be set at a time, and [No value for “GDBN”] will reject this command if more than two are used. Delete or disable unused hardware breakpoints before setting new ones (see Section 5.1.5 [Disabling], page 41). See Section 5.1.6 [Break conditions], page 42.

thbreak args

Set a hardware-assisted breakpoint enabled only for one stop. args are the same as for the hbreak command and the breakpoint is set in the same way. However, like the tbreak command, the breakpoint is automatically deleted after the first time your program stops there. Also, like the hbreak command, the breakpoint requires hardware support and some target hardware may not have this support. See Section 5.1.5 [Disabling breakpoints], page 41. See also Section 5.1.6 [Break conditions], page 42.

rbreak regex

Set breakpoints on all functions matching the regular expression regex. This command sets an unconditional breakpoint on all matches, printing a list of all breakpoints it set. Once these breakpoints are set, they are treated just like the breakpoints set with the break command. You can delete them, disable them, or make them conditional the same way as any other breakpoint.

The syntax of the regular expression is the standard one used with tools like ‘grep’. Note that this is different from the syntax used by shells, so for instance foo* matches all functions that include an fo followed by zero or more os. There is an implicit .* leading and trailing the regular expression you supply, so to match only functions that begin with foo, use ^foo. When debugging C++ programs, rbreak is useful for setting breakpoints on overloaded functions that are not members of any special classes.

info breakpoints [n]
info break [n]
info watchpoints [n]

Print a table of all breakpoints, watchpoints, and catchpoints set and not deleted, with the following columns for each breakpoint:

Breakpoint Numbers Type Breakpoint, watchpoint, or catchpoint.

Disposition Whether the breakpoint is marked to be disabled or deleted when hit.

Enabled or Disabled Enabled breakpoints are marked with ‘y’. ‘n’ marks breakpoints that are not enabled.

Address Where the breakpoint is in your program, as a memory address.

What Where the breakpoint is in the source for your program, as a file and line number.

If a breakpoint is conditional, info break shows the condition on the line following the affected breakpoint; breakpoint commands, if any, are listed after that. info break with a breakpoint number n as argument lists only that breakpoint. The convenience variable $_and the default examining-address for the x command are set to the address of the last breakpoint listed (see Section 8.5 [Examining memory], page 69). info break displays a count of the number of times the breakpoint has been hit. This is especially useful in conjunction with the ignore command. You can ignore a large number of breakpoint hits, look at the breakpoint info to see how many times the breakpoint was hit, and then run again, ignoring one less than that number. This will get you quickly to the last hit of that breakpoint.

[No value for “GDBN”] allows you to set any number of breakpoints at the same place in your program. There is nothing silly or meaningless about this. When the breakpoints are conditional, this is even useful (see Section 5.1.6 [Break conditions], page 42).

[No value for “GDBN”] itself sometimes sets breakpoints in your program for special purposes, such as proper handling of longjmp (in C programs). These internal breakpoints are assigned negative numbers, starting with -1; ‘info breakpoints’ does not display them.

You can see these breakpoints with the [No value for “GDBN”] maintenance command

‘maint info breakpoints’.

maint info breakpoints

Using the same format as ‘info breakpoints’, display both the breakpoints you’ve set explicitly, and those [No value for “GDBN”] is using for internal purposes. Internal breakpoints are shown with negative breakpoint numbers. The type column identifies what kind of breakpoint is shown:

breakpoint

Normal, explicitly set breakpoint.

watchpoint

Normal, explicitly set watchpoint.

longjmp Internal breakpoint, used to handle correctly stepping through

          longjmp calls.
longjmp resume

Internal breakpoint at the target of a longjmp.

until Temporary internal breakpoint used by the [No value for “GDBN”] until command.

finish Temporary internal breakpoint used by the [No value for “GDBN”] finish command.

shlib events

Shared library events.

5.1.2 Setting watchpoints

You can use a watchpoint to stop execution whenever the value of an expression changes, without having to predict a particular place where this may happen.

Depending on your system, watchpoints may be implemented in software or hardware. [No value for “GDBN”] does software watchpointing by single-stepping your program and testing the variable’s value each time, which is hundreds of times slower than normal execution. (But this may still be worth it, to catch errors where you have no clue what part of your program is the culprit.)

On some systems, such as HP-UX, Linux and some other x86-based targets, [No value for “GDBN”] includes support for hardware watchpoints, which do not slow down the running of your program.

watch expr

Set a watchpoint for an expression. [No value for “GDBN”] will break when expr is written into by the program and its value changes.

rwatch expr

Set a watchpoint that will break when watch expr is read by the program.

awatch expr

Set a watchpoint that will break when expr is either read or written into by the program.

info watchpoints

This command prints a list of watchpoints, breakpoints, and catchpoints; it is the same as info break.

[No value for “GDBN”] sets a hardware watchpoint if possible. Hardware watchpoints execute very quickly, and the debugger reports a change in value at the exact instruction where the change occurs. If [No value for “GDBN”] cannot set a hardware watchpoint, it sets a software watchpoint, which executes more slowly and reports the change in value at the next statement, not the instruction, after the change occurs.

When you issue the watch command, [No value for “GDBN”] reports

     Hardware watchpoint num: expr

if it was able to set a hardware watchpoint.

Currently, the awatch and rwatch commands can only set hardware watchpoints, because accesses to data that don’t change the value of the watched expression cannot be detected without examining every instruction as it is being executed, and [No value for “GDBN”] does not do that currently. If [No value for “GDBN”] finds that it is unable to set a hardware breakpoint with the awatch or rwatch command, it will print a message like this:

     Expression cannot be implemented with read/access watchpoint.

Sometimes, [No value for “GDBN”] cannot set a hardware watchpoint because the data type of the watched expression is wider than what a hardware watchpoint on the target machine can handle. For example, some systems can only watch regions that are up to 4 bytes wide; on such systems you cannot set hardware watchpoints for an expression that yields a double-precision floating-point number (which is typically 8 bytes wide). As a work-around, it might be possible to break the large region into a series of smaller ones and watch them with separate watchpoints.

If you set too many hardware watchpoints, [No value for “GDBN”] might be unable to insert all of them when you resume the execution of your program. Since the precise number of active watchpoints is unknown until such time as the program is about to be resumed, [No value for “GDBN”] might not be able to warn you about this when you set the watchpoints, and the warning will be printed only when the program is resumed:

     Hardware watchpoint num: Could not insert watchpoint

If this happens, delete or disable some of the watchpoints.

The SPARClite DSU will generate traps when a program accesses some data or instruction address that is assigned to the debug registers. For the data addresses, DSU facilitates the watch command. However the hardware breakpoint registers can only take two data watchpoints, and both watchpoints must be the same kind. For example, you can set two watchpoints with watch commands, two with rwatch commands, or two with awatch commands, but you cannot set one watchpoint with one command and the other with a different command. [No value for “GDBN”] will reject the command if you try to mix watchpoints. Delete or disable unused watchpoint commands before setting new ones.

If you call a function interactively using print or call, any watchpoints you have set will be inactive until [No value for “GDBN”] reaches another kind of breakpoint or the call completes.

[No value for “GDBN”] automatically deletes watchpoints that watch local (automatic) variables, or expressions that involve such variables, when they go out of scope, that is, when the execution leaves the block in which these variables were defined. In particular, when the program being debugged terminates, all local variables go out of scope, and so only watchpoints that watch global variables remain set. If you rerun the program, you will need to set all such watchpoints again. One way of doing that would be to set a code breakpoint at the entry to the main function and when it breaks, set all the watchpoints.

Warning: In multi-thread programs, watchpoints have only limited usefulness. With the current watchpoint implementation, [No value for “GDBN”] can only watch the value of an expression in a single thread. If you are confident that the expression can only change due to the current thread’s activity (and if you are also confident that no other thread can become current), then you can use watchpoints as usual. However, [No value for “GDBN”] may not notice when a non-current thread’s activity changes the expression. HP-UX Warning: In multi-thread programs, software watchpoints have only limited usefulness. If [No value for “GDBN”] creates a software watchpoint, it can only watch the value of an expression in a single thread. If you are confident that the expression can only change due to the current thread’s activity (and if you are also confident that no other thread can become current), then you can use software watchpoints as usual. However, [No value for “GDBN”] may not notice when a non-current thread’s activity changes the expression. (Hardware watchpoints, in contrast, watch an expression in all threads.)

5.1.3 Setting catchpoints

You can use catchpoints to cause the debugger to stop for certain kinds of program events, such as C++ exceptions or the loading of a shared library. Use the catch command to set a catchpoint.

catch event

Stop when event occurs. event can be any of the following:

throw The throwing of a C++ exception.

catch The catching of a C++ exception.

exec A call to exec. This is currently only available for HP-UX.

fork A call to fork. This is currently only available for HP-UX.

vfork A call to vfork. This is currently only available for HP-UX.

     load

load libname The dynamic loading of any shared library, or the loading of the library libname. This is currently only available for HP-UX.

     unload

unload libname The unloading of any dynamically loaded shared library, or the unloading of the library libname. This is currently only available for HP-UX.

tcatch event

Set a catchpoint that is enabled only for one stop. The catchpoint is automatically deleted after the first time the event is caught.

Use the info break command to list the current catchpoints.

There are currently some limitations to C++ exception handling (catch throw and catch catch) in [No value for “GDBN”]:

• If you call a function interactively, [No value for “GDBN”] normally returns control to you when the function has finished executing. If the call raises an exception, however, the call may bypass the mechanism that returns control to you and cause your program either to abort or to simply continue running until it hits a breakpoint, catches a signal that [No value for “GDBN”] is listening for, or exits. This is the case even if you set a catchpoint for the exception; catchpoints on exceptions are disabled within interactive calls.

• You cannot raise an exception interactively.

• You cannot install an exception handler interactively.

Sometimes catch is not the best way to debug exception handling: if you need to know exactly where an exception is raised, it is better to stop before the exception handler is called, since that way you can see the stack before any unwinding takes place. If you set a breakpoint in an exception handler instead, it may not be easy to find out where the exception was raised.

To stop just before an exception handler is called, you need some knowledge of the implementation. In the case of gnu C++, exceptions are raised by calling a library function named __raise_exception which has the following ANSI C interface:

         /* addr is where the exception identifier is stored.
            id is the exception identifier.  */
         void __raise_exception (void **addr, void *id);

To make the debugger catch all exceptions before any stack unwinding takes place, set a breakpoint on __raise_exception (see Section 5.1 [Breakpoints; watchpoints; and exceptions], page 33).

With a conditional breakpoint (see Section 5.1.6 [Break conditions], page 42) that depends on the value of id, you can stop your program when a specific exception is raised. You can use multiple conditional breakpoints to stop your program when any of a number of exceptions are raised.

5.1.4 Deleting breakpoints

It is often necessary to eliminate a breakpoint, watchpoint, or catchpoint once it has done its job and you no longer want your program to stop there. This is called deleting the breakpoint. A breakpoint that has been deleted no longer exists; it is forgotten.

With the clear command you can delete breakpoints according to where they are in your program. With the delete command you can delete individual breakpoints, watchpoints, or catchpoints by specifying their breakpoint numbers.

It is not necessary to delete a breakpoint to proceed past it. [No value for “GDBN”] automatically ignores breakpoints on the first instruction to be executed when you continue execution without changing the execution address. clear Delete any breakpoints at the next instruction to be executed in the selected

stack frame (see Section 6.3 [Selecting a frame], page 55). When the innermost frame is selected, this is a good way to delete a breakpoint where your program just stopped.

clear function clear filename:function

Delete any breakpoints set at entry to the function function.

clear linenum clear filename:linenum

Delete any breakpoints set at or within the code of the specified line.

delete [breakpoints] [range...]

Delete the breakpoints, watchpoints, or catchpoints of the breakpoint ranges specified as arguments. If no argument is specified, delete all breakpoints ([No value for “GDBN”] asks confirmation, unless you have set confirm off). You can abbreviate this command as d.

5.1.5 Disabling breakpoints

Rather than deleting a breakpoint, watchpoint, or catchpoint, you might prefer to disable it. This makes the breakpoint inoperative as if it had been deleted, but remembers the information on the breakpoint so that you can enable it again later.

You disable and enable breakpoints, watchpoints, and catchpoints with the enable and disable commands, optionally specifying one or more breakpoint numbers as arguments. Use info break or info watch to print a list of breakpoints, watchpoints, and catchpoints if you do not know which numbers to use.

A breakpoint, watchpoint, or catchpoint can have any of four different states of enablement:

• Enabled. The breakpoint stops your program. A breakpoint set with the break command starts out in this state.

• Disabled. The breakpoint has no effect on your program.

• Enabled once. The breakpoint stops your program, but then becomes disabled.

• Enabled for deletion. The breakpoint stops your program, but immediately after it does so it is deleted permanently. A breakpoint set with the tbreak command starts out in this state.

You can use the following commands to enable or disable breakpoints, watchpoints, and catchpoints:

disable [breakpoints] [range...]

Disable the specified breakpoints—or all breakpoints, if none are listed. A disabled breakpoint has no effect but is not forgotten. All options such as ignore-counts, conditions and commands are remembered in case the breakpoint is enabled again later. You may abbreviate disable as dis.

enable [breakpoints] [range...]

Enable the specified breakpoints (or all defined breakpoints). They become effective once again in stopping your program.

enable [breakpoints] once range...

Enable the specified breakpoints temporarily. [No value for “GDBN”] disables any of these breakpoints immediately after stopping your program.

enable [breakpoints] delete range...

Enable the specified breakpoints to work once, then die. [No value for “GDBN”] deletes any of these breakpoints as soon as your program stops there.

Except for a breakpoint set with tbreak (see Section 5.1.1 [Setting breakpoints], page 34), breakpoints that you set are initially enabled; subsequently, they become disabled or enabled only when you use one of the commands above. (The command until can set and delete a breakpoint of its own, but it does not change the state of your other breakpoints; see Section 5.2 [Continuing and stepping], page 46.)

5.1.6 Break conditions

The simplest sort of breakpoint breaks every time your program reaches a specified place. You can also specify a condition for a breakpoint. A condition is just a Boolean expression in your programming language (see Section 8.1 [Expressions], page 65). A breakpoint with a condition evaluates the expression each time your program reaches it, and your program stops only if the condition is true.

This is the converse of using assertions for program validation; in that situation, you want to stop when the assertion is violated—that is, when the condition is false. In C, if you want to test an assertion expressed by the condition assert, you should set the condition ‘! assert’ on the appropriate breakpoint.

Conditions are also accepted for watchpoints; you may not need them, since a watchpoint is inspecting the value of an expression anyhow—but it might be simpler, say, to just set a watchpoint on a variable name, and specify a condition that tests whether the new value is an interesting one.

Break conditions can have side effects, and may even call functions in your program. This can be useful, for example, to activate functions that log program progress, or to use your own print functions to format special data structures. The effects are completely predictable unless there is another enabled breakpoint at the same address. (In that case, [No value for “GDBN”] might see the other breakpoint first and stop your program without checking the condition of this one.) Note that breakpoint commands are usually more convenient and flexible than break conditions for the purpose of performing side effects when a breakpoint is reached (see Section 5.1.7 [Breakpoint command lists], page 43).

Break conditions can be specified when a breakpoint is set, by using ‘if’ in the arguments to the break command. See Section 5.1.1 [Setting breakpoints], page 34. They can also be changed at any time with the condition command.

You can also use the if keyword with the watch command. The catch command does not recognize the if keyword; condition is the only way to impose a further condition on a catchpoint.

condition bnum expression Specify expression as the break condition for breakpoint, watchpoint, or catchpoint number bnum. After you set a condition, breakpoint bnum stops your program only if the value of expression is true (nonzero, in C). When you use condition, [No value for “GDBN”] checks expression immediately for syntactic correctness, and to determine whether symbols in it have referents in the context of your breakpoint. If expression uses symbols not referenced in the context of the breakpoint, [No value for “GDBN”] prints an error message:

     No symbol "foo" in current context.

[No value for “GDBN”] does not actually evaluate expression at the time the condition command (or a command that sets a breakpoint with a condition, like break if ...) is given, however. See Section 8.1 [Expressions], page 65.

condition bnum

Remove the condition from breakpoint number bnum. It becomes an ordinary unconditional breakpoint.

A special case of a breakpoint condition is to stop only when the breakpoint has been reached a certain number of times. This is so useful that there is a special way to do it, using the ignore count of the breakpoint. Every breakpoint has an ignore count, which is an integer. Most of the time, the ignore count is zero, and therefore has no effect. But if your program reaches a breakpoint whose ignore count is positive, then instead of stopping, it just decrements the ignore count by one and continues. As a result, if the ignore count value is n, the breakpoint does not stop the next n times your program reaches it.

ignore bnum count Set the ignore count of breakpoint number bnum to count. The next count times the breakpoint is reached, your program’s execution does not stop; other than to decrement the ignore count, [No value for “GDBN”] takes no action.

To make the breakpoint stop the next time it is reached, specify a count of zero.

When you use continue to resume execution of your program from a breakpoint, you can specify an ignore count directly as an argument to continue, rather than using ignore. See Section 5.2 [Continuing and stepping], page 46.

If a breakpoint has a positive ignore count and a condition, the condition is not checked. Once the ignore count reaches zero, [No value for “GDBN”] resumes checking the condition.

You could achieve the effect of the ignore count with a condition such as ‘$foo-- <= 0’ using a debugger convenience variable that is decremented each time. See Section 8.9 [Convenience variables], page 78.

Ignore counts apply to breakpoints, watchpoints, and catchpoints.

5.1.7 Breakpoint command lists

You can give any breakpoint (or watchpoint or catchpoint) a series of commands to execute when your program stops due to that breakpoint. For example, you might want to print the values of certain expressions, or enable other breakpoints.

commands [bnum]

... command-list ... end Specify a list of commands for breakpoint number bnum. The commands them-

selves appear on the following lines. Type a line containing just end to terminate the commands. To remove all commands from a breakpoint, type commands and follow it immediately with end; that is, give no commands. With no bnum argument, commands refers to the last breakpoint, watchpoint, or catchpoint set (not to the breakpoint most recently encountered).

Pressing ^{⟨}\underline{\overline{RET}}^{⟩} as a means of repeating the last [No value for “GDBN”] command is disabled within a command-list.

You can use breakpoint commands to start your program up again. Simply use the continue command, or step, or any other command that resumes execution.

Any other commands in the command list, after a command that resumes execution, are ignored. This is because any time you resume execution (even with a simple next or step), you may encounter another breakpoint—which could have its own command list, leading to ambiguities about which list to execute.

If the first command you specify in a command list is silent, the usual message about stopping at a breakpoint is not printed. This may be desirable for breakpoints that are to print a specific message and then continue. If none of the remaining commands print anything, you see no sign that the breakpoint was reached. silent is meaningful only at the beginning of a breakpoint command list.

The commands echo, output, and printf allow you to print precisely controlled output, and are often useful in silent breakpoints. See Section 17.4 [Commands for controlled output], page 180.

For example, here is how you could use breakpoint commands to print the value of x at entry to foo whenever x is positive.

     break foo if x>0
     commands
     silent
     printf "x is %d\n",x
     cont
     end

One application for breakpoint commands is to compensate for one bug so you can test for another. Put a breakpoint just after the erroneous line of code, give it a condition to detect the case in which something erroneous has been done, and give it commands to assign correct values to any variables that need them. End with the continue command so that your program does not stop, and start with the silent command so that no output is produced. Here is an example:

     break 403
     commands
     silent
     set x = y + 4
     cont
     end

5.1.8 Breakpoint menus

Some programming languages (notably C++) permit a single function name to be defined several times, for application in different contexts. This is called overloading. When a function name is overloaded, ‘break function’ is not enough to tell [No value for “GDBN”] where you want a breakpoint. If you realize this is a problem, you can use something like ‘break function(types)’ to specify which particular version of the function you want. Otherwise, [No value for “GDBN”] offers you a menu of numbered choices for different possible breakpoints, and waits for your selection with the prompt ‘>’. The first two options are always ‘[0] cancel’ and ‘[1] all’. Typing 1 sets a breakpoint at each definition of function, and typing 0 aborts the break command without setting any new breakpoints.

For example, the following session excerpt shows an attempt to set a breakpoint at the overloaded symbol String::after. We choose three particular definitions of that function name:

([No value for ``GDBP'']) b String::after
[0] cancel
[1] all
[2] file:String.cc; line number:867
[3] file:String.cc; line number:860
[4] file:String.cc; line number:875
[5] file:String.cc; line number:853
[6] file:String.cc; line number:846
[7] file:String.cc; line number:735
> 2 4 6
Breakpoint 1 at 0xb26c: file String.cc, line 867.
Breakpoint 2 at 0xb344: file String.cc, line 875.
Breakpoint 3 at 0xafcc: file String.cc, line 846.
Multiple breakpoints were set.
Use the "delete" command to delete unwanted
 breakpoints.
([No value for ``GDBP''])

5.1.9 “Cannot insert breakpoints”

Under some operating systems, breakpoints cannot be used in a program if any other process is running that program. In this situation, attempting to run or continue a program with a breakpoint causes [No value for “GDBN”] to print an error message:

Cannot insert breakpoints.
The same program may be running in another process.

When this happens, you have three ways to proceed:

1. Remove or disable the breakpoints, then continue.

2. Suspend [No value for “GDBN”], and copy the file containing your program to a new name. Resume [No value for “GDBN”] and use the exec-file command to specify that [No value for “GDBN”] should run your program under that name. Then start your program again.

3. Relink your program so that the text segment is nonsharable, using the linker option ‘-N’. The operating system limitation may not apply to nonsharable executables.

A similar message can be printed if you request too many active hardware-assisted breakpoints and watchpoints:

     Stopped; cannot insert breakpoints.
     You may have requested too many hardware breakpoints and watchpoints.

This message is printed when you attempt to resume the program, since only then [No value for “GDBN”] knows exactly how many hardware breakpoints and watchpoints it needs to insert.

When this message is printed, you need to disable or remove some of the hardwareassisted breakpoints and watchpoints, and then continue.

5.2 Continuing and stepping

Continuing means resuming program execution until your program completes normally. In contrast, stepping means executing just one more “step” of your program, where “step” may mean either one line of source code, or one machine instruction (depending on what particular command you use). Either when continuing or when stepping, your program may stop even sooner, due to a breakpoint or a signal. (If it stops due to a signal, you may want to use handle, or use ‘signal 0’ to resume execution. See Section 5.3 [Signals], page 49.)

continue [ignore-count] c [ignore-count] fg [ignore-count]

Resume program execution, at the address where your program last stopped; any breakpoints set at that address are bypassed. The optional argument ignore-count allows you to specify a further number of times to ignore a breakpoint at this location; its effect is like that of ignore (see Section 5.1.6 [Break conditions], page 42). The argument ignore-count is meaningful only when your program stopped due to a breakpoint. At other times, the argument to continue is ignored. The synonyms c and fg (for foreground, as the debugged program is deemed to be the foreground program) are provided purely for convenience, and have exactly the same behavior as continue.

To resume execution at a different place, you can use return (see Section 12.4 [Returning from a function], page 121) to go back to the calling function; or jump (see Section 12.2 [Continuing at a different address], page 120) to go to an arbitrary location in your program.

A typical technique for using stepping is to set a breakpoint (see Section 5.1 [Breakpoints; watchpoints; and catchpoints], page 33) at the beginning of the function or the section of your program where a problem is believed to lie, run your program until it stops at that breakpoint, and then step through the suspect area, examining the variables that are interesting, until you see the problem happen.

step Continue running your program until control reaches a different source line, then stop it and return control to [No value for “GDBN”]. This command is abbreviated s.

Warning: If you use the step command while control is within a function that was compiled without debugging information, execution proceeds until control reaches a function that does have debugging information. Likewise, it will not step into a function which is compiled without debugging information. To step through functions without debugging information, use the stepi command, described below.

The step command only stops at the first instruction of a source line. This prevents the multiple stops that could otherwise occur in switch statements, for loops, etc. step continues to stop if a function that has debugging information is called within the line. In other words, step steps inside any functions called within the line. Also, the step command only enters a function if there is line number information for the function. Otherwise it acts like the next command. This avoids problems when using cc -gl on MIPS machines. Previously, step entered subroutines if there was any debugging information about the routine.

step count Continue running as in step, but do so count times. If a breakpoint is reached, or a signal not related to stepping occurs before count steps, stepping stops right away.

next [count] Continue to the next source line in the current (innermost) stack frame. This is similar to step, but function calls that appear within the line of code are executed without stopping. Execution stops when control reaches a different line of code at the original stack level that was executing when you gave the next command. This command is abbreviated n. An argument count is a repeat count, as for step. The next command only stops at the first instruction of a source line. This prevents multiple stops that could otherwise occur in switch statements, for loops, etc.

set step-mode
set step-mode on

The set step-mode on command causes the step command to stop at the first instruction of a function which contains no debug line information rather than stepping over it. This is useful in cases where you may be interested in inspecting the machine instructions of a function which has no symbolic info and do not want [No value for “GDBN”] to automatically skip over this function.

set step-mode off

Causes the step command to step over any functions which contains no debug information. This is the default.

finish Continue running until just after function in the selected stack frame returns. Print the returned value (if any).

Contrast this with the return command (see Section 12.4 [Returning from a function], page 121).

until

u Continue running until a source line past the current line, in the current stack frame, is reached. This command is used to avoid single stepping through a loop more than once. It is like the next command, except that when until encounters a jump, it automatically continues execution until the program counter is greater than the address of the jump. This means that when you reach the end of a loop after single stepping though it, until makes your program continue execution until it exits the loop. In contrast, a next command at the end of a loop simply steps back to the beginning of the loop, which forces you to step through the next iteration. until always stops your program if it attempts to exit the current stack frame. until may produce somewhat counterintuitive results if the order of machine code does not match the order of the source lines. For example, in the following excerpt from a debugging session, the f (frame) command shows that execution is stopped at line 206; yet when we use until, we get to line 195:

     ([No value for ``GDBP'']) f
     #0  main (argc=4, argv=0xf7fffae8) at m4.c:206
     206                 expand_input();
     ([No value for ``GDBP'']) until
     195             for ( ; argc > 0; NEXTARG) {

This happened because, for execution efficiency, the compiler had generated code for the loop closure test at the end, rather than the start, of the loop— even though the test in a C for-loop is written before the body of the loop. The until command appeared to step back to the beginning of the loop when it advanced to this expression; however, it has not really gone to an earlier statement—not in terms of the actual machine code. until with no argument works by means of single instruction stepping, and hence is slower than until with an argument.

until location u location Continue running your program until either the specified location is reached,

or the current stack frame returns. location is any of the forms of argument acceptable to break (see Section 5.1.1 [Setting breakpoints], page 34). This form of the command uses breakpoints, and hence is quicker than until without an argument.

stepi
stepi arg

si Execute one machine instruction, then stop and return to the debugger. It is often useful to do ‘display/i $pc’ when stepping by machine instructions. This makes [No value for “GDBN”] automatically display the next instruction to be executed, each time your program stops. See Section 8.6 [Automatic display], page 70. An argument is a repeat count, as in step.

nexti
nexti arg

ni Execute one machine instruction, but if it is a function call, proceed until the function returns. An argument is a repeat count, as in next.

5.3 Signals

A signal is an asynchronous event that can happen in a program. The operating system defines the possible kinds of signals, and gives each kind a name and a number. For example, in Unix SIGINT is the signal a program gets when you type an interrupt character (often Cc); SIGSEGV is the signal a program gets from referencing a place in memory far away from all the areas in use; SIGALRM occurs when the alarm clock timer goes off (which happens only if your program has requested an alarm).

Some signals, including SIGALRM, are a normal part of the functioning of your program. Others, such as SIGSEGV, indicate errors; these signals are fatal (they kill your program immediately) if the program has not specified in advance some other way to handle the signal. SIGINT does not indicate an error in your program, but it is normally fatal so it can carry out the purpose of the interrupt: to kill the program.

[No value for “GDBN”] has the ability to detect any occurrence of a signal in your program. You can tell [No value for “GDBN”] in advance what to do for each kind of signal.

Normally, [No value for “GDBN”] is set up to let the non-erroneous signals like SIGALRM be silently passed to your program (so as not to interfere with their role in the program’s functioning) but to stop your program immediately whenever an error signal happens. You can change these settings with the handle command.

info signals
info handle

Print a table of all the kinds of signals and how [No value for “GDBN”] has been told to handle each one. You can use this to see the signal numbers of all the defined types of signals.

          info handle is an alias for info signals.

handle signal keywords...

Change the way [No value for “GDBN”] handles signal signal. signal can be the number of a signal or its name (with or without the ‘SIG’ at the beginning); a list of signal numbers of the form ‘low-high’; or the word ‘all’, meaning all the known signals. The keywords say what change to make.

The keywords allowed by the handle command can be abbreviated. Their full names are:

nostop [No value for “GDBN”] should not stop your program when this signal happens. It may still print a message telling you that the signal has come in.

stop [No value for “GDBN”] should stop your program when this signal happens. This implies the print keyword as well. print [No value for “GDBN”] should print a message when this signal happens.

noprint [No value for “GDBN”] should not mention the occurrence of the signal at all. This implies the nostop keyword as well.

pass

noignore [No value for “GDBN”] should allow your program to see this signal; your program can handle the signal, or else it may terminate if the signal is fatal and not handled. pass and noignore are synonyms.

nopass

ignore [No value for “GDBN”] should not allow your program to see this signal. nopass and ignore are synonyms.

When a signal stops your program, the signal is not visible to the program until you continue. Your program sees the signal then, if pass is in effect for the signal in question at that time. In other words, after [No value for “GDBN”] reports a signal, you can use the handle command with pass or nopass to control whether your program sees that signal when you continue.

The default is set to nostop, noprint, pass for non-erroneous signals such as SIGALRM, SIGWINCH and SIGCHLD, and to stop, print, pass for the erroneous signals.

You can also use the signal command to prevent your program from seeing a signal, or cause it to see a signal it normally would not see, or to give it any signal at any time. For example, if your program stopped due to some sort of memory reference error, you might store correct values into the erroneous variables and continue, hoping to see more execution; but your program would probably terminate immediately as a result of the fatal signal once it saw the signal. To prevent this, you can continue with ‘signal 0’. See Section 12.3 [Giving your program a signal], page 121.

5.4 Stopping and starting multi-thread programs

When your program has multiple threads (see Section 4.9 [Debugging programs with multiple threads], page 28), you can choose whether to set breakpoints on all threads, or on a particular thread.

break linespec thread threadno

break linespec thread threadno if ...

linespec specifies source lines; there are several ways of writing them, but the effect is always to specify some source line. Use the qualifier ‘thread threadno’ with a breakpoint command to specify that you only want [No value for “GDBN”] to stop the program when a particular thread reaches this breakpoint. threadno is one of the numeric thread identifiers assigned by [No value for “GDBN”], shown in the first column of the ‘info threads’ display. If you do not specify ‘thread threadno’ when you set a breakpoint, the breakpoint applies to all threads of your program. You can use the thread qualifier on conditional breakpoints as well; in this case, place ‘thread threadno’ before the breakpoint condition, like this:

               ([No value for ``GDBP'']) break frik.c:13 thread 28 if bartab > lim

Whenever your program stops under [No value for “GDBN”] for any reason, all threads of execution stop, not just the current thread. This allows you to examine the overall state of the program, including switching between threads, without worrying that things may change underfoot.

Conversely, whenever you restart the program, all threads start executing. This is true even when single-stepping with commands like step or next.

In particular, [No value for “GDBN”] cannot single-step all threads in lockstep. Since thread scheduling is up to your debugging target’s operating system (not controlled by [No value for “GDBN”]), other threads may execute more than one statement while the current thread completes a single step. Moreover, in general other threads stop in the middle of a statement, rather than at a clean statement boundary, when the program stops.

You might even find your program stopped in another thread after continuing or even single-stepping. This happens whenever some other thread runs into a breakpoint, a signal, or an exception before the first thread completes whatever you requested.

On some OSes, you can lock the OS scheduler and thus allow only a single thread to run.

set scheduler-locking mode

Set the scheduler locking mode. If it is off, then there is no locking and any thread may run at any time. If on, then only the current thread may run when the inferior is resumed. The step mode optimizes for single-stepping. It stops other threads from “seizing the prompt” by preempting the current thread while you are stepping. Other threads will only rarely (or never) get a chance to run when you step. They are more likely to run when you ‘next’ over a function call, and they are completely free to run when you use commands like ‘continue’, ‘until’, or ‘finish’. However, unless another thread hits a breakpoint during its timeslice, they will never steal the [No value for “GDBN”] prompt away from the thread that you are debugging.

show scheduler-locking

Display the current scheduler locking mode.

52 Debugging with [No value for “GDBN”]

6 Examining the Stack

When your program has stopped, the first thing you need to know is where it stopped and how it got there.

Each time your program performs a function call, information about the call is generated. That information includes the location of the call in your program, the arguments of the call, and the local variables of the function being called. The information is saved in a block of data called a stack frame. The stack frames are allocated in a region of memory called the call stack.

When your program stops, the [No value for “GDBN”] commands for examining the stack allow you to see all of this information.

One of the stack frames is selected by [No value for “GDBN”] and many [No value for “GDBN”] commands refer implicitly to the selected frame. In particular, whenever you ask [No value for “GDBN”] for the value of a variable in your program, the value is found in the selected frame. There are special [No value for “GDBN”] commands to select whichever frame you are interested in. See Section 6.3 [Selecting a frame], page 55.

When your program stops, [No value for “GDBN”] automatically selects the currently executing frame and describes it briefly, similar to the frame command (see Section 6.4 [Information about a frame], page 56).

6.1 Stack frames

The call stack is divided up into contiguous pieces called stack frames, or frames for short; each frame is the data associated with one call to one function. The frame contains the arguments given to the function, the function’s local variables, and the address at which the function is executing.

When your program is started, the stack has only one frame, that of the function main. This is called the initial frame or the outermost frame. Each time a function is called, a new frame is made. Each time a function returns, the frame for that function invocation is eliminated. If a function is recursive, there can be many frames for the same function. The frame for the function in which execution is actually occurring is called the innermost frame. This is the most recently created of all the stack frames that still exist.

Inside your program, stack frames are identified by their addresses. A stack frame consists of many bytes, each of which has its own address; each kind of computer has a convention for choosing one byte whose address serves as the address of the frame. Usually this address is kept in a register called the frame pointer register while execution is going on in that frame.

[No value for “GDBN”] assigns numbers to all existing stack frames, starting with zero for the innermost frame, one for the frame that called it, and so on upward. These numbers do not really exist in your program; they are assigned by [No value for “GDBN”] to give you a way of designating stack frames in [No value for “GDBN”] commands.

Some compilers provide a way to compile functions so that they operate without stack frames. (For example, the [No value for “GCC”] option

     `-fomit-frame-pointer'

generates functions without a frame.) This is occasionally done with heavily used library functions to save the frame setup time. [No value for “GDBN”] has limited facilities for dealing with these function invocations. If the innermost function invocation has no stack frame, [No value for “GDBN”] nevertheless regards it as though it had a separate frame, which is numbered zero as usual, allowing correct tracing of the function call chain. However, [No value for “GDBN”] has no provision for frameless functions elsewhere in the stack.

frame args

The frame command allows you to move from one stack frame to another, and to print the stack frame you select. args may be either the address of the frame or the stack frame number. Without an argument, frame prints the current stack frame.

select-frame

The select-frame command allows you to move from one stack frame to another without printing the frame. This is the silent version of frame.

6.2 Backtraces

A backtrace is a summary of how your program got where it is. It shows one line per frame, for many frames, starting with the currently executing frame (frame zero), followed by its caller (frame one), and on up the stack.

backtrace

bt Print a backtrace of the entire stack: one line per frame for all frames in the stack. You can stop the backtrace at any time by typing the system interrupt character, normally C-c.

backtrace n

bt n Similar, but print only the innermost n frames.

backtrace -n

bt -n Similar, but print only the outermost n frames.

The names where and info stack (abbreviated info s) are additional aliases for

backtrace.

Each line in the backtrace shows the frame number and the function name. The program counter value is also shown—unless you use set print address off. The backtrace also shows the source file name and line number, as well as the arguments to the function. The program counter value is omitted if it is at the beginning of the code for that line number.

Here is an example of a backtrace. It was made with the command ‘bt 3’, so it shows the innermost three frames.

     #0  m4_traceon (obs=0x24eb0, argc=1, argv=0x2b8c8)
         at builtin.c:993
     #1  0x6e38 in expand_macro (sym=0x2b600) at macro.c:242
     #2  0x6840 in expand_token (obs=0x0, t=177664, td=0xf7fffb08)
         at macro.c:71
     (More stack frames follow...)

The display for frame zero does not begin with a program counter value, indicating that your program has stopped at the beginning of the code for line 993 of builtin.c.

6.3 Selecting a frame

Most commands for examining the stack and other data in your program work on whichever stack frame is selected at the moment. Here are the commands for selecting a stack frame; all of them finish by printing a brief description of the stack frame just selected.

frame n

f n Select frame number n. Recall that frame zero is the innermost (currently executing) frame, frame one is the frame that called the innermost one, and so on. The highest-numbered frame is the one for main.

frame addr

f addr Select the frame at address addr. This is useful mainly if the chaining of stack frames has been damaged by a bug, making it impossible for [No value for “GDBN”] to assign numbers properly to all frames. In addition, this can be useful when your program has multiple stacks and switches between them. On the SPARC architecture, frame needs two addresses to select an arbitrary frame: a frame pointer and a stack pointer. On the MIPS and Alpha architecture, it needs two addresses: a stack pointer and a program counter. On the 29k architecture, it needs three addresses: a register stack pointer, a program counter, and a memory stack pointer.

up n Move n frames up the stack. For positive numbers n, this advances toward the outermost frame, to higher frame numbers, to frames that have existed longer. n defaults to one.

down n Move n frames down the stack. For positive numbers n, this advances toward the innermost frame, to lower frame numbers, to frames that were created more recently. n defaults to one. You may abbreviate down as do.

All of these commands end by printing two lines of output describing the frame. The first line shows the frame number, the function name, the arguments, and the source file and line number of execution in that frame. The second line shows the text of that source line.

For example:

     ([No value for ``GDBP'']) up
     #1  0x22f0 in main (argc=1, argv=0xf7fffbf4, env=0xf7fffbfc)
         at env.c:10
     10              read_input_file (argv[i]);

After such a printout, the list command with no arguments prints ten lines centered on the point of execution in the frame. See Section 7.1 [Printing source lines], page 59.

up-silently n
down-silently n

These two commands are variants of up and down, respectively; they differ in that they do their work silently, without causing display of the new frame. They are intended primarily for use in [No value for “GDBN”] command scripts, where the output might be unnecessary and distracting.

6.4 Information about a frame

There are several other commands to print information about the selected stack frame.

frame

f When used without any argument, this command does not change which frame is selected, but prints a brief description of the currently selected stack frame. It can be abbreviated f. With an argument, this command is used to select a stack frame. See Section 6.3 [Selecting a frame], page 55.

info frame

info f This command prints a verbose description of the selected stack frame, includ-

ing:

• the address of the frame

• the address of the next frame down (called by this frame)

• the address of the next frame up (caller of this frame)

• the language in which the source code corresponding to this frame is written • the address of the frame’s arguments

• the address of the frame’s local variables

• the program counter saved in it (the address of execution in the caller frame)

• which registers were saved in the frame

The verbose description is useful when something has gone wrong that has made the stack format fail to fit the usual conventions.

info frame addr
info f addr

Print a verbose description of the frame at address addr, without selecting that frame. The selected frame remains unchanged by this command. This requires the same kind of address (more than one for some architectures) that you specify in the frame command. See Section 6.3 [Selecting a frame], page 55.

info args Print the arguments of the selected frame, each on a separate line.

info locals

Print the local variables of the selected frame, each on a separate line. These are all variables (declared either static or automatic) accessible at the point of execution of the selected frame.

info catch

Print a list of all the exception handlers that are active in the current stack frame at the current point of execution. To see other exception handlers, visit Chapter 6: Examining the Stack 57

the associated frame (using the up, down, or frame commands); then type info catch. See Section 5.1.3 [Setting catchpoints], page 39.

58 Debugging with [No value for “GDBN”]

7 Examining Source Files

[No value for “GDBN”] can print parts of your program’s source, since the debugging information recorded in the program tells [No value for “GDBN”] what source files were used to build it. When your program stops, [No value for “GDBN”] spontaneously prints the line where it stopped. Likewise, when you select a stack frame (see Section 6.3 [Selecting a frame], page 55), [No value for “GDBN”] prints the line where execution in that frame has stopped. You can print other portions of source files by explicit command.

If you use [No value for “GDBN”] through its gnu Emacs interface, you may prefer to use Emacs facilities to view source; see Chapter 19 [Using [No value for “GDBN”] under gnu Emacs], page 187.

7.1 Printing source lines

To print lines from a source file, use the list command (abbreviated l). By default, ten lines are printed. There are several ways to specify what part of the file you want to print.

Here are the forms of the list command most commonly used:

list linenum Print lines centered around line number linenum in the current source file.

list function Print lines centered around the beginning of function function.

list Print more lines. If the last lines printed were printed with a list command, this prints lines following the last lines printed; however, if the last line printed was a solitary line printed as part of displaying a stack frame (see Chapter 6 [Examining the Stack], page 53), this prints lines centered around that line.

list - Print lines just before the lines last printed.

By default, [No value for “GDBN”] prints ten source lines with any of these forms of the list command. You can change this using set listsize:

set listsize count

Make the list command display count source lines (unless the list argument explicitly specifies some other number).

show listsize

Display the number of lines that list prints.

Repeating a list command with ^{⟨}\underline{\overline{RET}}^{⟩} discards the argument, so it is equivalent to typing just list. This is more useful than listing the same lines again. An exception is made for an argument of ‘-’; that argument is preserved in repetition so that each repetition moves up in the source file.

In general, the list command expects you to supply zero, one or two linespecs. Linespecs specify source lines; there are several ways of writing them, but the effect is always to specify some source line. Here is a complete description of the possible arguments for list:

list linespec Print lines centered around the line specified by linespec. list first,last

Print lines from first to last. Both arguments are linespecs.

list ,last Print lines ending with last.

list first,

Print lines starting with first.

list + Print lines just after the lines last printed.

list - Print lines just before the lines last printed.

list As described in the preceding table.

Here are the ways of specifying a single source line—all the kinds of linespec.

number Specifies line number of the current source file. When a list command has two linespecs, this refers to the same source file as the first linespec.

+offset Specifies the line offset lines after the last line printed. When used as the second linespec in a list command that has two, this specifies the line offset lines down from the first linespec.

-offset Specifies the line offset lines before the last line printed.

filename:number Specifies line number in the source file filename.

function Specifies the line that begins the body of the function function. For example: in C, this is the line with the open brace.

filename:function Specifies the line of the open-brace that begins the body of the function function in the file filename. You only need the file name with a function name to avoid ambiguity when there are identically named functions in different source files.

*address Specifies the line containing the program address address. address may be any expression.

7.2 Searching source files

There are two commands for searching through the current source file for a regular expression.

forward-search regexp
search regexp

The command ‘forward-search regexp’ checks each line, starting with the one following the last line listed, for a match for regexp. It lists the line that is found. You can use the synonym ‘search regexp’ or abbreviate the command name as

          fo.
reverse-search regexp

The command ‘reverse-search regexp’ checks each line, starting with the one before the last line listed and going backward, for a match for regexp. It lists the line that is found. You can abbreviate this command as rev.

7.3 Specifying source directories

Executable programs sometimes do not record the directories of the source files from which they were compiled, just the names. Even when they do, the directories could be moved between the compilation and your debugging session. [No value for “GDBN”] has a list of directories to search for source files; this is called the source path. Each time [No value for “GDBN”] wants a source file, it tries all the directories in the list, in the order they are present in the list, until it finds a file with the desired name. Note that the executable search path is not used for this purpose. Neither is the current working directory, unless it happens to be in the source path.

If [No value for “GDBN”] cannot find a source file in the source path, and the object program records a directory, [No value for “GDBN”] tries that directory too. If the source path is empty, and there is no record of the compilation directory, [No value for “GDBN”] looks in the current directory as a last resort.

Whenever you reset or rearrange the source path, [No value for “GDBN”] clears out any information it has cached about where source files are found and where each line is in the file.

When you start [No value for “GDBN”], its source path includes only ‘cdir’ and ‘cwd’, in that order. To add other directories, use the directory command.

directory dirname ...

dir dirname ...

Add directory dirname to the front of the source path. Several directory names may be given to this command, separated by ‘:’ (‘;’ on MS-DOS and MS-Windows, where ‘:’ usually appears as part of absolute file names) or whitespace. You may specify a directory that is already in the source path; this moves it forward, so [No value for “GDBN”] searches it sooner. You can use the string ‘$cdir’ to refer to the compilation directory (if one is recorded), and ‘$cwd’ to refer to the current working directory. ‘$cwd’ is not the same as ‘.’—the former tracks the current working directory as it changes during your [No value for “GDBN”] session, while the latter is immediately expanded to the current directory at the time you add an entry to the source path.

directory

Reset the source path to empty again. This requires confirmation.

show directories

Print the source path: show which directories it contains.

If your source path is cluttered with directories that are no longer of interest, [No value for “GDBN”] may sometimes cause confusion by finding the wrong versions of source. You can correct the situation as follows:

1. Use directory with no argument to reset the source path to empty.

2. Use directory with suitable arguments to reinstall the directories you want in the source path. You can add all the directories in one command.

7.4 Source and machine code

You can use the command info line to map source lines to program addresses (and vice versa), and the command disassemble to display a range of addresses as machine instructions. When run under gnu Emacs mode, the info line command causes the arrow to point to the line specified. Also, info line prints addresses in symbolic form as well as hex.

info line linespec

Print the starting and ending addresses of the compiled code for source line linespec. You can specify source lines in any of the ways understood by the list command (see Section 7.1 [Printing source lines], page 59).

For example, we can use info line to discover the location of the object code for the first line of function m4_changequote:

     ([No value for ``GDBP'']) info line m4_changequote
     Line 895 of "builtin.c" starts at pc 0x634c and ends at 0x6350.

We can also inquire (using *addr as the form for linespec) what source line covers a particular address:

     ([No value for ``GDBP'']) info line *0x63ff
     Line 926 of "builtin.c" starts at pc 0x63e4 and ends at 0x6404.

After info line, the default address for the x command is changed to the starting address of the line, so that ‘x/i’ is sufficient to begin examining the machine code (see Section 8.5 [Examining memory], page 69). Also, this address is saved as the value of the convenience variable $_(see Section 8.9 [Convenience variables], page 78).

disassemble

This specialized command dumps a range of memory as machine instructions. The default memory range is the function surrounding the program counter of the selected frame. A single argument to this command is a program counter value; [No value for “GDBN”] dumps the function surrounding this value. Two arguments specify a range of addresses (first inclusive, second exclusive) to dump.

The following example shows the disassembly of a range of addresses of HP PA-RISC 2.0 code:

     ([No value for ``GDBP'']) disas 0x32c4 0x32e4
     Dump of assembler code from 0x32c4 to 0x32e4:
     0x32c4 <main+204>:      addil 0,dp
     0x32c8 <main+208>:      ldw 0x22c(sr0,r1),r26
     0x32cc <main+212>:      ldil 0x3000,r31
     0x32d0 <main+216>:      ble 0x3f8(sr4,r31)
     0x32d4 <main+220>:      ldo 0(r31),rp
     0x32d8 <main+224>:      addil -0x800,dp
     0x32dc <main+228>:      ldo 0x588(r1),r26
     0x32e0 <main+232>:      ldil 0x3000,r31
     End of assembler dump.

Some architectures have more than one commonly-used set of instruction mnemonics or other syntax.

set disassembly-flavor instruction-set

Select the instruction set to use when disassembling the program via the

disassemble or x/i commands.

Currently this command is only defined for the Intel x86 family. You can set instruction-set to either intel or att. The default is att, the AT&T flavor used by default by Unix assemblers for x86-based targets.

64 Debugging with [No value for “GDBN”]

8 Examining Data

The usual way to examine data in your program is with the print command (abbreviated p), or its synonym inspect. It evaluates and prints the value of an expression of the language your program is written in (see Chapter 10 [Using [No value for “GDBN”] with Different Languages], page 93).

print expr
print /f expr

expr is an expression (in the source language). By default the value of expr is printed in a format appropriate to its data type; you can choose a different format by specifying ‘/f ’, where f is a letter specifying the format; see Section 8.4 [Output formats], page 68.

print

print /f If you omit expr, [No value for “GDBN”] displays the last value again (from the value history; see Section 8.8 [Value history], page 77). This allows you to conveniently inspect the same value in an alternative format.

A more low-level way of examining data is with the x command. It examines data in memory at a specified address and prints it in a specified format. See Section 8.5 [Examining memory], page 69.

If you are interested in information about types, or about how the fields of a struct or a class are declared, use the ptype exp command rather than print. See Chapter 11 [Examining the Symbol Table], page 115.

8.1 Expressions

print and many other [No value for “GDBN”] commands accept an expression and compute its value. Any kind of constant, variable or operator defined by the programming language you are using is valid in an expression in [No value for “GDBN”]. This includes conditional expressions, function calls, casts and string constants. It unfortunately does not include symbols defined by preprocessor #define commands.

[No value for “GDBN”] supports array constants in expressions input by the user. The syntax is {element, element…}. For example, you can use the command print {1, 2, 3} to build up an array in memory that is malloced in the target program.

Because C is so widespread, most of the expressions shown in examples in this manual are in C. See Chapter 10 [Using [No value for “GDBN”] with Different Languages], page 93, for information on how to use expressions in other languages.

In this section, we discuss operators that you can use in [No value for “GDBN”] expressions regardless of your programming language.

Casts are supported in all languages, not just in C, because it is so useful to cast a number into a pointer in order to examine a structure at that address in memory.

[No value for “GDBN”] supports these operators, in addition to those common to programming languages:

@ ‘@’ is a binary operator for treating parts of memory as arrays. See Section 8.3 [Artificial arrays], page 67, for more information.

:: ‘::’ allows you to specify a variable in terms of the file or function where it is defined. See Section 8.2 [Program variables], page 66.

{type} addr Refers to an object of type type stored at address addr in memory. addr may be any expression whose value is an integer or pointer (but parentheses are required around binary operators, just as in a cast). This construct is allowed regardless of what kind of data is normally supposed to reside at addr.

8.2 Program variables

The most common kind of expression to use is the name of a variable in your program. Variables in expressions are understood in the selected stack frame (see Section 6.3

[Selecting a frame], page 55); they must be either:

• global (or file-static)

or

• visible according to the scope rules of the programming language from the point of execution in that frame

This means that in the function

     foo (a)
          int a;
     {
       bar (a);
       {
         int b = test ();
         bar (b);
       }
     }

you can examine and use the variable a whenever your program is executing within the function foo, but you can only use or examine the variable b while your program is executing inside the block where b is declared.

There is an exception: you can refer to a variable or function whose scope is a single source file even if the current execution point is not in this file. But it is possible to have more than one such variable or function with the same name (in different source files). If that happens, referring to that name has unpredictable effects. If you wish, you can specify a static variable in a particular function or file, using the colon-colon notation:

file::variable function::variable

Here file or function is the name of the context for the static variable. In the case of file names, you can use quotes to make sure [No value for “GDBN”] parses the file name as a single word—for example, to print a global value of x defined in ‘f2.c’:

     ([No value for ``GDBP'']) p 'f2.c'::x

This use of ‘::’ is very rarely in conflict with the very similar use of the same notation in C++. [No value for “GDBN”] also supports use of the C++ scope resolution operator in [No value for “GDBN”] expressions.

Warning: Occasionally, a local variable may appear to have the wrong value at certain points in a function—just after entry to a new scope, and just before exit.

You may see this problem when you are stepping by machine instructions. This is because, on most machines, it takes more than one instruction to set up a stack frame (including local variable definitions); if you are stepping by machine instructions, variables may appear to have the wrong values until the stack frame is completely built. On exit, it usually also takes more than one machine instruction to destroy a stack frame; after you begin stepping through that group of instructions, local variable definitions may be gone.

This may also happen when the compiler does significant optimizations. To be sure of always seeing accurate values, turn off all optimization when compiling.

Another possible effect of compiler optimizations is to optimize unused variables out of existence, or assign variables to registers (as opposed to memory addresses). Depending on the support for such cases offered by the debug info format used by the compiler, [No value for “GDBN”] might not be able to display values for such local variables. If that happens, [No value for “GDBN”] will print a message like this:

     No symbol "foo" in current context.

To solve such problems, either recompile without optimizations, or use a different debug info format, if the compiler supports several such formats. For example, [No value for “NGCC”], the gnu C/C++ compiler usually supports the ‘-gstabs’ option. ‘-gstabs’ produces debug info in a format that is superior to formats such as COFF. You may be able to use DWARF2 (‘-gdwarf-2’), which is also an effective form for debug info. See section “Options for Debugging Your Program or gnu CC” in Using gnu CC, for more information.

8.3 Artificial arrays

It is often useful to print out several successive objects of the same type in memory; a section of an array, or an array of dynamically determined size for which only a pointer exists in the program.

You can do this by referring to a contiguous span of memory as an artificial array, using the binary operator ‘@’. The left operand of ‘@’ should be the first element of the desired array and be an individual object. The right operand should be the desired length of the array. The result is an array value whose elements are all of the type of the left argument. The first element is actually the left argument; the second element comes from bytes of memory immediately following those that hold the first element, and so on. Here is an example. If a program says

     int *array = (int *) malloc (len * sizeof (int));

you can print the contents of array with

     p *array@len

The left operand of ‘@’ must reside in memory. Array values made with ‘@’ in this way behave just like other arrays in terms of subscripting, and are coerced to pointers when used in expressions. Artificial arrays most often appear in expressions via the value history (see Section 8.8 [Value history], page 77), after printing one out.

Another way to create an artificial array is to use a cast. This re-interprets a value as if it were an array. The value need not be in memory:

     ([No value for ``GDBP'']) p/x (short[2])0x12345678
     $1 = {0x1234, 0x5678}

As a convenience, if you leave the array length out (as in ‘(type[])value’) [No value for “GDBN”] calculates the size to fill the value (as ‘sizeof(value)/sizeof(type)’:

     ([No value for ``GDBP'']) p/x (short[])0x12345678
     $2 = {0x1234, 0x5678}

Sometimes the artificial array mechanism is not quite enough; in moderately complex data structures, the elements of interest may not actually be adjacent—for example, if you are interested in the values of pointers in an array. One useful work-around in this situation is to use a convenience variable (see Section 8.9 [Convenience variables], page 78) as a counter in an expression that prints the first interesting value, and then repeat that expression via^{⟨}\underline{\overline{RET}}^{⟩}. For instance, suppose you have an array dtab of pointers to structures, and you are interested in the values of a field fv in each structure. Here is an example of what you might type:

     set $i = 0
     p dtab[$i++]->fv

⟨\underline{\overline{RET}}⟩ ⟨\underline{\overline{RET}}⟩ ...

8.4 Output formats

By default, [No value for “GDBN”] prints a value according to its data type. Sometimes this is not what you want. For example, you might want to print a number in hex, or a pointer in decimal. Or you might want to view data in memory at a certain address as a character string or as an instruction. To do these things, specify an output format when you print a value.

The simplest use of output formats is to say how to print a value already computed. This is done by starting the arguments of the print command with a slash and a format letter. The format letters supported are:

x Regard the bits of the value as an integer, and print the integer in hexadecimal.

d Print as integer in signed decimal.

u Print as integer in unsigned decimal.

o Print as integer in octal.

t Print as integer in binary. The letter ‘t’ stands for “two”.^{1}

a Print as an address, both absolute in hexadecimal and as an offset from the nearest preceding symbol. You can use this format used to discover where (in what function) an unknown address is located:

1 ‘b’ cannot be used because these format letters are also used with the x command, where ‘b’ stands for

“byte”; see Section 8.5 [Examining memory], page 69.

               ([No value for ``GDBP'']) p/a 0x54320
               $3 = 0x54320 <_initialize_vx+396>

The command info symbol 0x54320 yields similar results. See Chapter 11 [Symbols], page 115.

c Regard as an integer and print it as a character constant.

f Regard the bits of the value as a floating point number and print using typical floating point syntax.

For example, to print the program counter in hex (see Section 8.10 [Registers], page 79), type

     p/x $pc

Note that no space is required before the slash; this is because command names in [No value for “GDBN”] cannot contain a slash.

To reprint the last value in the value history with a different format, you can use the print command with just a format and no expression. For example, ‘p/x’ reprints the last value in hex.

8.5 Examining memory

You can use the command x (for “examine”) to examine memory in any of several formats, independently of your program’s data types.

x/nfu addr x addr x Use the x command to examine memory.

n, f, and u are all optional parameters that specify how much memory to display and how to format it; addr is an expression giving the address where you want to start displaying memory. If you use defaults for nfu, you need not type the slash ‘/’. Several commands set convenient defaults for addr.

n, the repeat count The repeat count is a decimal integer; the default is 1. It specifies how much memory (counting by units u) to display.

f, the display format The display format is one of the formats used by print, ‘s’ (null-terminated string), or ‘i’ (machine instruction). The default is ‘x’ (hexadecimal) initially. The default changes each time you use either x or print.

u, the unit size The unit size is any of

b Bytes.

h Halfwords (two bytes).

w Words (four bytes). This is the initial default.

g Giant words (eight bytes).

Each time you specify a unit size with x, that size becomes the default unit the next time you use x. (For the ‘s’ and ‘i’ formats, the unit size is ignored and is normally not written.)

addr, starting display address addr is the address where you want [No value for “GDBN”] to begin displaying memory. The expression need not have a pointer value (though it may); it is always interpreted as an integer address of a byte of memory. See Section 8.1 [Expressions], page 65, for more information on expressions. The default for addr is usually just after the last address examined—but several other commands also set the default address: info breakpoints (to the address of the last breakpoint listed), info line (to the starting address of a line), and print (if you use it to display a value from memory).

For example, ‘x/3uh 0x54320’ is a request to display three halfwords (h) of memory, formatted as unsigned decimal integers (‘u’), starting at address 0x54320. ‘x/4xw $sp’ prints the four words (‘w’) of memory above the stack pointer (here, ‘$sp’; see Section 8.10 [Registers], page 79) in hexadecimal (‘x’).

Since the letters indicating unit sizes are all distinct from the letters specifying output formats, you do not have to remember whether unit size or format comes first; either order works. The output specifications ‘4xw’ and ‘4wx’ mean exactly the same thing. (However, the count n must come first; ‘wx4’ does not work.)

Even though the unit size u is ignored for the formats ‘s’ and ‘i’, you might still want to use a count n; for example, ‘3i’ specifies that you want to see three machine instructions, including any operands. The command disassemble gives an alternative way of inspecting machine instructions; see Section 7.4 [Source and machine code], page 62.

All the defaults for the arguments to x are designed to make it easy to continue scanning memory with minimal specifications each time you use x. For example, after you have inspected three machine instructions with ‘x/3i addr’, you can inspect the next seven with just ‘x/7’. If you use ^{⟨}\underline{\overline{RET}}^{⟩} to repeat the x command, the repeat count n is used again; the other arguments default as for successive uses of x.

The addresses and contents printed by the x command are not saved in the value history because there is often too much of them and they would get in the way. Instead, [No value for “GDBN”] makes these values available for subsequent use in expressions as values of the convenience variables $_and $__. After an x command, the last address examined is available for use in expressions in the convenience variable $_. The contents of that address, as examined, are available in the convenience variable $__.

If the x command has a repeat count, the address and contents saved are from the last memory unit printed; this is not the same as the last address printed if several units were printed on the last line of output.

8.6 Automatic display

If you find that you want to print the value of an expression frequently (to see how it changes), you might want to add it to the automatic display list so that [No value for “GDBN”] prints its value each time your program stops. Each expression added to the list is given a number to identify it; to remove an expression from the list, you specify that number. The automatic display looks like this:

2: foo = 38
3: bar[5] = (struct hack *) 0x3804

This display shows item numbers, expressions and their current values. As with displays you request manually using x or print, you can specify the output format you prefer; in fact, display decides whether to use print or x depending on how elaborate your format specification is—it uses x if you specify a unit size, or one of the two formats (‘i’ and ‘s’) that are only supported by x; otherwise it uses print.

display expr

Add the expression expr to the list of expressions to display each time your program stops. See Section 8.1 [Expressions], page 65. display does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} again after using it.

display/fmt expr

For fmt specifying only a display format and not a size or count, add the expression expr to the auto-display list but arrange to display it each time in the specified format fmt. See Section 8.4 [Output formats], page 68.

display/fmt addr

For fmt ‘i’ or ‘s’, or including a unit-size or a number of units, add the expression addr as a memory address to be examined each time your program stops. Examining means in effect doing ‘x/fmt addr’. See Section 8.5 [Examining memory], page 69.

For example, ‘display/i $pc’ can be helpful, to see the machine instruction about to be executed each time execution stops (‘$pc’ is a common name for the program counter; see Section 8.10 [Registers], page 79).

undisplay dnums...
delete display dnums...

Remove item numbers dnums from the list of expressions to display. undisplay does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} after using it. (Otherwise you would just get the error ‘No display number ...’.)

disable display dnums...

Disable the display of item numbers dnums. A disabled display item is not printed automatically, but is not forgotten. It may be enabled again later.

enable display dnums...

Enable display of item numbers dnums. It becomes effective once again in auto display of its expression, until you specify otherwise.

display Display the current values of the expressions on the list, just as is done when your program stops.

info display

Print the list of expressions previously set up to display automatically, each one with its item number, but without showing the values. This includes disabled expressions, which are marked as such. It also includes expressions which would not be displayed right now because they refer to automatic variables not currently available.

If a display expression refers to local variables, then it does not make sense outside the lexical context for which it was set up. Such an expression is disabled when execution enters a context where one of its variables is not defined. For example, if you give the command display last_char while inside a function with an argument last_char, [No value for “GDBN”] displays this argument while your program continues to stop inside that function. When it stops elsewhere—where there is no variable last_char—the display is disabled automatically. The next time your program stops where last_char is meaningful, you can enable the display expression once again.

8.7 Print settings

[No value for “GDBN”] provides the following ways to control how arrays, structures, and symbols are printed. These settings are useful for debugging programs in any language:

set print address
set print address on

[No value for “GDBN”] prints memory addresses showing the location of stack traces, structure values, pointer values, breakpoints, and so forth, even when it also displays the contents of those addresses. The default is on. For example, this is what a stack frame display looks like with set print address on:

               ([No value for ``GDBP'']) f
               #0  set_quotes (lq=0x34c78 "<<", rq=0x34c88 ">>")
                   at input.c:530
               530         if (lquote != def_lquote)
set print address off

Do not print addresses when displaying their contents. For example, this is the same stack frame displayed with set print address off:

               ([No value for ``GDBP'']) set print addr off
               ([No value for ``GDBP'']) f
               #0  set_quotes (lq="<<", rq=">>") at input.c:530
               530         if (lquote != def_lquote)

You can use ‘set print address off’ to eliminate all machine dependent displays from the [No value for “GDBN”] interface. For example, with print address off, you should get the same text for backtraces on all machines— whether or not they involve pointer arguments.

show print address

Show whether or not addresses are to be printed.

When [No value for “GDBN”] prints a symbolic address, it normally prints the closest earlier symbol plus an offset. If that symbol does not uniquely identify the address (for example, it is a name whose scope is a single source file), you may need to clarify. One way to do this is with info line, for example ‘info line *0x4537’. Alternately, you can set [No value for “GDBN”] to print the source file and line number when it prints a symbolic address:

set print symbol-filename on

Tell [No value for “GDBN”] to print the source file name and line number of a symbol in the symbolic form of an address.

set print symbol-filename off

Do not print source file name and line number of a symbol. This is the default.

show print symbol-filename

Show whether or not [No value for “GDBN”] will print the source file name and line number of a symbol in the symbolic form of an address.

Another situation where it is helpful to show symbol filenames and line numbers is when disassembling code; [No value for “GDBN”] shows you the line number and source file that corresponds to each instruction.

Also, you may wish to see the symbolic form only if the address being printed is reasonably close to the closest earlier symbol:

set print max-symbolic-offset max-offset

Tell [No value for “GDBN”] to only display the symbolic form of an address if the offset between the closest earlier symbol and the address is less than maxoffset. The default is 0, which tells [No value for “GDBN”] to always print the symbolic form of an address if any symbol precedes it.

show print max-symbolic-offset

Ask how large the maximum offset is that [No value for “GDBN”] prints in a symbolic address.

If you have a pointer and you are not sure where it points, try ‘set print symbol-filename on’. Then you can determine the name and source file location of the variable where it points, using ‘p/a pointer’. This interprets the address in symbolic form. For example, here [No value for “GDBN”] shows that a variable ptt points at another variable t, defined in ‘hi2.c’:

     ([No value for ``GDBP'']) set print symbol-filename on
     ([No value for ``GDBP'']) p/a ptt
     $4 = 0xe008 <t in hi2.c>

Warning: For pointers that point to a local variable, ‘p/a’ does not show the symbol name and filename of the referent, even with the appropriate set print options turned on.

Other settings control how different kinds of objects are printed:

set print array
set print array on

Pretty print arrays. This format is more convenient to read, but uses more space. The default is off.

set print array off

Return to compressed format for arrays.

show print array

Show whether compressed or pretty format is selected for displaying arrays. set print elements number-of-elements

Set a limit on how many elements of an array [No value for “GDBN”] will print. If [No value for “GDBN”] is printing a large array, it stops printing after it has printed the number of elements set by the set print elements command. This limit also applies to the display of strings. When [No value for “GDBN”] starts, this limit is set to 200. Setting number-of-elements to zero means that the printing is unlimited.

show print elements

Display the number of elements of a large array that [No value for “GDBN”] will print. If the number is 0, then the printing is unlimited.

set print null-stop

Cause [No value for “GDBN”] to stop printing the characters of an array when the first null is encountered. This is useful when large arrays actually contain only short strings. The default is off.

set print pretty on

Cause [No value for “GDBN”] to print structures in an indented format with one member per line, like this:

     $1 = {
       next = 0x0,
       flags = {
         sweet = 1,
         sour = 1
       },
       meat = 0x54 "Pork"
     }
set print pretty off

Cause [No value for “GDBN”] to print structures in a compact format, like this:

     $1 = {next = 0x0, flags = {sweet = 1, sour = 1}, \
     meat = 0x54 "Pork"}

This is the default format.

show print pretty

Show which format [No value for “GDBN”] is using to print structures.

set print sevenbit-strings on

Print using only seven-bit characters; if this option is set, [No value for “GDBN”] displays any eight-bit characters (in strings or character values) using the notation \nnn. This setting is best if you are working in English (ascii) and you use the high-order bit of characters as a marker or “meta” bit.

set print sevenbit-strings off

Print full eight-bit characters. This allows the use of more international character sets, and is the default.

show print sevenbit-strings

Show whether or not [No value for “GDBN”] is printing only seven-bit characters.

set print union on

Tell [No value for “GDBN”] to print unions which are contained in structures. This is the default setting.

set print union off

Tell [No value for “GDBN”] not to print unions which are contained in structures.

show print union

Ask [No value for “GDBN”] whether or not it will print unions which are contained in structures. For example, given the declarations

     typedef enum {Tree, Bug} Species;
     typedef enum {Big_tree, Acorn, Seedling} Tree_forms;
     typedef enum {Caterpillar, Cocoon, Butterfly}
                   Bug_forms;
     struct thing {
       Species it;
       union {
         Tree_forms tree;
         Bug_forms bug;
       } form;
     };
     struct thing foo = {Tree, {Acorn}};

with set print union on in effect ‘p foo’ would print

     $1 = {it = Tree, form = {tree = Acorn, bug = Cocoon}}

and with set print union off in effect it would print

     $1 = {it = Tree, form = {...}}

These settings are of interest when debugging C++ programs:

set print demangle
set print demangle on

Print C++ names in their source form rather than in the encoded (“mangled”) form passed to the assembler and linker for type-safe linkage. The default is on.

show print demangle

Show whether C++ names are printed in mangled or demangled form.

set print asm-demangle
set print asm-demangle on

Print C++ names in their source form rather than their mangled form, even in assembler code printouts such as instruction disassemblies. The default is off.

show print asm-demangle

Show whether C++ names in assembly listings are printed in mangled or demangled form.

set demangle-style style

Choose among several encoding schemes used by different compilers to represent C++ names. The choices for style are currently:

auto Allow [No value for “GDBN”] to choose a decoding style by inspecting your program.

gnu Decode based on the gnu C++ compiler (g++) encoding algorithm. This is the default.

hp Decode based on the HP ANSI C++ (aCC) encoding algorithm.

lucid Decode based on the Lucid C++ compiler (lcc) encoding algorithm.

arm Decode using the algorithm in the C++ Annotated Reference Manual. Warning: this setting alone is not sufficient to allow debugging cfront-generated executables. [No value for “GDBN”] would require further enhancement to permit that.

If you omit style, you will see a list of possible formats.

show demangle-style

Display the encoding style currently in use for decoding C++ symbols.

set print object
set print object on

When displaying a pointer to an object, identify the actual (derived) type of the object rather than the declared type, using the virtual function table.

set print object off

Display only the declared type of objects, without reference to the virtual function table. This is the default setting.

show print object

Show whether actual, or declared, object types are displayed.

set print static-members
set print static-members on

Print static members when displaying a C++ object. The default is on.

set print static-members off

Do not print static members when displaying a C++ object.

show print static-members

Show whether C++ static members are printed, or not.

set print vtbl
set print vtbl on

Pretty print C++ virtual function tables. The default is off. (The vtbl commands do not work on programs compiled with the HP ANSI C++ compiler (aCC).)

set print vtbl off

Do not pretty print C++ virtual function tables.

show print vtbl

Show whether C++ virtual function tables are pretty printed, or not.

8.8 Value history

Values printed by the print command are saved in the [No value for “GDBN”] value history. This allows you to refer to them in other expressions. Values are kept until the symbol table is re-read or discarded (for example with the file or symbol-file commands). When the symbol table changes, the value history is discarded, since the values may contain pointers back to the types defined in the symbol table.

The values printed are given history numbers by which you can refer to them. These are successive integers starting with one. print shows you the history number assigned to a value by printing ‘$num = ’ before the value; here num is the history number.

To refer to any previous value, use ‘$’ followed by the value’s history number. The way print labels its output is designed to remind you of this. Just $ refers to the most recent value in the history, and $$ refers to the value before that. $$n refers to the nth value from the end; $$2 is the value just prior to $$, $$1 is equivalent to $$, and $$0 is equivalent to

$.

For example, suppose you have just printed a pointer to a structure and want to see the contents of the structure. It suffices to type

     p *$

If you have a chain of structures where the component next points to the next one, you can print the contents of the next one with this:

     p *$.next

You can print successive links in the chain by repeating this command—which you can do by just typing ^{⟨}\underline{\overline{RET}}^{⟩}.

Note that the history records values, not expressions. If the value of x is 4 and you type these commands:

     print x
     set x=5

then the value recorded in the value history by the print command remains 4 even though the value of x has changed.

show values

Print the last ten values in the value history, with their item numbers. This is like ‘p $$9’ repeated ten times, except that show values does not change the history.

show values n

Print ten history values centered on history item number n.

show values +

Print ten history values just after the values last printed. If no more values are available, show values + produces no display.

Pressing ^{⟨}\underline{\overline{RET}}^{⟩} to repeat show values n has exactly the same effect as ‘show values +’.

8.9 Convenience variables

[No value for “GDBN”] provides convenience variables that you can use within [No value for “GDBN”] to hold on to a value and refer to it later. These variables exist entirely within [No value for “GDBN”]; they are not part of your program, and setting a convenience variable has no direct effect on further execution of your program. That is why you can use them freely.

Convenience variables are prefixed with ‘$’. Any name preceded by ‘$’ can be used for a convenience variable, unless it is one of the predefined machine-specific register names (see Section 8.10 [Registers], page 79). (Value history references, in contrast, are numbers preceded by ‘$’. See Section 8.8 [Value history], page 77.)

You can save a value in a convenience variable with an assignment expression, just as you would set a variable in your program. For example:

     set $foo = *object_ptr

would save in $foo the value contained in the object pointed to by object_ptr.

Using a convenience variable for the first time creates it, but its value is void until you assign a new value. You can alter the value with another assignment at any time.

Convenience variables have no fixed types. You can assign a convenience variable any type of value, including structures and arrays, even if that variable already has a value of a different type. The convenience variable, when used as an expression, has the type of its current value.

show convenience

Print a list of convenience variables used so far, and their values. Abbreviated

          show conv.

One of the ways to use a convenience variable is as a counter to be incremented or a pointer to be advanced. For example, to print a field from successive elements of an array of structures:

     set $i = 0
     print bar[$i++]->contents

Repeat that command by typing ^{⟨}\underline{\overline{RET}}^{⟩}.

Some convenience variables are created automatically by [No value for “GDBN”] and given values likely to be useful.

$_The variable $_is automatically set by the x command to the last address examined (see Section 8.5 [Examining memory], page 69). Other commands which provide a default address for x to examine also set $_to that address; these commands include info line and info breakpoint. The type of $_is void * except when set by the x command, in which case it is a pointer to the type of $__.

$__The variable $__is automatically set by the x command to the value found in the last address examined. Its type is chosen to match the format in which the data was printed.

$_exitcode

The variable $_exitcode is automatically set to the exit code when the program being debugged terminates.

On HP-UX systems, if you refer to a function or variable name that begins with a dollar sign, [No value for “GDBN”] searches for a user or system name first, before it searches for a convenience variable.

8.10 Registers

You can refer to machine register contents, in expressions, as variables with names starting with ‘$’. The names of registers are different for each machine; use info registers to see the names used on your machine.

info registers

Print the names and values of all registers except floating-point registers (in the selected stack frame).

info all-registers

Print the names and values of all registers, including floating-point registers.

info registers regname ...

Print the relativized value of each specified register regname. As discussed in detail below, register values are normally relative to the selected stack frame. regname may be any register name valid on the machine you are using, with or without the initial ‘$’.

[No value for “GDBN”] has four “standard” register names that are available (in expressions) on most machines—whenever they do not conflict with an architecture’s canonical mnemonics for registers. The register names $pc and $sp are used for the program counter register and the stack pointer. $fp is used for a register that contains a pointer to the current stack frame, and $ps is used for a register that contains the processor status. For example, you could print the program counter in hex with

     p/x $pc

or print the instruction to be executed next with

     x/i $pc

or add four to the stack pointer^{2} with

     set $sp += 4

Whenever possible, these four standard register names are available on your machine even though the machine has different canonical mnemonics, so long as there is no conflict. The info registers command shows the canonical names. For example, on the SPARC, info registers displays the processor status register as $psr but you can also refer to it as $ps; and on x86-based machines $ps is an alias for the eflags register.

2 This is a way of removing one word from the stack, on machines where stacks grow downward in memory

(most machines, nowadays). This assumes that the innermost stack frame is selected; setting $sp is not

allowed when other stack frames are selected. To pop entire frames off the stack, regardless of machine

architecture, use return; see Section 12.4 [Returning from a function], page 121.

[No value for “GDBN”] always considers the contents of an ordinary register as an integer when the register is examined in this way. Some machines have special registers which can hold nothing but floating point; these registers are considered to have floating point values. There is no way to refer to the contents of an ordinary register as floating point value (although you can print it as a floating point value with ‘print/f $regname’).

Some registers have distinct “raw” and “virtual” data formats. This means that the data format in which the register contents are saved by the operating system is not the same one that your program normally sees. For example, the registers of the 68881 floating point coprocessor are always saved in “extended” (raw) format, but all C programs expect to work with “double” (virtual) format. In such cases, [No value for “GDBN”] normally works with the virtual format only (the format that makes sense for your program), but the info registers command prints the data in both formats.

Normally, register values are relative to the selected stack frame (see Section 6.3 [Selecting a frame], page 55). This means that you get the value that the register would contain if all stack frames farther in were exited and their saved registers restored. In order to see the true contents of hardware registers, you must select the innermost frame (with ‘frame 0’).

However, [No value for “GDBN”] must deduce where registers are saved, from the machine code generated by your compiler. If some registers are not saved, or if [No value for “GDBN”] is unable to locate the saved registers, the selected stack frame makes no difference.

8.11 Floating point hardware

Depending on the configuration, [No value for “GDBN”] may be able to give you more information about the status of the floating point hardware.

info float

Display hardware-dependent information about the floating point unit. The exact contents and layout vary depending on the floating point chip. Currently, ‘info float’ is supported on the ARM and x86 machines.

8.12 Memory Region Attributes

Memory region attributes allow you to describe special handling required by regions of your target’s memory. [No value for “GDBN”] uses attributes to determine whether to allow certain types of memory accesses; whether to use specific width accesses; and whether to cache target memory.

Defined memory regions can be individually enabled and disabled. When a memory region is disabled, [No value for “GDBN”] uses the default attributes when accessing memory in that region. Similarly, if no memory regions have been defined, [No value for “GDBN”] uses the default attributes when accessing all memory.

When a memory region is defined, it is given a number to identify it; to enable, disable, or remove a memory region, you specify that number. mem address1 address1 attributes...

Define memory region bounded by address1 and address2 with attributes attributes. . ..

delete mem nums...

Remove memory region numbers nums.

disable mem nums...

Disable memory region numbers nums. A disabled memory region is not forgotten. It may be enabled again later.

enable mem nums...

Enable memory region numbers nums.

info mem Print a table of all defined memory regions, with the following columns for each region.

Memory Region Number Enabled or Disabled.

Enabled memory regions are marked with ‘y’. Disabled memory regions are marked with ‘n’.

Lo Address The address defining the inclusive lower bound of the memory region.

Hi Address The address defining the exclusive upper bound of the memory region.

Attributes The list of attributes set for this memory region.

8.12.1 Attributes

8.12.1.1 Memory Access Mode

The access mode attributes set whether [No value for “GDBN”] may make read or write accesses to a memory region.

While these attributes prevent [No value for “GDBN”] from performing invalid memory accesses, they do nothing to prevent the target system, I/O DMA, etc. from accessing memory.

ro Memory is read only.

wo Memory is write only.

rw Memory is read/write (default).

8.12.1.2 Memory Access Size

The acccess size attributes tells [No value for “GDBN”] to use specific sized accesses in the memory region. Often memory mapped device registers require specific sized accesses. If no access size attribute is specified, [No value for “GDBN”] may use accesses of any size.

8 Use 8 bit memory accesses. 16 Use 16 bit memory accesses. 32 Use 32 bit memory accesses. 64 Use 64 bit memory accesses.

8.12.1.3 Data Cache

The data cache attributes set whether [No value for “GDBN”] will cache target memory. While this generally improves performance by reducing debug protocol overhead, it can lead to incorrect results because [No value for “GDBN”] does not know about volatile variables or memory mapped device registers. cache Enable [No value for “GDBN”] to cache target memory.

nocache (default)

Disable [No value for “GDBN”] from caching target memory.

9 Tracepoints

In some applications, it is not feasible for the debugger to interrupt the program’s execution long enough for the developer to learn anything helpful about its behavior. If the program’s correctness depends on its real-time behavior, delays introduced by a debugger might cause the program to change its behavior drastically, or perhaps fail, even when the code itself is correct. It is useful to be able to observe the program’s behavior without interrupting it.

Using [No value for “GDBN”]’s trace and collect commands, you can specify locations in the program, called tracepoints, and arbitrary expressions to evaluate when those tracepoints are reached. Later, using the tfind command, you can examine the values those expressions had when the program hit the tracepoints. The expressions may also denote objects in memory—structures or arrays, for example—whose values [No value for “GDBN”] should record; while visiting a particular tracepoint, you may inspect those objects as if they were in memory at that moment. However, because [No value for “GDBN”] records these values without interacting with you, it can do so quickly and unobtrusively, hopefully not disturbing the program’s behavior.

The tracepoint facility is currently available only for remote targets. See Chapter 14 [Targets], page 129.

This chapter describes the tracepoint commands and features.

9.1 Commands to Set Tracepoints

Before running such a trace experiment, an arbitrary number of tracepoints can be set. Like a breakpoint (see Section 5.1.1 [Set Breaks], page 34), a tracepoint has a number assigned to it by [No value for “GDBN”]. Like with breakpoints, tracepoint numbers are successive integers starting from one. Many of the commands associated with tracepoints take the tracepoint number as their argument, to identify which tracepoint to work on.

For each tracepoint, you can specify, in advance, some arbitrary set of data that you want the target to collect in the trace buffer when it hits that tracepoint. The collected data can include registers, local variables, or global data. Later, you can use [No value for “GDBN”] commands to examine the values these data had at the time the tracepoint was hit.

This section describes commands to set tracepoints and associated conditions and actions.

9.1.1 Create and Delete Tracepoints

trace The trace command is very similar to the break command. Its argument can be a source line, a function name, or an address in the target program. See Section 5.1.1 [Set Breaks], page 34. The trace command defines a tracepoint, which is a point in the target program where the debugger will briefly stop, collect some data, and then allow the program to continue. Setting a tracepoint or changing its commands doesn’t take effect until the next tstart command; thus, you cannot change the tracepoint attributes once a trace experiment is running. Here are some examples of using the trace command:

     ([No value for ``GDBP'']) trace foo.c:121  // a source file and line number
     ([No value for ``GDBP'']) trace +2  // 2 lines forward
     ([No value for ``GDBP'']) trace my function // first source line of function
     ([No value for ``GDBP'']) trace *my function // EXACT start address of function
     ([No value for ``GDBP'']) trace *0x2117c4  // an address

You can abbreviate trace as tr. The convenience variable $tpnum records the tracepoint number of the most recently set tracepoint.

delete tracepoint [num]

Permanently delete one or more tracepoints. With no argument, the default is to delete all tracepoints. Examples:

     ([No value for ``GDBP'']) delete trace 1 2 3 // remove three tracepoints
     ([No value for ``GDBP'']) delete trace  // remove all tracepoints

You can abbreviate this command as del tr.

9.1.2 Enable and Disable Tracepoints

disable tracepoint [num]

Disable tracepoint num, or all tracepoints if no argument num is given. A disabled tracepoint will have no effect during the next trace experiment, but it is not forgotten. You can re-enable a disabled tracepoint using the enable

tracepoint command.
enable tracepoint [num]

Enable tracepoint num, or all tracepoints. The enabled tracepoints will become effective the next time a trace experiment is run.

9.1.3 Tracepoint Passcounts

passcount [n [num]]

Set the passcount of a tracepoint. The passcount is a way to automatically stop a trace experiment. If a tracepoint’s passcount is n, then the trace experiment will be automatically stopped on the n’th time that tracepoint is hit. If the tracepoint number num is not specified, the passcount command sets the passcount of the most recently defined tracepoint. If no passcount is given, the trace experiment will run until stopped explicitly by the user. Examples:

([No value for ``GDBP'']) passcount 5 2 // Stop on the 5th execution of tracepoint 2
([No value for ``GDBP'']) passcount 12 // Stop on the 12th execution of the
                                // most recently defined tracepoint.
([No value for ``GDBP'']) trace foo
([No value for ``GDBP'']) pass 3
([No value for ``GDBP'']) trace bar
([No value for ``GDBP'']) pass 2
([No value for ``GDBP'']) trace baz
([No value for ``GDBP'']) pass 1  // Stop tracing when foo has been
                                 // executed 3 times OR when bar has
                                 // been executed 2 times
                                 // OR when baz has been executed 1 time.

9.1.4 Tracepoint Action Lists

actions [num]

This command will prompt for a list of actions to be taken when the tracepoint is hit. If the tracepoint number num is not specified, this command sets the actions for the one that was most recently defined (so that you can define a tracepoint and then say actions without bothering about its number). You specify the actions themselves on the following lines, one action at a time, and terminate the actions list with a line containing just end. So far, the only

defined actions are collect and while-stepping.

To remove all actions from a tracepoint, type ‘actions num’ and follow it immediately with ‘end’.

([No value for ``GDBP'']) collect data // collect some data
([No value for ``GDBP'']) while-stepping 5 // single-step 5 times and collect data
([No value for ``GDBP'']) end  // signals the end of actions.

In the following example, the action list begins with collect commands indicating the things to be collected when the tracepoint is hit. Then, in order to single-step and collect additional data following the tracepoint, a whilestepping command is used, followed by the list of things to be collected while stepping. The while-stepping command is terminated by its own separate end command. Lastly, the action list is terminated by an end command.

([No value for ``GDBP'']) trace foo
([No value for ``GDBP'']) actions
Enter actions for tracepoint 1, one per line:
> collect bar,baz
> collect $regs
> while-stepping 12
  > collect $fp, $sp
  > end
end

collect expr1, expr2, ...

Collect values of the given expressions when the tracepoint is hit. This command accepts a comma-separated list of any valid expressions. In addition to global, static, or local variables, the following special arguments are supported:

$regs collect all registers

$args collect all function arguments

$locals collect all local variables.

You can give several consecutive collect commands, each one with a single argument, or one collect command with several arguments separated by commas: the effect is the same. The command info scope (see Chapter 11 [Symbols], page 115) is particularly useful for figuring out what data to collect.

while-stepping n

Perform n single-step traces after the tracepoint, collecting new data at each step. The while-stepping command is followed by the list of what to collect while stepping (followed by its own end command):

     > while-stepping 12
       > collect $regs, myglobal
       > end
     >
You may abbreviate while-stepping as ws or stepping.

9.1.5 Listing Tracepoints

info tracepoints [num]

Display information the tracepoint num. If you don’t specify a tracepoint number displays information about all the tracepoints defined so far. For each tracepoint, the following information is shown:

• its number

• whether it is enabled or disabled

• its address

• its passcount as given by the passcount n command

• its step count as given by the while-stepping n command

• where in the source files is the tracepoint set

• its action list as given by the actions command

     ([No value for ``GDBP'']) info trace
     Num Enb Address    PassC StepC What
     1   y   0x002117c4 0     0     <gdb_asm>
     2   y   0x0020dc64 0     0     in gdb_test at gdb_test.c:375
     3   y   0x0020b1f4 0     0     in collect_data at ../foo.c:1741
     ([No value for ``GDBP''])

This command can be abbreviated info tp.

9.1.6 Starting and Stopping Trace Experiment

tstart This command takes no arguments. It starts the trace experiment, and begins collecting data. This has the side effect of discarding all the data collected in the trace buffer during the previous trace experiment.

tstop This command takes no arguments. It ends the trace experiment, and stops collecting data. Note: a trace experiment and data collection may stop automatically if any tracepoint’s passcount is reached (see Section 9.1.3 [Tracepoint Passcounts], page 84), or if the trace buffer becomes full.

tstatus This command displays the status of the current trace data collection.

Here is an example of the commands we described so far:

     ([No value for ``GDBP'']) trace gdb c test
     ([No value for ``GDBP'']) actions
     Enter actions for tracepoint #1, one per line.
     > collect $regs,$locals,$args
     > while-stepping 11
       > collect $regs
       > end
     > end
     ([No value for ``GDBP'']) tstart
     [time passes ...]
     ([No value for ``GDBP'']) tstop

9.2 Using the collected data

After the tracepoint experiment ends, you use [No value for “GDBN”] commands for examining the trace data. The basic idea is that each tracepoint collects a trace snapshot every time it is hit and another snapshot every time it single-steps. All these snapshots are consecutively numbered from zero and go into a buffer, and you can examine them later. The way you examine them is to focus on a specific trace snapshot. When the remote stub is focused on a trace snapshot, it will respond to all [No value for “GDBN”] requests for memory and registers by reading from the buffer which belongs to that snapshot, rather than from real memory or registers of the program being debugged. This means that all [No value for “GDBN”] commands (print, info registers, backtrace, etc.) will behave as if we were currently debugging the program state as it was when the tracepoint occurred. Any requests for data that are not in the buffer will fail.

9.2.1 tfind n

The basic command for selecting a trace snapshot from the buffer is tfind n, which finds trace snapshot number n, counting from zero. If no argument n is given, the next snapshot is selected.

Here are the various forms of using the tfind command.

tfind start

Find the first snapshot in the buffer. This is a synonym for tfind 0 (since 0 is the number of the first snapshot).

tfind none

Stop debugging trace snapshots, resume live debugging.

tfind end Same as ‘tfind none’.

tfind No argument means find the next trace snapshot.

tfind - Find the previous trace snapshot before the current one. This permits retracing earlier steps.

tfind tracepoint num

Find the next snapshot associated with tracepoint num. Search proceeds forward from the last examined trace snapshot. If no argument num is given, it means find the next snapshot collected for the same tracepoint as the current snapshot.

tfind pc addr

Find the next snapshot associated with the value addr of the program counter. Search proceeds forward from the last examined trace snapshot. If no argument addr is given, it means find the next snapshot with the same value of PC as the current snapshot.

tfind outside addr1, addr2

Find the next snapshot whose PC is outside the given range of addresses.

tfind range addr1, addr2

Find the next snapshot whose PC is between addr1 and addr2.

tfind line [file:]n

Find the next snapshot associated with the source line n. If the optional argument file is given, refer to line n in that source file. Search proceeds forward from the last examined trace snapshot. If no argument n is given, it means find the next line other than the one currently being examined; thus saying tfind line repeatedly can appear to have the same effect as stepping from line to line in a live debugging session.

The default arguments for the tfind commands are specifically designed to make it easy to scan through the trace buffer. For instance, tfind with no argument selects the next trace snapshot, and tfind - with no argument selects the previous trace snapshot. So, by giving one tfind command, and then simply hitting ^{⟨}\underline{\overline{RET}}^{⟩} repeatedly you can examine all the trace snapshots in order. Or, by saying tfind - and then hitting ^{⟨}\underline{\overline{RET}}^{⟩} repeatedly you can examine the snapshots in reverse order. The tfind line command with no argument selects the snapshot for the next source line executed. The tfind pc command with no argument selects the next snapshot with the same program counter (PC) as the current frame. The tfind tracepoint command with no argument selects the next trace snapshot collected by the same tracepoint as the current one.

In addition to letting you scan through the trace buffer manually, these commands make it easy to construct [No value for “GDBN”] scripts that scan through the trace buffer and print out whatever collected data you are interested in. Thus, if we want to examine the PC, FP, and SP registers from each trace frame in the buffer, we can say this:

([No value for ``GDBP'']) tfind start
([No value for ``GDBP'']) while ($trace frame != -1)
> printf "Frame %d, PC = %08X, SP = %08X, FP = %08X\n", \
          $trace_frame, $pc, $sp, $fp
> tfind
> end
Frame 0, PC = 0020DC64, SP = 0030BF3C, FP = 0030BF44
Frame 1, PC = 0020DC6C, SP = 0030BF38, FP = 0030BF44
Frame 2, PC = 0020DC70, SP = 0030BF34, FP = 0030BF44
Frame 3, PC = 0020DC74, SP = 0030BF30, FP = 0030BF44
Frame 4, PC = 0020DC78, SP = 0030BF2C, FP = 0030BF44
Frame 5, PC = 0020DC7C, SP = 0030BF28, FP = 0030BF44
Frame 6, PC = 0020DC80, SP = 0030BF24, FP = 0030BF44
Frame 7, PC = 0020DC84, SP = 0030BF20, FP = 0030BF44
Frame 8, PC = 0020DC88, SP = 0030BF1C, FP = 0030BF44
Frame 9, PC = 0020DC8E, SP = 0030BF18, FP = 0030BF44
Frame 10, PC = 00203F6C, SP = 0030BE3C, FP = 0030BF14

Or, if we want to examine the variable X at each source line in the buffer:

([No value for ``GDBP'']) tfind start
([No value for ``GDBP'']) while ($trace frame != -1)
> printf "Frame %d, X == %d\n", $trace_frame, X
> tfind line
> end
Frame 0, X = 1
Frame 7, X = 2
Frame 13, X = 255
9.2.2 tdump

This command takes no arguments. It prints all the data collected at the current trace snapshot.

([No value for ``GDBP'']) trace 444
([No value for ``GDBP'']) actions
Enter actions for tracepoint #2, one per line:
> collect $regs, $locals, $args, gdb_long_test
> end
([No value for ``GDBP'']) tstart
([No value for ``GDBP'']) tfind line 444
#0  gdb_test (p1=0x11, p2=0x22, p3=0x33, p4=0x44, p5=0x55, p6=0x66)
at gdb_test.c:444
444        printp( "%s: arguments = 0x%X 0x%X 0x%X 0x%X 0x%X 0x%X\n", )

([No value for ``GDBP'']) tdump
Data collected at tracepoint 2, trace frame 1:
d0             0xc4aa0085       -995491707
d1             0x18     24
d2             0x80     128
d3             0x33     51
d4             0x71aea3d        119204413
d5             0x22     34
d6             0xe0     224
d7             0x380035 3670069
a0             0x19e24a 1696330
a1             0x3000668        50333288
a2             0x100    256
a3             0x322000 3284992
a4             0x3000698        50333336
a5             0x1ad3cc 1758156
fp             0x30bf3c 0x30bf3c
sp             0x30bf34 0x30bf34
ps             0x0      0
pc             0x20b2c8 0x20b2c8
fpcontrol      0x0      0
fpstatus       0x0      0
fpiaddr        0x0      0
p = 0x20e5b4 "gdb-test"
p1 = (void *) 0x11
p2 = (void *) 0x22
p3 = (void *) 0x33
p4 = (void *) 0x44
p5 = (void *) 0x55
p6 = (void *) 0x66
gdb_long_test = 17 '\021'
([No value for ``GDBP''])
9.2.3 save-tracepoints filename

This command saves all current tracepoint definitions together with their actions and passcounts, into a file ‘filename’ suitable for use in a later debugging session. To read the saved tracepoint definitions, use the source command (see Section 17.3 [Command Files], page 179).

9.3 Convenience Variables for Tracepoints

(int) $trace_frame

The current trace snapshot (a.k.a. frame) number, or -1 if no snapshot is selected.

(int) $tracepoint

The tracepoint for the current trace snapshot.

(int) $trace_line

The line number for the current trace snapshot.

(char []) $trace_file

The source file for the current trace snapshot.

(char []) $trace_func

The name of the function containing $tracepoint.

Note: $trace_file is not suitable for use in printf, use output instead. Here’s a simple example of using these convenience variables for stepping through all the

trace snapshots and printing some of their data.

([No value for ``GDBP'']) tfind start
([No value for ``GDBP'']) while $trace frame != -1
> output $trace_file
> printf ", line %d (tracepoint #%d)\n", $trace_line, $tracepoint
> tfind
> end

92 Debugging with [No value for “GDBN”]

10 Using [No value for “GDBN”] with Different Languages

Although programming languages generally have common aspects, they are rarely expressed in the same manner. For instance, in ANSI C, dereferencing a pointer p is accomplished by *p, but in Modula-2, it is accomplished by p^. Values can also be represented (and displayed) differently. Hex numbers in C appear as ‘0x1ae’, while in Modula-2 they appear as ‘1AEH’.

Language-specific information is built into [No value for “GDBN”] for some languages, allowing you to express operations like the above in your program’s native language, and allowing [No value for “GDBN”] to output values in a manner consistent with the syntax of your program’s native language. The language you use to build expressions is called the working language.

10.1 Switching between source languages

There are two ways to control the working language—either have [No value for “GDBN”] set it automatically, or select it manually yourself. You can use the set language command for either purpose. On startup, [No value for “GDBN”] defaults to setting the language automatically. The working language is used to determine how expressions you type are interpreted, how values are printed, etc.

In addition to the working language, every source file that [No value for “GDBN”] knows about has its own working language. For some object file formats, the compiler might indicate which language a particular source file is in. However, most of the time [No value for “GDBN”] infers the language from the name of the file. The language of a source file controls whether C++ names are demangled—this way backtrace can show each frame appropriately for its own language. There is no way to set the language of a source file from within [No value for “GDBN”], but you can set the language associated with a filename extension. See Section 10.2 [Displaying the language], page 95.

This is most commonly a problem when you use a program, such as cfront or f2c, that generates C but is written in another language. In that case, make the program use #line directives in its C output; that way [No value for “GDBN”] will know the correct language of the source code of the original program, and will display that source code, not the generated C code.

10.1.1 List of filename extensions and languages

If a source file name ends in one of the following extensions, then [No value for “GDBN”] infers that its language is the one indicated.

‘.c’ C source file

‘.C’
‘.cc’
‘.cp’
‘.cpp’
‘.cxx’

‘.c++’ C++ source file

‘.f’

‘.F’ Fortran source file

‘.ch’
‘.c186’

‘.c286’ CHILL source file

‘.mod’ Modula-2 source file

‘.s’

‘.S’ Assembler source file. This actually behaves almost like C, but [No value for “GDBN”] does not skip over function prologues when stepping.

In addition, you may set the language associated with a filename extension. See Section 10.2 [Displaying the language], page 95.

10.1.2 Setting the working language

If you allow [No value for “GDBN”] to set the language automatically, expressions are interpreted the same way in your debugging session and your program.

If you wish, you may set the language manually. To do this, issue the command ‘set language lang’, where lang is the name of a language, such as c or modula-2. For a list of the supported languages, type ‘set language’.

Setting the language manually prevents [No value for “GDBN”] from updating the working language automatically. This can lead to confusion if you try to debug a program when the working language is not the same as the source language, when an expression is acceptable to both languages—but means different things. For instance, if the current source file were written in C, and [No value for “GDBN”] was parsing Modula-2, a command such as:

     print a = b + c

might not have the effect you intended. In C, this means to add b and c and place the result in a. The result printed would be the value of a. In Modula-2, this means to compare a to the result of b+c, yielding a BOOLEAN value.

10.1.3 Having [No value for “GDBN”] infer the source language

To have [No value for “GDBN”] set the working language automatically, use ‘set language local’ or ‘set language auto’. [No value for “GDBN”] then infers the working language. That is, when your program stops in a frame (usually by encountering a breakpoint), [No value for “GDBN”] sets the working language to the language recorded for the function in that frame. If the language for a frame is unknown (that is, if the function or block corresponding to the frame was defined in a source file that does not have a recognized extension), the current working language is not changed, and [No value for “GDBN”] issues a warning.

This may not seem necessary for most programs, which are written entirely in one source language. However, program modules and libraries written in one source language can be used by a main program written in a different source language. Using ‘set language auto’ in this case frees you from having to set the working language manually.

10.2 Displaying the language

The following commands help you find out which language is the working language, and also what language source files were written in.

show language

Display the current working language. This is the language you can use with commands such as print to build and compute expressions that may involve variables in your program.

info frame

Display the source language for this frame. This language becomes the working language if you use an identifier from this frame. See Section 6.4 [Information about a frame], page 56, to identify the other information listed here.

info source

Display the source language of this source file. See Chapter 11 [Examining the Symbol Table], page 115, to identify the other information listed here.

In unusual circumstances, you may have source files with extensions not in the standard list. You can then set the extension associated with a language explicitly:

set extension-language .ext language

Set source files with extension .ext to be assumed to be in the source language language.

info extensions

List all the filename extensions and the associated languages.

10.3 Type and range checking

Warning: In this release, the [No value for “GDBN”] commands for type and range checking are included, but they do not yet have any effect. This section documents the intended facilities.

Some languages are designed to guard you against making seemingly common errors through a series of compile- and run-time checks. These include checking the type of arguments to functions and operators, and making sure mathematical overflows are caught at run time. Checks such as these help to ensure a program’s correctness once it has been compiled by eliminating type mismatches, and providing active checks for range errors when your program is running.

[No value for “GDBN”] can check for conditions like the above if you wish. Although [No value for “GDBN”] does not check the statements in your program, it can check expressions entered directly into [No value for “GDBN”] for evaluation via the print command, for example. As with the working language, [No value for “GDBN”] can also decide whether or not to check automatically based on your program’s source language. See Section 10.4 [Supported languages], page 97, for the default settings of supported languages.

10.3.1 An overview of type checking

Some languages, such as Modula-2, are strongly typed, meaning that the arguments to operators and functions have to be of the correct type, otherwise an error occurs. These checks prevent type mismatch errors from ever causing any run-time problems. For example,

     1 + 2 ⇒ 3

but

      error 1 + 2.3

The second example fails because the CARDINAL 1 is not type-compatible with the REAL 2.3.

For the expressions you use in [No value for “GDBN”] commands, you can tell the [No value for “GDBN”] type checker to skip checking; to treat any mismatches as errors and abandon the expression; or to only issue warnings when type mismatches occur, but evaluate the expression anyway. When you choose the last of these, [No value for “GDBN”] evaluates expressions like the second example above, but also issues a warning.

Even if you turn type checking off, there may be other reasons related to type that prevent [No value for “GDBN”] from evaluating an expression. For instance, [No value for “GDBN”] does not know how to add an int and a struct foo. These particular type errors have nothing to do with the language in use, and usually arise from expressions, such as the one described above, which make little sense to evaluate anyway.

Each language defines to what degree it is strict about type. For instance, both Modula-2 and C require the arguments to arithmetical operators to be numbers. In C, enumerated types and pointers can be represented as numbers, so that they are valid arguments to mathematical operators. See Section 10.4 [Supported languages], page 97, for further details on specific languages.

[No value for “GDBN”] provides some additional commands for controlling the type checker:

set check type auto

Set type checking on or off based on the current working language. See Section 10.4 [Supported languages], page 97, for the default settings for each language.

set check type on
set check type off

Set type checking on or off, overriding the default setting for the current working language. Issue a warning if the setting does not match the language default. If any type mismatches occur in evaluating an expression while type checking is on, [No value for “GDBN”] prints a message and aborts evaluation of the expression.

set check type warn

Cause the type checker to issue warnings, but to always attempt to evaluate the expression. Evaluating the expression may still be impossible for other reasons. For example, [No value for “GDBN”] cannot add numbers and structures.

show type Show the current setting of the type checker, and whether or not [No value for “GDBN”] is setting it automatically.

10.3.2 An overview of range checking

In some languages (such as Modula-2), it is an error to exceed the bounds of a type; this is enforced with run-time checks. Such range checking is meant to ensure program correctness by making sure computations do not overflow, or indices on an array element access do not exceed the bounds of the array.

For expressions you use in [No value for “GDBN”] commands, you can tell [No value for “GDBN”] to treat range errors in one of three ways: ignore them, always treat them as errors and abandon the expression, or issue warnings but evaluate the expression anyway.

A range error can result from numerical overflow, from exceeding an array index bound, or when you type a constant that is not a member of any type. Some languages, however, do not treat overflows as an error. In many implementations of C, mathematical overflow causes the result to “wrap around” to lower values—for example, if m is the largest integer value, and s is the smallest, then

m + 1 ⇒ s This, too, is specific to individual languages, and in some cases specific to individual

compilers or machines. See Section 10.4 [Supported languages], page 97, for further details on specific languages.

[No value for “GDBN”] provides some additional commands for controlling the range checker:

set check range auto

Set range checking on or off based on the current working language. See Section 10.4 [Supported languages], page 97, for the default settings for each language.

set check range on
set check range off

Set range checking on or off, overriding the default setting for the current working language. A warning is issued if the setting does not match the language default. If a range error occurs and range checking is on, then a message is printed and evaluation of the expression is aborted.

set check range warn

Output messages when the [No value for “GDBN”] range checker detects a range error, but attempt to evaluate the expression anyway. Evaluating the expression may still be impossible for other reasons, such as accessing memory that the process does not own (a typical example from many Unix systems).

show range

Show the current setting of the range checker, and whether or not it is being set automatically by [No value for “GDBN”].

10.4 Supported languages

[No value for “GDBN”] supports C, C++, Fortran, Java, Chill, assembly, and Modula-2. Some [No value for “GDBN”] features may be used in expressions regardless of the language you use: the [No value for “GDBN”] @ and :: operators, and the ‘{type}addr’ construct (see Section 8.1 [Expressions], page 65) can be used with the constructs of any supported language.

The following sections detail to what degree each source language is supported by [No value for “GDBN”]. These sections are not meant to be language tutorials or references, but serve only as a reference guide to what the [No value for “GDBN”] expression parser accepts, and what input and output formats should look like for different languages. There are many good books written on each of these languages; please look to these for a language reference or tutorial.

10.4.1 C and C++

Since C and C++ are so closely related, many features of [No value for “GDBN”] apply to both languages. Whenever this is the case, we discuss those languages together.

The C++ debugging facilities are jointly implemented by the C++ compiler and [No value for “GDBN”]. Therefore, to debug your C++ code effectively, you must compile your C++ programs with a supported C++ compiler, such as gnu g++, or the HP ANSI C++ compiler

(aCC).

For best results when using gnu C++, use the stabs debugging format. You can select that format explicitly with the g++ command-line options ‘-gstabs’ or ‘-gstabs+’. See section “Options for Debugging Your Program or gnu CC” in Using gnu CC, for more information.

10.4.1.1 C and C++ operators

Operators must be defined on values of specific types. For instance, + is defined on numbers, but not on structures. Operators are often defined on groups of types.

For the purposes of C and C++, the following definitions hold:

• Integral types include int with any of its storage-class specifiers; char; enum; and, for

    C++, bool.

• Floating-point types include float, double, and long double (if supported by the target platform).

• Pointer types include all types defined as (type *).

• Scalar types include all of the above.

The following operators are supported. They are listed here in order of increasing precedence:

, The comma or sequencing operator. Expressions in a comma-separated list are evaluated from left to right, with the result of the entire expression being the last expression evaluated.

= Assignment. The value of an assignment expression is the value assigned. Defined on scalar types.

op= Used in an expression of the form a op= b, and translated to a = a op b. op= and = have the same precedence. op is any one of the operators |, ^, &, <<, >>,

          +, -, *, /, %.

?: The ternary operator. a ? b : c can be thought of as: if a then b else c. a should be of an integral type.

|| Logical or. Defined on integral types.

&& Logical and. Defined on integral types.

| Bitwise or. Defined on integral types. ^Bitwise exclusive-or. Defined on integral types. & Bitwise and. Defined on integral types. ==, != Equality and inequality. Defined on scalar types. The value of these expressions

is 0 for false and non-zero for true.

<, >, <=, >=

Less than, greater than, less than or equal, greater than or equal. Defined on scalar types. The value of these expressions is 0 for false and non-zero for true.

<<, >> left shift, and right shift. Defined on integral types. @ The [No value for “GDBN”] “artificial array” operator (see Section 8.1 [Expres-

sions], page 65).

+, - Addition and subtraction. Defined on integral types, floating-point types and

pointer types.

*, /, % Multiplication, division, and modulus. Multiplication and division are defined

on integral and floating-point types. Modulus is defined on integral types.

++, -- Increment and decrement. When appearing before a variable, the operation is performed before the variable is used in an expression; when appearing after it, the variable’s value is used before the operation takes place.

* Pointer dereferencing. Defined on pointer types. Same precedence as ++. & Address operator. Defined on variables. Same precedence as ++.

For debugging C++, [No value for “GDBN”] implements a use of ‘&’ beyond what is allowed in the C++ language itself: you can use ‘&(&ref )’ (or, if you prefer, simply ‘&&ref ’) to examine the address where a C++ reference variable (declared with ‘&ref ’) is stored.

- Negative. Defined on integral and floating-point types. Same precedence as ++. ! Logical negation. Defined on integral types. Same precedence as ++. ~ Bitwise complement operator. Defined on integral types. Same precedence as

          ++.

., -> Structure member, and pointer-to-structure member. For convenience, [No value for “GDBN”] regards the two as equivalent, choosing whether to dereference a pointer based on the stored type information. Defined on struct and

          union data.

.*, ->* Dereferences of pointers to members. [] Array indexing. a[i] is defined as *(a+i). Same precedence as ->. () Function parameter list. Same precedence as ->. :: C++ scope resolution operator. Defined on struct, union, and class types. :: Doubled colons also represent the [No value for “GDBN”] scope operator (see

Section 8.1 [Expressions], page 65). Same precedence as ::, above. If an operator is redefined in the user code, [No value for “GDBN”] usually attempts to

invoke the redefined version instead of using the operator’s predefined meaning.

10.4.1.2 C and C++ constants

[No value for “GDBN”] allows you to express the constants of C and C++ in the following ways:

• Integer constants are a sequence of digits. Octal constants are specified by a leading ‘0’ (i.e. zero), and hexadecimal constants by a leading ‘0x’ or ‘0X’. Constants may also end with a letter ‘l’, specifying that the constant should be treated as a long value.

• Floating point constants are a sequence of digits, followed by a decimal point, followed by a sequence of digits, and optionally followed by an exponent. An exponent is of the form: ‘e[[+]|-]nnn’, where nnn is another sequence of digits. The ‘+’ is optional for positive exponents. A floating-point constant may also end with a letter ‘f’ or ‘F’, specifying that the constant should be treated as being of the float (as opposed to the default double) type; or with a letter ‘l’ or ‘L’, which specifies a long double constant.

• Enumerated constants consist of enumerated identifiers, or their integral equivalents.

• Character constants are a single character surrounded by single quotes ('), or a number—the ordinal value of the corresponding character (usually its ascii value). Within quotes, the single character may be represented by a letter or by escape sequences, which are of the form ‘\nnn’, where nnn is the octal representation of the character’s ordinal value; or of the form ‘\x’, where ‘x’ is a predefined special character—for example, ‘\n’ for newline.

• String constants are a sequence of character constants surrounded by double quotes ("). Any valid character constant (as described above) may appear. Double quotes within the string must be preceded by a backslash, so for instance ‘"a\"b'c"’ is a string of five characters.

• Pointer constants are an integral value. You can also write pointers to constants using the C operator ‘&’.

• Array constants are comma-separated lists surrounded by braces ‘{’ and ‘}’; for example, ‘{1,2,3}’ is a three-element array of integers, ‘{{1,2}, {3,4}, {5,6}}’ is a three-by-two array, and ‘{&"hi", &"there", &"fred"}’ is a three-element array of pointers.

10.4.1.3 C++ expressions

[No value for “GDBN”] expression handling can interpret most C++ expressions. Warning: [No value for “GDBN”] can only debug C++ code if you use the proper compiler. Typically, C++ debugging depends on the use of additional debugging information in the symbol table, and thus requires special support. In particular, if your compiler generates a.out, MIPS ecoff, RS/6000 xcoff, or elf with stabs extensions to the symbol table, these facilities are all available. (With gnu CC, you can use the ‘-gstabs’ option to request stabs debugging extensions explicitly.) Where the object code format is standard coff or dwarf in elf, on the other hand, most of the C++ support in [No value for “GDBN”] does not work.

1. Member function calls are allowed; you can use expressions like

     count = aml->GetOriginal(x, y)

2. While a member function is active (in the selected stack frame), your expressions have the same namespace available as the member function; that is, [No value for “GDBN”] allows implicit references to the class instance pointer this following the same rules as

C++.

3. You can call overloaded functions; [No value for “GDBN”] resolves the function call to the right definition, with some restrictions. [No value for “GDBN”] does not perform overload resolution involving user-defined type conversions, calls to constructors, or instantiations of templates that do not exist in the program. It also cannot handle ellipsis argument lists or default arguments. It does perform integral conversions and promotions, floating-point promotions, arithmetic conversions, pointer conversions, conversions of class objects to base classes, and standard conversions such as those of functions or arrays to pointers; it requires an exact match on the number of function arguments. Overload resolution is always performed, unless you have specified set overloadresolution off. See Section 10.4.1.7 [[No value for “GDBN”] features for C++], page 102. You must specify set overload-resolution off in order to use an explicit function signature to call an overloaded function, as in

     p 'foo(char,int)'('x', 13)

The [No value for “GDBN”] command-completion facility can simplify this; see Section 3.2 [Command completion], page 17.

4. [No value for “GDBN”] understands variables declared as C++ references; you can use them in expressions just as you do in C++ source—they are automatically dereferenced. In the parameter list shown when [No value for “GDBN”] displays a frame, the values of reference variables are not displayed (unlike other variables); this avoids clutter, since references are often used for large structures. The address of a reference variable is always shown, unless you have specified ‘set print address off’.

5. [No value for “GDBN”] supports the C++ name resolution operator ::—your expressions can use it just as expressions in your program do. Since one scope may be defined in another, you can use :: repeatedly if necessary, for example in an expression like ‘scope1::scope2::name’. [No value for “GDBN”] also allows resolving name scope by reference to source files, in both C and C++ debugging (see Section 8.2 [Program variables], page 66).

In addition, when used with HP’s C++ compiler, [No value for “GDBN”] supports calling virtual functions correctly, printing out virtual bases of objects, calling functions in a base subobject, casting objects, and invoking user-defined operators.

10.4.1.4 C and C++ defaults

If you allow [No value for “GDBN”] to set type and range checking automatically, they both default to off whenever the working language changes to C or C++. This happens regardless of whether you or [No value for “GDBN”] selects the working language.

If you allow [No value for “GDBN”] to set the language automatically, it recognizes source files whose names end with ‘.c’, ‘.C’, or ‘.cc’, etc, and when [No value for “GDBN”] enters code compiled from one of these files, it sets the working language to C or C++. See Section 10.1.3 [Having [No value for “GDBN”] infer the source language], page 94, for further details.

10.4.1.5 C and C++ type and range checks

By default, when [No value for “GDBN”] parses C or C++ expressions, type checking is not used. However, if you turn type checking on, [No value for “GDBN”] considers two variables type equivalent if:

• The two variables are structured and have the same structure, union, or enumerated tag.

• The two variables have the same type name, or types that have been declared equivalent through typedef.

Range checking, if turned on, is done on mathematical operations. Array indices are not checked, since they are often used to index a pointer that is not itself an array.

10.4.1.6 [No value for “GDBN”] and C

The set print union and show print union commands apply to the union type. When set to ‘on’, any union that is inside a struct or class is also printed. Otherwise, it appears

as ‘{...}’.

The @ operator aids in the debugging of dynamic arrays, formed with pointers and a memory allocation function. See Section 8.1 [Expressions], page 65.

10.4.1.7 [No value for “GDBN”] features for C++

Some [No value for “GDBN”] commands are particularly useful with C++, and some are designed specifically for use with C++. Here is a summary:

breakpoint menus When you want a breakpoint in a function whose name is overloaded, [No value for “GDBN”] breakpoint menus help you specify which function definition you want. See Section 5.1.8 [Breakpoint menus], page 45.

rbreak regex

Setting breakpoints using regular expressions is helpful for setting breakpoints on overloaded functions that are not members of any special classes. See Section 5.1.1 [Setting breakpoints], page 34.

catch throw
catch catch

Debug C++ exception handling using these commands. See Section 5.1.3 [Setting catchpoints], page 39.

ptype typename Print inheritance relationships as well as other information for type typename. See Chapter 11 [Examining the Symbol Table], page 115.

set print demangle
show print demangle
set print asm-demangle
show print asm-demangle

Control whether C++ symbols display in their source form, both when displaying code as C++ source and when displaying disassemblies. See Section 8.7 [Print settings], page 72.

set print object
show print object

Choose whether to print derived (actual) or declared types of objects. See Section 8.7 [Print settings], page 72.

set print vtbl
show print vtbl

Control the format for printing virtual function tables. See Section 8.7 [Print settings], page 72. (The vtbl commands do not work on programs compiled with the HP ANSI C++ compiler (aCC).)

set overload-resolution on

Enable overload resolution for C++ expression evaluation. The default is on. For overloaded functions, [No value for “GDBN”] evaluates the arguments and searches for a function whose signature matches the argument types, using the standard C++ conversion rules (see Section 10.4.1.3 [C++ expressions], page 100, for details). If it cannot find a match, it emits a message.

set overload-resolution off

Disable overload resolution for C++ expression evaluation. For overloaded functions that are not class member functions, [No value for “GDBN”] chooses the first function of the specified name that it finds in the symbol table, whether or not its arguments are of the correct type. For overloaded functions that are class member functions, [No value for “GDBN”] searches for a function whose signature exactly matches the argument types.

Overloaded symbol names You can specify a particular definition of an overloaded symbol, using the same notation that is used to declare such symbols in C++: type symbol(types) rather than just symbol. You can also use the [No value for “GDBN”] command-line word completion facilities to list the available choices, or to finish the type list for you. See Section 3.2 [Command completion], page 17, for details on how to do this.

10.4.2 Modula-2

The extensions made to [No value for “GDBN”] to support Modula-2 only support output from the gnu Modula-2 compiler (which is currently being developed). Other Modula-2 compilers are not currently supported, and attempting to debug executables produced by them is most likely to give an error as [No value for “GDBN”] reads in the executable’s symbol table.

10.4.2.1 Operators

Operators must be defined on values of specific types. For instance, + is defined on numbers, but not on structures. Operators are often defined on groups of types. For the purposes of Modula-2, the following definitions hold:

• Integral types consist of INTEGER, CARDINAL, and their subranges.

• Character types consist of CHAR and its subranges.

• Floating-point types consist of REAL.

• Pointer types consist of anything declared as POINTER TO type.

• Scalar types consist of all of the above.

• Set types consist of SET and BITSET types.

• Boolean types consist of BOOLEAN. The following operators are supported, and appear in order of increasing precedence: , Function argument or array index separator. := Assignment. The value of var := value is value. <, > Less than, greater than on integral, floating-point, or enumerated types. <=, >= Less than or equal to, greater than or equal to on integral, floating-point and

enumerated types, or set inclusion on set types. Same precedence as <.

=, <>, # Equality and two ways of expressing inequality, valid on scalar types. Same precedence as <. In [No value for “GDBN”] scripts, only <> is available for inequality, since # conflicts with the script comment character.

IN Set membership. Defined on set types and the types of their members. Same

precedence as <.

OR Boolean disjunction. Defined on boolean types. AND, & Boolean conjunction. Defined on boolean types. @ The [No value for “GDBN”] “artificial array” operator (see Section 8.1 [Expres-

sions], page 65).

+, - Addition and subtraction on integral and floating-point types, or union and

difference on set types.

* Multiplication on integral and floating-point types, or set intersection on set types.

/ Division on floating-point types, or symmetric set difference on set types. Same

precedence as *.

DIV, MOD Integer division and remainder. Defined on integral types. Same precedence as

          *.

- Negative. Defined on INTEGER and REAL data. ^Pointer dereferencing. Defined on pointer types. NOT Boolean negation. Defined on boolean types. Same precedence as ^.

. RECORD field selector. Defined on RECORD data. Same precedence as ^. [] Array indexing. Defined on ARRAY data. Same precedence as ^. () Procedure argument list. Defined on PROCEDURE objects. Same precedence as

          ^.

::, . [No value for “GDBN”] and Modula-2 scope operators.

Warning: Sets and their operations are not yet supported, so [No value for “GDBN”] treats the use of the operator IN, or the use of operators +, -, *, /, =, , <>, #, <=, and >= on sets as an error.

10.4.2.2 Built-in functions and procedures

Modula-2 also makes available several built-in procedures and functions. In describing these, the following metavariables are used: a represents an ARRAY variable. c represents a CHAR constant or variable. i represents a variable or constant of integral type. m represents an identifier that belongs to a set. Generally used in the same func-

tion with the metavariable s. The type of s should be SET OF mtype (where mtype is the type of m).

n represents a variable or constant of integral or floating-point type. r represents a variable or constant of floating-point type. t represents a type. v represents a variable. x represents a variable or constant of one of many types. See the explanation of

the function for details. All Modula-2 built-in procedures also return a result, described below.

ABS(n) Returns the absolute value of n. CAP(c) If c is a lower case letter, it returns its upper case equivalent, otherwise it

returns its argument.

CHR(i) Returns the character whose ordinal value is i. DEC(v) Decrements the value in the variable v by one. Returns the new value. DEC(v,i) Decrements the value in the variable v by i. Returns the new value.

EXCL(m,s)

Removes the element m from the set s. Returns the new set.

FLOAT(i) Returns the floating point equivalent of the integer i. HIGH(a) Returns the index of the last member of a. INC(v) Increments the value in the variable v by one. Returns the new value.

INC(v,i) Increments the value in the variable v by i. Returns the new value.

INCL(m,s)

Adds the element m to the set s if it is not already there. Returns the new set.

MAX(t) Returns the maximum value of the type t.

MIN(t) Returns the minimum value of the type t.

ODD(i) Returns boolean TRUE if i is an odd number.

ORD(x) Returns the ordinal value of its argument. For example, the ordinal value of a character is its ascii value (on machines supporting the ascii character set). x must be of an ordered type, which include integral, character and enumerated types.

SIZE(x) Returns the size of its argument. x can be a variable or a type.

TRUNC(r) Returns the integral part of r.

VAL(t,i) Returns the member of the type t whose ordinal value is i.

Warning: Sets and their operations are not yet supported, so [No value for “GDBN”] treats the use of procedures INCL and EXCL as an error.

10.4.2.3 Constants

[No value for “GDBN”] allows you to express the constants of Modula-2 in the following ways:

• Integer constants are simply a sequence of digits. When used in an expression, a constant is interpreted to be type-compatible with the rest of the expression. Hexadecimal integers are specified by a trailing ‘H’, and octal integers by a trailing ‘B’.

• Floating point constants appear as a sequence of digits, followed by a decimal point and another sequence of digits. An optional exponent can then be specified, in the form ‘E[+|-]nnn’, where ‘[+|-]nnn’ is the desired exponent. All of the digits of the floating point constant must be valid decimal (base 10) digits.

• Character constants consist of a single character enclosed by a pair of like quotes, either single (') or double ("). They may also be expressed by their ordinal value (their ascii value, usually) followed by a ‘C’.

• String constants consist of a sequence of characters enclosed by a pair of like quotes, either single (') or double ("). Escape sequences in the style of C are also allowed. See Section 10.4.1.2 [C and C++ constants], page 100, for a brief explanation of escape sequences.

• Enumerated constants consist of an enumerated identifier.

• Boolean constants consist of the identifiers TRUE and FALSE.

• Pointer constants consist of integral values only.

• Set constants are not yet supported.

10.4.2.4 Modula-2 defaults

If type and range checking are set automatically by [No value for “GDBN”], they both default to on whenever the working language changes to Modula-2. This happens regardless of whether you or [No value for “GDBN”] selected the working language.

If you allow [No value for “GDBN”] to set the language automatically, then entering code compiled from a file whose name ends with ‘.mod’ sets the working language to Modula-2. See Section 10.1.3 [Having [No value for “GDBN”] set the language automatically], page 94, for further details.

10.4.2.5 Deviations from standard Modula-2

A few changes have been made to make Modula-2 programs easier to debug. This is done primarily via loosening its type strictness:

• Unlike in standard Modula-2, pointer constants can be formed by integers. This allows you to modify pointer variables during debugging. (In standard Modula-2, the actual address contained in a pointer variable is hidden from you; it can only be modified through direct assignment to another pointer variable or expression that returned a pointer.)

• C escape sequences can be used in strings and characters to represent non-printable characters. [No value for “GDBN”] prints out strings with these escape sequences embedded. Single non-printable characters are printed using the ‘CHR(nnn)’ format.

• The assignment operator (:=) returns the value of its right-hand argument.

• All built-in procedures both modify and return their argument.

10.4.2.6 Modula-2 type and range checks

Warning: in this release, [No value for “GDBN”] does not yet perform type or range checking.

[No value for “GDBN”] considers two Modula-2 variables type equivalent if:

• They are of types that have been declared equivalent via a TYPE t1 = t2 statement • They have been declared on the same line. (Note: This is true of the gnu Modula-2

compiler, but it may not be true of other compilers.)

As long as type checking is enabled, any attempt to combine variables whose types are not equivalent is an error.

Range checking is done on all mathematical operations, assignment, array index bounds, and all built-in functions and procedures.

10.4.2.7 The scope operators :: and .

There are a few subtle differences between the Modula-2 scope operator (.) and the [No value for “GDBN”] scope operator (::). The two have similar syntax:

module . id scope :: id where scope is the name of a module or a procedure, module the name of a module, and id is any declared identifier within your program, except another module.

Using the :: operator makes [No value for “GDBN”] search the scope specified by scope for the identifier id. If it is not found in the specified scope, then [No value for “GDBN”] searches all scopes enclosing the one specified by scope.

Using the . operator makes [No value for “GDBN”] search the current scope for the identifier specified by id that was imported from the definition module specified by module. With this operator, it is an error if the identifier id was not imported from definition module module, or if id is not an identifier in module.

10.4.2.8 [No value for “GDBN”] and Modula-2

Some [No value for “GDBN”] commands have little use when debugging Modula-2 programs. Five subcommands of set print and show print apply specifically to C and C++: ‘vtbl’, ‘demangle’, ‘asm-demangle’, ‘object’, and ‘union’. The first four apply to C++, and the last to the C union type, which has no direct analogue in Modula-2.

The @ operator (see Section 8.1 [Expressions], page 65), while available with any language, is not useful with Modula-2. Its intent is to aid the debugging of dynamic arrays, which cannot be created in Modula-2 as they can in C or C++. However, because an address can be specified by an integral constant, the construct ‘{type}adrexp’ is still useful.

In [No value for “GDBN”] scripts, the Modula-2 inequality operator # is interpreted as the beginning of a comment. Use <> instead.

10.4.3 Chill

The extensions made to [No value for “GDBN”] to support Chill only support output from the gnu Chill compiler. Other Chill compilers are not currently supported, and attempting to debug executables produced by them is most likely to give an error as [No value for “GDBN”] reads in the executable’s symbol table.

This section covers the Chill related topics and the features of [No value for “GDBN”] which support these topics.

10.4.3.1 How modes are displayed

The Chill Datatype- (Mode) support of [No value for “GDBN”] is directly related with the functionality of the gnu Chill compiler, and therefore deviates slightly from the standard specification of the Chill language. The provided modes are:

Discrete modes:

• Integer Modes which are predefined by BYTE, UBYTE, INT, UINT, LONG,

              ULONG,

• Boolean Mode which is predefined by BOOL,

• Character Mode which is predefined by CHAR,

• Set Mode which is displayed by the keyword SET.

                   ([No value for ``GDBP'']) ptype x
                   type = SET (karli = 10, susi = 20, fritzi = 100)

If the type is an unnumbered set the set element values are omitted.

• Range Mode which is displayed by

         type = <basemode>(<lower bound> : <upper bound>)

where <lower bound>, <upper bound> can be of any discrete literal expression (e.g. set element names).

Powerset Mode:

A Powerset Mode is displayed by the keyword POWERSET followed by the member mode of the powerset. The member mode can be any discrete mode.

     ([No value for ``GDBP'']) ptype x
     type = POWERSET SET (egon, hugo, otto)

Reference Modes:

• Bound Reference Mode which is displayed by the keyword REF followed by the mode name to which the reference is bound.

• Free Reference Mode which is displayed by the keyword PTR.

Procedure mode

The procedure mode is displayed by type = PROC(<parameter list>) <return
mode> EXCEPTIONS (<exception list>). The <parameter list> is a list of

the parameter modes. <return mode> indicates the mode of the result of the procedure if any. The exceptionlist lists all possible exceptions which can be raised by the procedure.

Synchronization Modes:

• Event Mode which is displayed by

         EVENT (<event length>)

where (<event length>) is optional.

• Buffer Mode which is displayed by

         BUFFER (<buffer length>)<buffer element mode>
    where (<buffer length>) is optional.

Timing Modes:

• Duration Mode which is predefined by DURATION

• Absolute Time Mode which is predefined by TIME

Real Modes:

Real Modes are predefined with REAL and LONG_REAL.

String Modes:

• Character String Mode which is displayed by

         CHARS(<string length>)

followed by the keyword VARYING if the String Mode is a varying mode • Bit String Mode which is displayed by

         BOOLS(<string
         length>)

Array Mode:

The Array Mode is displayed by the keyword ARRAY(<range>) followed by the element mode (which may in turn be an array mode).

               ([No value for ``GDBP'']) ptype x
               type = ARRAY (1:42)
                         ARRAY (1:20)
                            SET (karli = 10, susi = 20, fritzi = 100)

Structure Mode The Structure mode is displayed by the keyword STRUCT(<field list>). The <field list> consists of names and modes of fields of the structure. Variant

          structures have the keyword CASE <field> OF <variant fields> ESAC in their

field list. Since the current version of the GNU Chill compiler doesn’t implement tag processing (no runtime checks of variant fields, and therefore no debugging info), the output always displays all variant fields.

               ([No value for ``GDBP'']) ptype str
               type = STRUCT (
                   as x,
                   bs x,
                   CASE bs OF
                   (karli):
                       cs a
                   (ott):
                       ds x
                   ESAC
               )

10.4.3.2 Locations and their accesses

A location in Chill is an object which can contain values.

A value of a location is generally accessed by the (declared) name of the location. The output conforms to the specification of values in Chill programs. How values are specified is the topic of the next section, Section 10.4.3.3 [Values and their Operations], page 111.

The pseudo-location RESULT (or result) can be used to display or change the result of a currently-active procedure:

     set result := EXPR

This does the same as the Chill action RESULT EXPR (which is not available in [No value for “GDBN”]).

Values of reference mode locations are printed by PTR(<hex value>) in case of a free reference mode, and by (REF <reference mode>) (<hex-value>) in case of a bound reference. <hex value> represents the address where the reference points to. To access the value of the location referenced by the pointer, use the dereference operator ‘->’.

Values of procedure mode locations are displayed by

     { PROC
     (<argument modes> ) <return mode> } <address> <name of procedure
     location>

<argument modes> is a list of modes according to the parameter specification of the procedure and <address> shows the address of the entry point.

Substructures of string mode-, array mode- or structure mode-values (e.g. array slices, fields of structure locations) are accessed using certain operations which are described in the next section, Section 10.4.3.3 [Values and their Operations], page 111.

A location value may be interpreted as having a different mode using the location conversion. This mode conversion is written as <mode name>(<location>). The user has to consider that the sizes of the modes have to be equal otherwise an error occurs. Furthermore, no range checking of the location against the destination mode is performed, and therefore the result can be quite confusing.

     ([No value for ``GDBP'']) print int (s(3 up 4)) XXX TO be filled in !! XXX

10.4.3.3 Values and their Operations

Values are used to alter locations, to investigate complex structures in more detail or to filter relevant information out of a large amount of data. There are several (mode dependent) operations defined which enable such investigations. These operations are not only applicable to constant values but also to locations, which can become quite useful when debugging complex structures. During parsing the command line (e.g. evaluating an expression) [No value for “GDBN”] treats location names as the values behind these locations.

This section describes how values have to be specified and which operations are legal to be used with such values.

Literal Values

Literal values are specified in the same manner as in gnu Chill programs. For detailed specification refer to the gnu Chill implementation Manual chapter 1.5.

Tuple Values

A tuple is specified by <mode name>[<tuple>], where <mode name> can be omitted if the mode of the tuple is unambiguous. This unambiguity is derived from the context of a evaluated expression. <tuple> can be one of the following:

• Powerset Tuple

• Array Tuple

• Structure Tuple Powerset tuples, array tuples and structure tuples are specified in the same manner as in Chill programs refer to z200/88 chpt 5.2.5.

String Element Value

A string element value is specified by

               <string value>(<index>)

where <index> is a integer expression. It delivers a character value which is equivalent to the character indexed by <index> in the string.

String Slice Value

A string slice value is specified by <string value>(<slice spec>), where <slice spec> can be either a range of integer expressions or specified by <start expr> up <size>. <size> denotes the number of elements which the slice contains. The delivered value is a string value, which is part of the specified string.

Array Element Values

An array element value is specified by <array value>(<expr>) and delivers a array element value of the mode of the specified array.

Array Slice Values
An array slice is specified by <array value>(<slice spec>), where <slice

spec> can be either a range specified by expressions or by <start expr> up <size>. <size> denotes the number of arrayelements the slice contains. The delivered value is an array value which is part of the specified array.

Structure Field Values

A structure field value is derived by <structure value>.<field name>, where <field name> indicates the name of a field specified in the mode definition of the structure. The mode of the delivered value corresponds to this mode definition in the structure definition.

Procedure Call Value

The procedure call value is derived from the return value of the procedure^{1}. Values of duration mode locations are represented by ULONG literals. Values of time mode locations appear as

     TIME(<secs>:<nsecs>)
Zero-adic Operator Value

The zero-adic operator value is derived from the instance value for the current active process.

Expression Values

The value delivered by an expression is the result of the evaluation of the specified expression. If there are error conditions (mode incompatibility, etc.) the evaluation of expressions is aborted with a corresponding error message. Expressions may be parenthesised which causes the evaluation of this expression before any other expression which uses the result of the parenthesised expression. The following operators are supported by [No value for “GDBN”]:

OR, ORIF, XOR
AND, ANDIF

NOT Logical operators defined over operands of boolean mode.

=, /= Equality and inequality operators defined over all modes.

>, >=

<, <= Relational operators defined over predefined modes.

+, -
*, /, MOD, REM

Arithmetic operators defined over predefined modes.

- Change sign operator.

// String concatenation operator.

1 If a procedure call is used for instance in an expression, then this procedure is called with all its side

effects. This can lead to confusing results if used carelessly.

() String repetition operator.

-> Referenced location operator which can be used either to take the address of a location (->loc), or to dereference a reference location

                    (loc->).
          OR, XOR
          AND

NOT Powerset and bitstring operators.

          >, >=

<, <= Powerset inclusion operators.

IN Membership operator.

10.4.3.4 Chill type and range checks

[No value for “GDBN”] considers two Chill variables mode equivalent if the sizes of the two modes are equal. This rule applies recursively to more complex datatypes which means that complex modes are treated equivalent if all element modes (which also can be complex modes like structures, arrays, etc.) have the same size.

Range checking is done on all mathematical operations, assignment, array index bounds and all built in procedures.

Strong type checks are forced using the [No value for “GDBN”] command set check strong. This enforces strong type and range checks on all operations where Chill constructs are used (expressions, built in functions, etc.) in respect to the semantics as defined in the z.200 language specification.

All checks can be disabled by the [No value for “GDBN”] command set check off.

10.4.3.5 Chill defaults

If type and range checking are set automatically by [No value for “GDBN”], they both default to on whenever the working language changes to Chill. This happens regardless of whether you or [No value for “GDBN”] selected the working language.

If you allow [No value for “GDBN”] to set the language automatically, then entering code compiled from a file whose name ends with ‘.ch’ sets the working language to Chill. See Section 10.1.3 [Having [No value for “GDBN”] set the language automatically], page 94, for further details.

114 Debugging with [No value for “GDBN”]

11 Examining the Symbol Table

The commands described in this chapter allow you to inquire about the symbols (names of variables, functions and types) defined in your program. This information is inherent in the text of your program and does not change as your program executes. [No value for “GDBN”] finds it in your program’s symbol table, in the file indicated when you started [No value for “GDBN”] (see Section 2.1.1 [Choosing files], page 12), or by one of the filemanagement commands (see Section 13.1 [Commands to specify files], page 123).

Occasionally, you may need to refer to symbols that contain unusual characters, which [No value for “GDBN”] ordinarily treats as word delimiters. The most frequent case is in referring to static variables in other source files (see Section 8.2 [Program variables], page 66). File names are recorded in object files as debugging symbols, but [No value for “GDBN”] would ordinarily parse a typical file name, like ‘foo.c’, as the three words ‘foo’ ‘.’ ‘c’. To allow [No value for “GDBN”] to recognize ‘foo.c’ as a single symbol, enclose it in single quotes; for example,

     p 'foo.c'::x

looks up the value of x in the scope of the file ‘foo.c’.

info address symbol

Describe where the data for symbol is stored. For a register variable, this says which register it is kept in. For a non-register local variable, this prints the stack-frame offset at which the variable is always stored. Note the contrast with ‘print &symbol’, which does not work at all for a register variable, and for a stack local variable prints the exact address of the current instantiation of the variable.

info symbol addr

Print the name of a symbol which is stored at the address addr. If no symbol is stored exactly at addr, [No value for “GDBN”] prints the nearest symbol and an offset from it:

               ([No value for ``GDBP'']) info symbol 0x54320
               _initialize_vx + 396 in section .text

This is the opposite of the info address command. You can use it to find out the name of a variable or a function given its address.

whatis expr

Print the data type of expression expr. expr is not actually evaluated, and any side-effecting operations (such as assignments or function calls) inside it do not take place. See Section 8.1 [Expressions], page 65.

whatis Print the data type of $, the last value in the value history.

ptype typename Print a description of data type typename. typename may be the name of a type, or for C code it may have the form ‘class class-name’, ‘struct structtag’, ‘union union-tag’ or ‘enum enum-tag’.

ptype expr

ptype Print a description of the type of expression expr. ptype differs from whatis by printing a detailed description, instead of just the name of the type.

For example, for this variable declaration:

     struct complex {double real; double imag;} v;

the two commands give this output:

     ([No value for ``GDBP'']) whatis v
     type = struct complex
     ([No value for ``GDBP'']) ptype v
     type = struct complex {
         double real;
         double imag;
     }

As with whatis, using ptype without an argument refers to the type of $, the last value in the value history.

info types regexp
info types

Print a brief description of all types whose names match regexp (or all types in your program, if you supply no argument). Each complete typename is matched as though it were a complete line; thus, ‘i type value’ gives information on all types in your program whose names include the string value, but ‘i type ^value$’ gives information only on types whose complete name is value.

This command differs from ptype in two ways: first, like whatis, it does not print a detailed description; second, it lists all source files where a type is defined.

info scope addr

List all the variables local to a particular scope. This command accepts a location—a function name, a source line, or an address preceded by a ‘*’, and prints all the variables local to the scope defined by that location. For example:

([No value for ``GDBP'']) info scope command line handler

     Scope for command_line_handler:
     Symbol rl is an argument at stack/frame offset 8, length 4.
     Symbol linebuffer is in static storage at address 0x150a18, length 4.
     Symbol linelength is in static storage at address 0x150a1c, length 4.
     Symbol p is a local variable in register $esi, length 4.
     Symbol p1 is a local variable in register $ebx, length 4.
     Symbol nline is a local variable in register $edx, length 4.
     Symbol repeat is a local variable at frame offset -8, length 4.

This command is especially useful for determining what data to collect during a trace experiment, see Section 9.1.4 [Tracepoint Actions], page 85.

info source

Show the name of the current source file—that is, the source file for the function containing the current point of execution—and the language it was written in.

info sources

Print the names of all source files in your program for which there is debugging information, organized into two lists: files whose symbols have already been read, and files whose symbols will be read when needed.

info functions

Print the names and data types of all defined functions.

info functions regexp

Print the names and data types of all defined functions whose names contain a match for regular expression regexp. Thus, ‘info fun step’ finds all functions whose names include step; ‘info fun ^step’ finds those whose names start with step.

info variables

Print the names and data types of all variables that are declared outside of functions (i.e., excluding local variables).

info variables regexp

Print the names and data types of all variables (except for local variables) whose names contain a match for regular expression regexp. Some systems allow individual object files that make up your program to be replaced without stopping and restarting your program. For example, in Vx-Works you can simply recompile a defective object file and keep on running. If you are running on one of these systems, you can allow [No value for “GDBN”] to reload the symbols for automatically relinked modules:

set symbol-reloading on

Replace symbol definitions for the corresponding source file when an object file with a particular name is seen again.

set symbol-reloading off

Do not replace symbol definitions when encountering object files of the same name more than once. This is the default state; if you are not running on a system that permits automatic relinking of modules, you should leave symbol-reloading off, since otherwise [No value for “GDBN”] may discard symbols when linking large programs, that may contain several modules (from different directories or libraries) with the same name.

show symbol-reloading

Show the current on or off setting.

set opaque-type-resolution on

Tell [No value for “GDBN”] to resolve opaque types. An opaque type is a type declared as a pointer to a struct, class, or union—for example, struct MyType *—that is used in one source file although the full declaration of struct MyType is in another source file. The default is on. A change in the setting of this subcommand will not take effect until the next time symbols for a file are loaded.

set opaque-type-resolution off

Tell [No value for “GDBN”] not to resolve opaque types. In this case, the type is printed as follows:

     {<no data fields>}
show opaque-type-resolution

Show whether opaque types are resolved or not.

maint print symbols filename
maint print psymbols filename
maint print msymbols filename

Write a dump of debugging symbol data into the file filename. These commands are used to debug the [No value for “GDBN”] symbol-reading code. Only symbols with debugging data are included. If you use ‘maint print symbols’, [No value for “GDBN”] includes all the symbols for which it has already collected full details: that is, filename reflects symbols for only those files whose symbols [No value for “GDBN”] has read. You can use the command info sources to find out which files these are. If you use ‘maint print psymbols’ instead, the dump shows information about symbols that [No value for “GDBN”] only knows partially—that is, symbols defined in files that [No value for “GDBN”] has skimmed, but not yet read completely. Finally, ‘maint print msymbols’ dumps just the minimal symbol information required for each object file from which [No value for “GDBN”] has read some symbols. See Section 13.1 [Commands to specify files], page 123, for a discussion of how [No value for “GDBN”] reads symbols (in the description of symbol-file).

12 Altering Execution

Once you think you have found an error in your program, you might want to find out for certain whether correcting the apparent error would lead to correct results in the rest of the run. You can find the answer by experiment, using the [No value for “GDBN”] features for altering execution of the program.

For example, you can store new values into variables or memory locations, give your program a signal, restart it at a different address, or even return prematurely from a function.

12.1 Assignment to variables

To alter the value of a variable, evaluate an assignment expression. See Section 8.1 [Expressions], page 65. For example,

     print x=4

stores the value 4 into the variable x, and then prints the value of the assignment expression (which is 4). See Chapter 10 [Using [No value for “GDBN”] with Different Languages], page 93, for more information on operators in supported languages.

If you are not interested in seeing the value of the assignment, use the set command instead of the print command. set is really the same as print except that the expression’s value is not printed and is not put in the value history (see Section 8.8 [Value history], page 77). The expression is evaluated only for its effects.

If the beginning of the argument string of the set command appears identical to a set subcommand, use the set variable command instead of just set. This command is identical to set except for its lack of subcommands. For example, if your program has a variable width, you get an error if you try to set a new value with just ‘set width=13’, because [No value for “GDBN”] has the command set width:

     ([No value for ``GDBP'']) whatis width
     type = double
     ([No value for ``GDBP'']) p width
     $4 = 13
     ([No value for ``GDBP'']) set width=47
     Invalid syntax in expression.

The invalid expression, of course, is ‘=47’. In order to actually set the program’s variable

width, use
     ([No value for ``GDBP'']) set var width=47

Because the set command has many subcommands that can conflict with the names of program variables, it is a good idea to use the set variable command instead of just set. For example, if your program has a variable g, you run into problems if you try to set a new value with just ‘set g=4’, because [No value for “GDBN”] has the command set

gnutarget, abbreviated set g:

([No value for ``GDBP'']) whatis g
type = double
([No value for ``GDBP'']) p g
$1 = 1
([No value for ``GDBP'']) set g=4
([No value for ``GDBP'']) p g
$2 = 1
([No value for ``GDBP'']) r
The program being debugged has been started already.
Start it from the beginning? (y or n) y
Starting program: /home/smith/cc_progs/a.out
"/home/smith/cc_progs/a.out": can't open to read symbols:
                                 Invalid bfd target.
([No value for ``GDBP'']) show g
The current BFD target is "=4".

The program variable g did not change, and you silently set the gnutarget to an invalid value. In order to set the variable g, use

([No value for ``GDBP'']) set var g=4

[No value for “GDBN”] allows more implicit conversions in assignments than C; you can freely store an integer value into a pointer variable or vice versa, and you can convert any structure to any other structure that is the same length or shorter.

To store values into arbitrary places in memory, use the ‘{...}’ construct to generate a value of specified type at a specified address (see Section 8.1 [Expressions], page 65). For example, {int}0x83040 refers to memory location 0x83040 as an integer (which implies a certain size and representation in memory), and

set {int}0x83040 = 4

stores the value 4 into that memory location.

12.2 Continuing at a different address

Ordinarily, when you continue your program, you do so at the place where it stopped, with the continue command. You can instead continue at an address of your own choosing, with the following commands:

jump linespec Resume execution at line linespec. Execution stops again immediately if there is a breakpoint there. See Section 7.1 [Printing source lines], page 59, for a description of the different forms of linespec. It is common practice to use the tbreak command in conjunction with jump. See Section 5.1.1 [Setting breakpoints], page 34. The jump command does not change the current stack frame, or the stack pointer, or the contents of any memory location or any register other than the program counter. If line linespec is in a different function from the one currently executing, the results may be bizarre if the two functions expect different patterns of arguments or of local variables. For this reason, the jump command requests confirmation if the specified line is not in the function currently executing. However, even bizarre results are predictable if you are well acquainted with the machine-language code of your program.

jump *address Resume execution at the instruction at address address.

On many systems, you can get much the same effect as the jump command by storing a new value into the register $pc. The difference is that this does not start your program running; it only changes the address of where it will run when you continue. For example,

     set $pc = 0x485

makes the next continue command or stepping command execute at address 0x485, rather than at the address where your program stopped. See Section 5.2 [Continuing and stepping], page 46.

The most common occasion to use the jump command is to back up—perhaps with more breakpoints set—over a portion of a program that has already executed, in order to examine its execution in more detail.

12.3 Giving your program a signal

signal signal

Resume execution where your program stopped, but immediately give it the signal signal. signal can be the name or the number of a signal. For example, on many systems signal 2 and signal SIGINT are both ways of sending an interrupt signal. Alternatively, if signal is zero, continue execution without giving a signal. This is useful when your program stopped on account of a signal and would ordinary see the signal when resumed with the continue command; ‘signal 0’ causes it to resume without a signal. signal does not repeat when you press ^{⟨}\underline{\overline{RET}}^{⟩} a second time after executing the command.

Invoking the signal command is not the same as invoking the kill utility from the shell. Sending a signal with kill causes [No value for “GDBN”] to decide what to do with the signal depending on the signal handling tables (see Section 5.3 [Signals], page 49). The signal command passes the signal directly to your program.

12.4 Returning from a function

return

return expression You can cancel execution of a function call with the return command. If you give an expression argument, its value is used as the function’s return value.

When you use return, [No value for “GDBN”] discards the selected stack frame (and all frames within it). You can think of this as making the discarded frame return prematurely. If you wish to specify a value to be returned, give that value as the argument to return.

This pops the selected stack frame (see Section 6.3 [Selecting a frame], page 55), and any other frames inside of it, leaving its caller as the innermost remaining frame. That frame becomes selected. The specified value is stored in the registers used for returning values of functions.

The return command does not resume execution; it leaves the program stopped in the state that would exist if the function had just returned. In contrast, the finish command (see Section 5.2 [Continuing and stepping], page 46) resumes execution until the selected stack frame returns naturally.

12.5 Calling program functions

call expr Evaluate the expression expr without displaying void returned values.

You can use this variant of the print command if you want to execute a function from your program, but without cluttering the output with void returned values. If the result is not void, it is printed and saved in the value history.

For the A29K, a user-controlled variable call_scratch_address, specifies the location of a scratch area to be used when [No value for “GDBN”] calls a function in the target. This is necessary because the usual method of putting the scratch area on the stack does not work in systems that have separate instruction and data spaces.

12.6 Patching programs

By default, [No value for “GDBN”] opens the file containing your program’s executable code (or the corefile) read-only. This prevents accidental alterations to machine code; but it also prevents you from intentionally patching your program’s binary.

If you’d like to be able to patch the binary, you can specify that explicitly with the set write command. For example, you might want to turn on internal debugging flags, or even to make emergency repairs.

set write on
set write off

If you specify ‘set write on’, [No value for “GDBN”] opens executable and core files for both reading and writing; if you specify ‘set write off’ (the default), [No value for “GDBN”] opens them read-only. If you have already loaded a file, you must load it again (using the exec-file or core-file command) after changing set write, for your new setting to take effect.

show write

Display whether executable files and core files are opened for writing as well as reading.

13 [No value for “GDBN”] Files

[No value for “GDBN”] needs to know the file name of the program to be debugged, both in order to read its symbol table and in order to start your program. To debug a core dump of a previous run, you must also tell [No value for “GDBN”] the name of the core dump file.

13.1 Commands to specify files

You may want to specify executable and core dump file names. The usual way to do this is at start-up time, using the arguments to [No value for “GDBN”]’s start-up commands (see Chapter 2 [Getting In and Out of [No value for “GDBN”]], page 11).

Occasionally it is necessary to change to a different file during a [No value for “GDBN”] session. Or you may run [No value for “GDBN”] and forget to specify a file you want to use. In these situations the [No value for “GDBN”] commands to specify new files are useful.

file filename Use filename as the program to be debugged. It is read for its symbols and for the contents of pure memory. It is also the program executed when you use the run command. If you do not specify a directory and the file is not found in the [No value for “GDBN”] working directory, [No value for “GDBN”] uses the environment variable PATH as a list of directories to search, just as the shell does when looking for a program to run. You can change the value of this variable, for both [No value for “GDBN”] and your program, using the path command. On systems with memory-mapped files, an auxiliary file named ‘filename.syms’ may hold symbol table information for filename. If so, [No value for “GDBN”] maps in the symbol table from ‘filename.syms’, starting up more quickly. See the descriptions of the file options ‘-mapped’ and ‘-readnow’ (available on the command line, and with the commands file, symbol-file, or add-symbol-file, described below), for more information.

file file with no argument makes [No value for “GDBN”] discard any information it has on both executable file and the symbol table.

exec-file [ filename ]

Specify that the program to be run (but not the symbol table) is found in filename. [No value for “GDBN”] searches the environment variable PATH if necessary to locate your program. Omitting filename means to discard information on the executable file.

symbol-file [ filename ]

Read symbol table information from file filename. PATH is searched when necessary. Use the file command to get both symbol table and program to run from the same file. symbol-file with no argument clears out [No value for “GDBN”] information on your program’s symbol table. The symbol-file command causes [No value for “GDBN”] to forget the contents of its convenience variables, the value history, and all breakpoints and auto-display expressions. This is because they may contain pointers to the internal data recording symbols and data types, which are part of the old symbol table data being discarded inside [No value for “GDBN”]. symbol-file does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} again after executing it once. When [No value for “GDBN”] is configured for a particular environment, it understands debugging information in whatever format is the standard generated for that environment; you may use either a gnu compiler, or other compilers that adhere to the local conventions. Best results are usually obtained from gnu compilers; for example, using [No value for ``GCC''] you can generate debugging information for optimized code. For most kinds of object files, with the exception of old SVR3 systems using COFF, the symbol-file command does not normally read the symbol table in full right away. Instead, it scans the symbol table quickly to find which source files and which symbols are present. The details are read later, one source file at a time, as they are needed. The purpose of this two-stage reading strategy is to make [No value for “GDBN”] start up faster. For the most part, it is invisible except for occasional pauses while the symbol table details for a particular source file are being read. (The set verbose command can turn these pauses into messages if desired. See Section 16.6 [Optional warnings and messages], page 174.) We have not implemented the two-stage strategy for COFF yet. When the symbol table is stored in COFF format, symbol-file reads the symbol table data in full right away. Note that “stabs-in-COFF” still does the two-stage strategy, since the debug info is actually in stabs format.

symbol-file filename [ -readnow ] [ -mapped ]
file filename [ -readnow ] [ -mapped ]

You can override the [No value for “GDBN”] two-stage strategy for reading symbol tables by using the ‘-readnow’ option with any of the commands that load symbol table information, if you want to be sure [No value for “GDBN”] has the entire symbol table available. If memory-mapped files are available on your system through the mmap system call, you can use another option, ‘-mapped’, to cause [No value for “GDBN”] to write the symbols for your program into a reusable file. Future [No value for “GDBN”] debugging sessions map in symbol information from this auxiliary symbol file (if the program has not changed), rather than spending time reading the symbol table from the executable program. Using the ‘-mapped’ option has the same effect as starting [No value for “GDBN”] with the ‘-mapped’ commandline option. You can use both options together, to make sure the auxiliary symbol file has all the symbol information for your program. The auxiliary symbol file for a program called myprog is called ‘myprog.syms’. Once this file exists (so long as it is newer than the corresponding executable), [No value for “GDBN”] always attempts to use it when you debug myprog; no special options or commands are needed.

The ‘.syms’ file is specific to the host machine where you run [No value for “GDBN”]. It holds an exact image of the internal [No value for “GDBN”] symbol table. It cannot be shared across multiple host platforms.

core-file [ filename ]

Specify the whereabouts of a core dump file to be used as the “contents of memory”. Traditionally, core files contain only some parts of the address space of the process that generated them; [No value for “GDBN”] can access the executable file itself for other parts.

core-file with no argument specifies that no core file is to be used.

Note that the core file is ignored when your program is actually running under [No value for “GDBN”]. So, if you have been running your program and you wish to debug a core file instead, you must kill the subprocess in which the program is running. To do this, use the kill command (see Section 4.8 [Killing the child process], page 28).

add-symbol-file filename address
add-symbol-file filename address [ -readnow ] [ -mapped ]

add-symbol-file filename -ssection address The add-symbol-file command reads additional symbol table information from the file filename. You would use this command when filename has been dynamically loaded (by some other means) into the program that is running. address should be the memory address at which the file has been loaded; [No value for “GDBN”] cannot figure this out for itself. You can additionally specify an arbitrary number of ‘-ssection address’ pairs, to give an explicit section name and base address for that section. You can specify any address as an expression.

The symbol table of the file filename is added to the symbol table originally read with the symbol-file command. You can use the add-symbol-file command any number of times; the new symbol data thus read keeps adding to the old. To discard all old symbol data instead, use the symbol-file command without any arguments.

add-symbol-file does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} after using it.

You can use the ‘-mapped’ and ‘-readnow’ options just as with the symbolfile command, to change how [No value for “GDBN”] manages the symbol table information for filename.

add-shared-symbol-file

The add-shared-symbol-file command can be used only under Harris’ CXUX operating system for the Motorola 88k. [No value for “GDBN”] automatically looks for shared libraries, however if [No value for “GDBN”] does not find yours, you can run add-shared-symbol-file. It takes no arguments.

section The section command changes the base address of section SECTION of the exec file to ADDR. This can be used if the exec file does not contain section addresses, (such as in the a.out format), or when the addresses specified in the file itself are wrong. Each section must be changed separately. The info files command, described below, lists all the sections and their addresses.

info files
info target

info files and info target are synonymous; both print the current target (see Chapter 14 [Specifying a Debugging Target], page 129), including the names of the executable and core dump files currently in use by [No value for “GDBN”], and the files from which symbols were loaded. The command help target lists all possible targets rather than current ones.

All file-specifying commands allow both absolute and relative file names as arguments. [No value for “GDBN”] always converts the file name to an absolute file name and remembers it that way.

[No value for “GDBN”] supports HP-UX, SunOS, SVr4, Irix 5, and IBM RS/6000 shared libraries.

[No value for “GDBN”] automatically loads symbol definitions from shared libraries when you use the run command, or when you examine a core file. (Before you issue the run command, [No value for “GDBN”] does not understand references to a function in a shared library, however—unless you are debugging a core file).

On HP-UX, if the program loads a library explicitly, [No value for “GDBN”] automatically loads the symbols at the time of the shl_load call.

info share
info sharedlibrary

Print the names of the shared libraries which are currently loaded.

sharedlibrary regex
share regex

Load shared object library symbols for files matching a Unix regular expression. As with files loaded automatically, it only loads shared libraries required by your program for a core file or after typing run. If regex is omitted all shared libraries required by your program are loaded.

On HP-UX systems, [No value for “GDBN”] detects the loading of a shared library and automatically reads in symbols from the newly loaded library, up to a threshold that is initially set but that you can modify if you wish.

Beyond that threshold, symbols from shared libraries must be explicitly loaded. To load these symbols, use the command sharedlibrary filename. The base address of the shared library is determined automatically by [No value for “GDBN”] and need not be specified.

To display or set the threshold, use the commands:

set auto-solib-add threshold

Set the autoloading size threshold, in megabytes. If threshold is nonzero, symbols from all shared object libraries will be loaded automatically when the inferior begins execution or when the dynamic linker informs [No value for “GDBN”] that a new library has been loaded, until the symbol table of the program and libraries exceeds this threshold. Otherwise, symbols must be loaded manually, using the sharedlibrary command. The default threshold is 100 megabytes.

show auto-solib-add

Display the current autoloading size threshold, in megabytes.

13.2 Errors reading symbol files

While reading a symbol file, [No value for “GDBN”] occasionally encounters problems, such as symbol types it does not recognize, or known bugs in compiler output. By default, [No value for “GDBN”] does not notify you of such problems, since they are relatively common and primarily of interest to people debugging compilers. If you are interested in seeing information about ill-constructed symbol tables, you can either ask [No value for “GDBN”] to print only one message about each such type of problem, no matter how many times the problem occurs; or you can ask [No value for “GDBN”] to print more messages, to see how many times the problems occur, with the set complaints command (see Section 16.6 [Optional warnings and messages], page 174).

The messages currently printed, and their meanings, include:

inner block not inside outer block in symbol

The symbol information shows where symbol scopes begin and end (such as at the start of a function or a block of statements). This error indicates that an inner scope block is not fully contained in its outer scope blocks.

[No value for “GDBN”] circumvents the problem by treating the inner block as if it had the same scope as the outer block. In the error message, symbol may be shown as “(don't know)” if the outer block is not a function.

block at address out of order

The symbol information for symbol scope blocks should occur in order of increasing addresses. This error indicates that it does not do so.

[No value for “GDBN”] does not circumvent this problem, and has trouble locating symbols in the source file whose symbols it is reading. (You can often determine what source file is affected by specifying set verbose on. See Section 16.6 [Optional warnings and messages], page 174.)

bad block start address patched

The symbol information for a symbol scope block has a start address smaller than the address of the preceding source line. This is known to occur in the SunOS 4.1.1 (and earlier) C compiler.

[No value for “GDBN”] circumvents the problem by treating the symbol scope block as starting on the previous source line.

bad string table offset in symbol n

Symbol number n contains a pointer into the string table which is larger than the size of the string table.

[No value for “GDBN”] circumvents the problem by considering the symbol to have the name foo, which may cause other problems if many symbols end up with this name.

unknown symbol type 0xnn

The symbol information contains new data types that [No value for “GDBN”] does not yet know how to read. 0xnn is the symbol type of the uncomprehended information, in hexadecimal. [No value for “GDBN”] circumvents the error by ignoring this symbol information. This usually allows you to debug your program, though certain symbols are not accessible. If you encounter such a problem and feel like debugging it, you can debug [No value for ``GDBP''] with itself, breakpoint on complain, then go up to the function read_dbx_symtab and examine *bufp to see the symbol.

stub type has NULL name

[No value for “GDBN”] could not find the full definition for a struct or class.

const/volatile indicator missing (ok if using g++ v1.x), got...

The symbol information for a C++ member function is missing some information that recent versions of the compiler should have output for it.

info mismatch between compiler and debugger

[No value for “GDBN”] could not parse a type specification output by the compiler.

14 Specifying a Debugging Target

A target is the execution environment occupied by your program.

Often, [No value for “GDBN”] runs in the same host environment as your program; in that case, the debugging target is specified as a side effect when you use the file or core commands. When you need more flexibility—for example, running [No value for “GDBN”] on a physically separate host, or controlling a standalone system over a serial port or a realtime system over a TCP/IP connection—you can use the target command to specify one of the target types configured for [No value for “GDBN”] (see Section 14.2 [Commands for managing targets], page 129).

14.1 Active targets

There are three classes of targets: processes, core files, and executable files. [No value for “GDBN”] can work concurrently on up to three active targets, one in each class. This allows you to (for example) start a process and inspect its activity without abandoning your work on a core file.

For example, if you execute ‘gdb a.out’, then the executable file a.out is the only active target. If you designate a core file as well—presumably from a prior run that crashed and coredumped—then [No value for “GDBN”] has two active targets and uses them in tandem, looking first in the corefile target, then in the executable file, to satisfy requests for memory addresses. (Typically, these two classes of target are complementary, since core files contain only a program’s read-write memory—variables and so on—plus machine status, while executable files contain only the program text and initialized data.)

When you type run, your executable file becomes an active process target as well. When a process target is active, all [No value for “GDBN”] commands requesting memory addresses refer to that target; addresses in an active core file or executable file target are obscured while the process target is active.

Use the core-file and exec-file commands to select a new core file or executable target (see Section 13.1 [Commands to specify files], page 123). To specify as a target a process that is already running, use the attach command (see Section 4.7 [Debugging an already-running process], page 27).

14.2 Commands for managing targets

target type parameters Connects the [No value for “GDBN”] host environment to a target machine or process. A target is typically a protocol for talking to debugging facilities. You use the argument type to specify the type or protocol of the target machine. Further parameters are interpreted by the target protocol, but typically include things like device names or host names to connect with, process numbers, and baud rates. The target command does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} again after executing the command.

help target

Displays the names of all targets available. To display targets currently selected, use either info target or info files (see Section 13.1 [Commands to specify files], page 123).

help target name

Describe a particular target, including any parameters necessary to select it.

set gnutarget args

[No value for “GDBN”] uses its own library BFD to read your files. [No value for “GDBN”] knows whether it is reading an executable, a core, or a .o file; however, you can specify the file format with the set gnutarget command. Unlike most target commands, with gnutarget the target refers to a program, not a machine.

Warning: To specify a file format with set gnutarget, you must know the actual BFD name.

See Section 13.1 [Commands to specify files], page 123.

show gnutarget

Use the show gnutarget command to display what file format gnutarget is set to read. If you have not set gnutarget, [No value for “GDBN”] will determine the file format for each file automatically, and show gnutarget displays ‘The

current BDF target is "auto"’.

Here are some common targets (available, or not, depending on the GDB configuration):

target exec program

An executable file. ‘target exec program’ is the same as ‘exec-file program’.

target core filename

A core dump file. ‘target core filename’ is the same as ‘core-file filename’.

target remote dev

Remote serial target in GDB-specific protocol. The argument dev specifies what serial device to use for the connection (e.g. ‘/dev/ttya’). See Section 14.4 [Remote debugging], page 132. target remote supports the load command. This is only useful if you have some other way of getting the stub to the target system, and you can put it somewhere in memory where it won’t get clobbered by the download.

target sim

Builtin CPU simulator. [No value for “GDBN”] includes simulators for most architectures. In general,

             target sim
             load
             run

works; however, you cannot assume that a specific memory map, device drivers, or even basic I/O is available, although some simulators do provide these. For info about any processor-specific simulator details, see the appropriate section in Section 15.3 [Embedded Processors], page 155.

Some configurations may include these targets as well:

target nrom dev

NetROM ROM emulator. This target only supports downloading.

Different targets are available on different configurations of [No value for “GDBN”]; your configuration may have more or fewer targets.

Many remote targets require you to download the executable’s code once you’ve successfully established a connection.

load filename Depending on what remote debugging facilities are configured into [No value for “GDBN”], the load command may be available. Where it exists, it is meant to make filename (an executable) available for debugging on the remote system—by downloading, or dynamic linking, for example. load also records the filename symbol table in [No value for “GDBN”], like the add-symbol-file command. If your [No value for “GDBN”] does not have a load command, attempting to execute it gets the error message “You can't do that when your target is

...”

The file is loaded at whatever address is specified in the executable. For some object file formats, you can specify the load address when you link the program; for other formats, like a.out, the object file format specifies a fixed address. load does not repeat if you press ^{⟨}\underline{\overline{RET}}^{⟩} again after using it.

14.3 Choosing target byte order

Some types of processors, such as the MIPS, PowerPC, and Hitachi SH, offer the ability to run either big-endian or little-endian byte orders. Usually the executable or symbol will include a bit to designate the endian-ness, and you will not need to worry about which to use. However, you may still find it useful to adjust [No value for “GDBN”]’s idea of processor endian-ness manually.

set endian big

Instruct [No value for “GDBN”] to assume the target is big-endian.

set endian little

Instruct [No value for “GDBN”] to assume the target is little-endian.

set endian auto

Instruct [No value for “GDBN”] to use the byte order associated with the executable.

show endian

Display [No value for “GDBN”]’s current idea of the target byte order.

Note that these commands merely adjust interpretation of symbolic data on the host, and that they have absolutely no effect on the target system.

14.4 Remote debugging

If you are trying to debug a program running on a machine that cannot run [No value for “GDBN”] in the usual way, it is often useful to use remote debugging. For example, you might use remote debugging on an operating system kernel, or on a small system which does not have a general purpose operating system powerful enough to run a full-featured debugger.

Some configurations of [No value for “GDBN”] have special serial or TCP/IP interfaces to make this work with particular debugging targets. In addition, [No value for “GDBN”] comes with a generic serial protocol (specific to [No value for “GDBN”], but not specific to any particular target system) which you can use if you write the remote stubs—the code that runs on the remote system to communicate with [No value for “GDBN”].

Other remote targets may be available in your configuration of [No value for “GDBN”]; use help target to list them.

14.4.1 The [No value for “GDBN”] remote serial protocol

To debug a program running on another machine (the debugging target machine), you must first arrange for all the usual prerequisites for the program to run by itself. For example, for a C program, you need:

1. A startup routine to set up the C runtime environment; these usually have a name like ‘crt0’. The startup routine may be supplied by your hardware supplier, or you may have to write your own.

2. A C subroutine library to support your program’s subroutine calls, notably managing input and output.

3. A way of getting your program to the other machine—for example, a download program. These are often supplied by the hardware manufacturer, but you may have to write your own from hardware documentation.

The next step is to arrange for your program to use a serial port to communicate with the machine where [No value for “GDBN”] is running (the host machine). In general terms, the scheme looks like this:

On the host, [No value for “GDBN”] already understands how to use this protocol; when everything else is set up, you can simply use the ‘target remote’ command (see Chapter 14 [Specifying a Debugging Target], page 129).

On the target, you must link with your program a few special-purpose subroutines that implement the [No value for “GDBN”] remote serial protocol. The file containing these subroutines is called a debugging stub. On certain remote targets, you can use an auxiliary program gdbserver instead of linking a stub into your program. See Section 14.4.1.5 [Using the gdbserver program], page 148, for details.

The debugging stub is specific to the architecture of the remote machine; for example, use ‘sparc-stub.c’ to debug programs on sparc boards.

These working remote stubs are distributed with [No value for “GDBN”]:

i386-stub.c

For Intel 386 and compatible architectures.

m68k-stub.c

For Motorola 680x0 architectures.

sh-stub.c

For Hitachi SH architectures.

sparc-stub.c

For sparc architectures.

sparcl-stub.c

For Fujitsu sparclite architectures.

The ‘README’ file in the [No value for “GDBN”] distribution may list other recently added stubs.

14.4.1.1 What the stub can do for you

The debugging stub for your architecture supplies these three subroutines:

set_debug_traps

This routine arranges for handle_exception to run when your program stops. You must call this subroutine explicitly near the beginning of your program.

handle_exception

This is the central workhorse, but your program never calls it explicitly—the setup code arranges for handle_exception to run when a trap is triggered. handle_exception takes control when your program stops during execution (for example, on a breakpoint), and mediates communications with [No value for “GDBN”] on the host machine. This is where the communications protocol is implemented; handle_exception acts as the [No value for “GDBN”] representative on the target machine. It begins by sending summary information on the state of your program, then continues to execute, retrieving and transmitting any information [No value for “GDBN”] needs, until you execute a [No value for “GDBN”] command that makes your program resume; at that point, handle_exception returns control to your own code on the target machine.

breakpoint

Use this auxiliary subroutine to make your program contain a breakpoint. Depending on the particular situation, this may be the only way for [No value for “GDBN”] to get control. For instance, if your target machine has some sort of interrupt button, you won’t need to call this; pressing the interrupt button transfers control to handle_exception—in effect, to [No value for “GDBN”]. On some machines, simply receiving characters on the serial port may also trigger a trap; again, in that situation, you don’t need to call breakpoint from your own program—simply running ‘target remote’ from the host [No value for “GDBN”] session gets control. Call breakpoint if none of these is true, or if you simply want to make certain your program stops at a predetermined point for the start of your debugging session.

14.4.1.2 What you must do for the stub

The debugging stubs that come with [No value for “GDBN”] are set up for a particular chip architecture, but they have no information about the rest of your debugging target machine.

First of all you need to tell the stub how to communicate with the serial port.

int getDebugChar()

Write this subroutine to read a single character from the serial port. It may be identical to getchar for your target system; a different name is used to allow you to distinguish the two if you wish.

void putDebugChar(int)

Write this subroutine to write a single character to the serial port. It may be identical to putchar for your target system; a different name is used to allow you to distinguish the two if you wish.

If you want [No value for “GDBN”] to be able to stop your program while it is running, you need to use an interrupt-driven serial driver, and arrange for it to stop when it receives a ^C (‘\003’, the control-C character). That is the character which [No value for “GDBN”] uses to tell the remote system to stop.

Getting the debugging target to return the proper status to [No value for “GDBN”] probably requires changes to the standard stub; one quick and dirty way is to just execute a breakpoint instruction (the “dirty” part is that [No value for “GDBN”] reports a SIGTRAP instead of a SIGINT).

Other routines you need to supply are:

void exceptionHandler (int exception number, void *exception address)

Write this function to install exception address in the exception handling tables. You need to do this because the stub does not have any way of knowing what the exception handling tables on your target system are like (for example, the processor’s table might be in rom, containing entries which point to a table in ram). exception number is the exception number which should be changed; its meaning is architecture-dependent (for example, different numbers might represent divide by zero, misaligned access, etc). When this exception occurs, control should be transferred directly to exception address, and the processor state (stack, registers, and so on) should be just as it is when a processor exception occurs. So if you want to use a jump instruction to reach exception address, it should be a simple jump, not a jump to subroutine. For the 386, exception address should be installed as an interrupt gate so that interrupts are masked while the handler runs. The gate should be at privilege level 0 (the most privileged level). The sparc and 68k stubs are able to mask interrupts themselves without help from exceptionHandler.

void flush_i_cache()

On sparc and sparclite only, write this subroutine to flush the instruction cache, if any, on your target machine. If there is no instruction cache, this subroutine may be a no-op.

On target machines that have instruction caches, [No value for “GDBN”] requires this function to make certain that the state of your program is stable.

You must also make sure this library routine is available:

void *memset(void *, int, int)

This is the standard library function memset that sets an area of memory to a known value. If you have one of the free versions of libc.a, memset can be found there; otherwise, you must either obtain it from your hardware manufacturer, or write your own.

If you do not use the GNU C compiler, you may need other standard library subroutines as well; this varies from one stub to another, but in general the stubs are likely to use any of the common library subroutines which [No value for ``GCC''] generates as inline code.

14.4.1.3 Putting it all together

In summary, when your program is ready to debug, you must follow these steps.

1. Make sure you have defined the supporting low-level routines (see Section 14.4.1.2

[What you must do for the stub], page 134):

     getDebugChar, putDebugChar,
     flush_i_cache, memset, exceptionHandler.

2. Insert these lines near the top of your program:

     set_debug_traps();
     breakpoint();

3. For the 680x0 stub only, you need to provide a variable called exceptionHook. Nor-

mally you just use:

     void (*exceptionHook)() = 0;

but if before calling set_debug_traps, you set it to point to a function in your program, that function is called when [No value for ``GDBN''] continues after stopping on a trap (for example, bus error). The function indicated by exceptionHook is called with one parameter: an int which is the exception number.

4. Compile and link together: your program, the [No value for “GDBN”] debugging stub for your target architecture, and the supporting subroutines.

5. Make sure you have a serial connection between your target machine and the [No value for “GDBN”] host, and identify the serial port on the host.

6. Download your program to your target machine (or get it there by whatever means the manufacturer provides), and start it.

7. To start remote debugging, run [No value for “GDBN”] on the host machine, and specify as an executable file the program that is running in the remote machine. This tells [No value for “GDBN”] how to find your program’s symbols and the contents of its pure text.

8. Establish communication using the target remote command. Its argument specifies how to communicate with the target machine—either via a devicename attached to a direct serial line, or a TCP port (usually to a terminal server which in turn has a serial line to the target). For example, to use a serial line connected to the device named

‘/dev/ttyb’:

         target remote /dev/ttyb

To use a TCP connection, use an argument of the form host:port. For example, to connect to port 2828 on a terminal server named manyfarms:

         target remote manyfarms:2828

Now you can use all the usual commands to examine and change data and to step and continue the remote program.

To resume the remote program and stop debugging it, use the detach command.

Whenever [No value for “GDBN”] is waiting for the remote program, if you type the interrupt character (often ^{⟨}\underline{\overline{C-C}}^{⟩}), [No value for “GDBN”] attempts to stop the program. This may or may not succeed, depending in part on the hardware and the serial drivers the remote system uses. If you type the interrupt character once again, [No value for “GDBN”] displays this prompt:

     Interrupted while waiting for the program.
     Give up (and stop debugging it)?  (y or n)

If you type y, [No value for “GDBN”] abandons the remote debugging session. (If you decide you want to try again later, you can use ‘target remote’ again to connect once more.) If you type n, [No value for “GDBN”] goes back to waiting.

14.4.1.4 Communication protocol

The stub files provided with [No value for “GDBN”] implement the target side of the communication protocol, and the [No value for “GDBN”] side is implemented in the [No value for “GDBN”] source file ‘remote.c’. Normally, you can simply allow these subroutines to communicate, and ignore the details. (If you’re implementing your own stub file, you can still ignore the details: start with one of the existing stub files. ‘sparc-stub.c’ is the best organized, and therefore the easiest to read.)

However, there may be occasions when you need to know something about the protocol— for example, if there is only one serial port to your target machine, you might want your program to do something special if it recognizes a packet meant for [No value for “GDBN”].

In the examples below, ‘<-’ and ‘->’ are used to indicate transmitted and received data respectfully.

All [No value for “GDBN”] commands and responses (other than acknowledgments) are sent as a packet. A packet is introduced with the character ‘$’, the actual packet-data, and the terminating character ‘#’ followed by a two-digit checksum:

$packet-data#checksum

The two-digit checksum is computed as the modulo 256 sum of all characters between the leading ‘$’ and the trailing ‘#’ (an eight bit unsigned checksum).

Implementors should note that prior to [No value for “GDBN”] 5.0 the protocol specification also included an optional two-digit sequence-id:

$sequence-id:packet-data#checksum That sequence-id was appended to the acknowledgment. [No value for “GDBN”] has never output sequence-ids. Stubs that handle packets added since [No value for “GDBN”] 5.0 must not accept sequence-id.

When either the host or the target machine receives a packet, the first response expected is an acknowledgment: either ‘+’ (to indicate the package was received correctly) or ‘-’ (to request retransmission):

<- $packet-data#checksum

     -> +

The host ([No value for “GDBN”]) sends commands, and the target (the debugging stub incorporated in your program) sends a response. In the case of step and continue commands, the response is only sent when the operation has completed (the target has again stopped).

packet-data consists of a sequence of characters with the exception of ‘#’ and ‘$’ (see ‘X’ packet for additional exceptions).

Fields within the packet should be separated using ‘,’ ‘;’ or ‘:’. Except where otherwise noted all numbers are represented in HEX with leading zeros suppressed.

Implementors should note that prior to [No value for “GDBN”] 5.0, the character ‘:’ could not appear as the third character in a packet (as it would potentially conflict with the sequence-id).

Response data can be run-length encoded to save space. A ‘*’ means that the next character is an ascii encoding giving a repeat count which stands for that many repetitions of the character preceding the ‘*’. The encoding is n+29, yielding a printable character where n >=3 (which is where rle starts to win). The printable characters ‘$’, ‘#’, ‘+’ and ‘-’ or with a numeric value greater than 126 should not be used.

Some remote systems have used a different run-length encoding mechanism loosely refered to as the cisco encoding. Following the ‘*’ character are two hex digits that indicate the size of the packet.

So:

     "0* "

means the same as "0000".

The error response returned for some packets includes a two character error number. That number is not well defined.

For any command not supported by the stub, an empty response (‘$#00’) should be returned. That way it is possible to extend the protocol. A newer [No value for “GDBN”] can tell if a packet is supported based on that response.

A stub is required to support the ‘g’, ‘G’, ‘m’, ‘M’, ‘c’, and ‘s’ commands. All other commands are optional.

Below is a complete list of all currently defined commands and their corresponding response data:

Packet  Request  Description

extended mode

!

Enable extended mode. In extended mode, the remote server is made persistent. The ‘R’ packet is used to restart the program being debugged.

reply ‘OK’ The remote target both supports and has enabled extended mode.

last signal

?

Indicate the reason the target halted. The reply is the same as for step and continue.

reply  see below
reserved  a  Reserved for future use

set program arguments (reserved)

Aarglen,argnum,arg,...

Initialized ‘argv[]’ array passed into program. arglen specifies the number of bytes in the hex encoded byte stream arg. See ‘gdbserver’ for more details.

reply OK reply ENN

set baud (deprecated) bbaud Change the serial line speed to baud. JTC: When does the transport layer state change? When it’s received, or after the ACK is transmitted. In either case, there are problems if the command or the acknowledgment packet is dropped. Stan: If people really wanted to add something like this, and get it working for the first time, they ought to modify serunix.c to send some kind of out-of-band message to a specially-setup stub and have the switch happen "in between" packets, so that from remote protocol’s point of view, nothing actually happened. set breakpoint (deprecated)

Baddr,mode Set (mode is ‘S’) or clear (mode is ‘C’) a breakpoint at addr. This has been replaced by the ‘Z’ and ‘z’ packets.

continue caddr addr is address to resume. If addr is omitted, resume at current address.

reply  see below

continue with signal Csig;addr Continue with signal sig (hex signal number). If ;addr is omitted, resume at same address.

reply  see below
toggle debug (deprecated) d  toggle debug flag.

detach

D

Detach [No value for “GDBN”] from the remote system. Sent to the remote target before [No value for “GDBN”] disconnects.

reply no response [No value for “GDBN”] does not check for any response after sending this packet.

reserved  e  Reserved for future use
reserved  E  Reserved for future use
reserved  f  Reserved for future use
reserved  F  Reserved for future use
read registers  g  Read general registers.

reply XX...

Each byte of register data is described by two hex digits. The bytes with the register are transmitted in target byte order. The size of each register and their position within the ‘g’ packet are determined by the [No value for “GDBN”] internal macros REG-ISTER RAW SIZE and REGIS-TER NAME macros. The specification of several standard g packets is specified below.

ENN  for an error.

write regs GXX...

See ‘g’ for a description of the XX... data.

reply OK  for success
reply ENN  for an error
reserved  h  Reserved for future use

set thread Hct...

Set thread for subsequent operations (‘m’, ‘M’, ‘g’, ‘G’, et.al.). c = ‘c’ for thread used in step and continue; t... can be -1 for all threads. c = ‘g’ for thread used in other operations. If zero, pick a thread, any thread.

reply OK  for success
reply ENN  for an error

cycle step (draft) iaddr,nnn Step the remote target by a single clock cycle. If ,nnn is present, cycle step nnn cycles. If addr is present, cycle step starting at that address.

signal then cycle step (reserved)

I

See ‘i’ and ‘S’ for likely syntax and semantics.

reserved  j  Reserved for future use
reserved  J  Reserved for future use

kill request

k

FIXME: There is no description of how operate when a specific thread context has been selected (ie. does ’k’ kill only that thread?).

reserved  l  Reserved for future use
reserved  L  Reserved for future use

read memory maddr,length Read length bytes of memory starting at address addr. Neither [No value for “GDBN”] nor the stub assume that sized memory transfers are assumed using word alligned accesses. FIXME: A word aligned memory transfer mechanism is needed.

reply XX...

XX... is mem contents. Can be fewer bytes than requested if able to read only part of the data. Neither [No value for “GDBN”] nor the stub assume that sized memory transfers are assumed using word alligned accesses. FIXME: A word aligned memory transfer mechanism is needed.

reply ENN  NN is errno

write mem Maddr,length:XX...

Write length bytes of memory starting at address addr. XX... is the data.

reply OK  for success

reply ENN for an error (this includes the case where only part of the data was written).

reserved  n  Reserved for future use
reserved  N  Reserved for future use
reserved  o  Reserved for future use
reserved  O  Reserved for future use
read reg (reserved)  pn...  See write register.

return r....

The hex encoded value of the register in target byte order.

write reg Pn...=r...

Write register n... with value r..., which contains two hex digits for each byte in the register (target byte order).

reply OK  for success
reply ENN  for an error

general query qquery Request info about query. In general [No value for “GDBN”] queries have a leading upper case letter. Custom vendor queries should use a company prefix (in lower case) ex: ‘qfsf.var’. query may optionally be followed by a ‘,’ or ‘;’ separated list. Stubs must ensure that they match the full query name.

reply XX...

Hex encoded data from query. The reply can not be empty.

reply ENN  error reply
reply ‘’  Indicating an unrecognized query.

general set Qvar=val Set value of var to val. See ‘q’ for a discussing of naming conventions.

reset (deprecated)  r  Reset the entire system.

remote restart RXX Restart the program being debugged. XX, while needed, is ignored. This packet is only available in extended mode.

no reply  The ‘R’ packet has no reply.

step saddr addr is address to resume. If addr is omitted, resume at same address.

reply  see below
step with signal  Ssig;addr  Like ‘C’ but step not continue.
reply  see below

search taddr:PP,MM Search backwards starting at address addr for a match with pattern PP and mask MM. PP and MM are 4 bytes. addr must be at least 3 digits.

thread alive  TXX  Find out if the thread XX is alive.
reply OK  thread is still alive
reply ENN  thread is dead
reserved  u  Reserved for future use

reserved  U  Reserved for future use
reserved  v  Reserved for future use
reserved  V  Reserved for future use
reserved  w  Reserved for future use
reserved  W  Reserved for future use
reserved  x  Reserved for future use

write mem (binary)

Xaddr,length:XX... addr is address, length is number of bytes, XX... is binary data. The characters $, #, and 0x7d are escaped using 0x7d.

reply OK  for success
reply ENN  for an error
reserved  y  Reserved for future use
reserved  Y  Reserved for future use

remove break or watchpoint (draft) zt,addr,length

See ‘Z’.

insert break or watchpoint (draft)

Zt,addr,length t is type: ‘0’ - software breakpoint, ‘1’ - hardware breakpoint, ‘2’ - write watchpoint, ‘3’ - read watchpoint, ‘4’ - access watchpoint; addr is address; length is in bytes. For a software breakpoint, length specifies the size of the instruction to be patched. For hardware breakpoints and watchpoints length specifies the memory region to be monitored. To avoid potential problems with duplicate packets, the operations should be implemented in an idempotent way.

reply ENN  for an error
reply OK  for success
‘’  If not supported.
reserved  <other>  Reserved for future use

The ‘C’, ‘c’, ‘S’, ‘s’ and ‘?’ packets can receive any of the below as a reply. In the case of the ‘C’, ‘c’, ‘S’ and ‘s’ packets, that reply is only returned when the target halts. In the below the exact meaning of ‘signal number’ is poorly defined. In general one of the UNIX signal numbering conventions is used.

SAA  AA is the signal number

TAAn...:r...;n...:r...;n...:r...; AA = two hex digit signal number; n... = register number (hex), r... = target byte ordered register contents, size defined by REGISTER_RAW_SIZE; n... = ‘thread’, r... = thread process ID, this is a hex integer; n... = other string not starting with valid hex digit. [No value for “GDBN”] should ignore this n..., r... pair and go on to the next. This way we can extend the protocol.

WAA The process exited, and AA is the exit status. This is only applicable for certains sorts of targets.

XAA  The process terminated with signal AA.

NAA;t...;d...;b... (obsolete)

AA = signal number; t... = address of symbol " start"; d... = base of data section; b... = base of bss section. Note: only used by Cisco Systems targets. The difference between this reply and the "qOffsets" query is that the ’N’ packet may arrive spontaneously whereas the ’qOffsets’ is a query initiated by the host debugger.

OXX...

XX... is hex encoding of ascii data. This can happen at any time while the program is running and the debugger should continue to wait for ’W’, ’T’, etc.

The following set and query packets have already been defined.

current thread  qC  Return the current thread id.
reply QCpid  Where pid is a HEX encoded 16 bit process id.
reply *  Any other reply implies the old pid.

all thread ids  qfThreadInfo

qsThreadInfo

Obtain a list of active thread ids from the target (OS). Since there may be too many active threads to fit into one reply packet, this query works iteratively: it may require more than one query/reply sequence to obtain the entire list of threads. The first query of the sequence will be the qfThreadInfo query; subsequent queries in the sequence will be the

qsThreadInfo query.

NOTE: replaces the qL query (see below).

reply m<id>  A single thread id

reply m<id>,<id>...

a comma-separated list of thread ids

reply l  (lower case ’el’) denotes end of list.

In response to each query, the target will reply with a list of one or more thread ids, in big-endian hex, separated by commas. GDB will respond to each reply with a request for more thread ids (using the qs form of the query), until the target responds with l (lower-case el, for 'last').

extra thread info qThreadExtraInfo,id

Where <id> is a thread-id in big-endian hex. Obtain a printable string description of a thread’s attributes from the target OS. This string may contain anything that the target OS thinks is interesting for [No value for “GDBN”] to tell the user about the thread. The string is displayed in [No value for “GDBN”]’s ‘info threads’ display. Some examples of possible thread extra info strings are "Runnable", or "Blocked on Mutex".

reply XX...

Where XX... is a hex encoding of ascii data, comprising the printable string containing the extra information about the thread’s attributes.

query LIST or threadLIST (deprecated) qLstartflagthreadcountnextthread

Obtain thread information from RTOS. Where: startflag (one hex digit) is one to indicate the first query and zero to indicate a subsequent query; threadcount (two hex digits) is the maximum number of threads the response packet can contain; and nextthread (eight hex digits), for subsequent queries (startflag is zero), is returned in the response as argthread. NOTE: this query is replaced by the qfThreadInfo query (see above).

reply qMcountdoneargthreadthread...

Where: count (two hex digits) is the number of threads being returned; done (one hex digit) is zero to indicate more threads and one indicates no further threads; argthreadid (eight hex digits) is nextthread from the request packet; thread... is a sequence of thread IDs from the target. threadid (eight

hex digits). See remote.c:parse_threadlist_
response().

compute CRC of memory block qCRC:addr,length

reply ENN  An error (such as memory fault)

reply CCRC32 A 32 bit cyclic redundancy check of the specified memory region.

query sect offs

qOffsets

Get section offsets that the target used when relocating the downloaded image. Note: while a Bss offset is included in the response, [No value for “GDBN”] ignores this and instead applies the Data offset to the Bss section.

reply

Text=xxx;Data=yyy;Bss=zzz

thread info request qPmodethreadid

Returns information on threadid. Where: mode is a hex encoded 32 bit mode; threadid is a hex encoded 64 bit thread ID.

reply *

See  remote.c:remote_unpack_thread_info_
response().

remote command qRcmd,COMMAND

COMMAND (hex encoded) is passed to the local interpreter for execution. Invalid commands should be reported using the output string. Before the final result packet, the target may also respond with a number of intermediate OOUTPUT console output packets. Implementors should note that providing access to a stubs’s interpreter may have security implications.

reply OK  A command response with no output.

reply OUTPUT A command response with the hex encoded output string OUTPUT.

reply ENN  Indicate a badly formed request.
reply ‘’  When ‘q’‘Rcmd’ is not recognized.

symbol lookup

qSymbol::

Notify the target that [No value for “GDBN”] is prepared to serve symbol lookup requests. Accept requests from the target for the values of symbols. reply OK The target does not need to look up any (more) symbols.

reply

qSymbol:sym name

The target requests the value of symbol sym name (hex encoded). [No value for “GDBN”] may provide the value by using the qSymbol:sym value:sym name message, described below.

symbol value  qSymbol:sym value:sym name Set the value of SYM NAME to SYM VALUE.

sym name (hex encoded) is the name of a symbol whose value the target has previously requested. sym value (hex) is the value for symbol sym name. If [No value for “GDBN”] cannot supply a value for sym name, then this field will be empty.

reply OK The target does not need to look up any (more) symbols.

reply

qSymbol:sym name

The target requests the value of a new symbol sym name (hex encoded). [No value for “GDBN”] will continue to supply the values of symbols (if available), until the target ceases to request them.

The following ‘g’/‘G’ packets have previously been defined. In the below, some thirty-two bit registers are transferred as sixty-four bits. Those registers should be zero/sign extended (which?) to fill the space allocated. Register bytes are transfered in target byte order. The two nibbles within a register byte are transfered most-significant - least-significant.

MIPS32 All registers are transfered as thirty-two bit quantities in the order: 32 general-purpose; sr; lo; hi; bad; cause; pc; 32 floating-point registers; fsr; fir; fp.

MIPS64 All registers are transfered as sixty-four bit quantities (including thirty-two bit registers such as sr). The ordering is the same

        as MIPS32.

Example sequence of a target being re-started. Notice how the restart does not get any direct output:

<- R00
-> +

target restarts

<- ?
-> +
-> T001:1234123412341234
<- +

Example sequence of a target being stepped by a single instruction:

<- G1445...
-> +

<- s
-> +

time passes

-> T001:1234123412341234
<- +
<- g
-> +
-> 1455...
<- +

14.4.1.5 Using the gdbserver program

gdbserver is a control program for Unix-like systems, which allows you to connect your program with a remote [No value for “GDBN”] via target remote—but without linking in the usual debugging stub.

gdbserver is not a complete replacement for the debugging stubs, because it requires essentially the same operating-system facilities that [No value for “GDBN”] itself does. In fact, a system that can run gdbserver to connect to a remote [No value for “GDBN”] could also run [No value for “GDBN”] locally! gdbserver is sometimes useful nevertheless, because it is a much smaller program than [No value for “GDBN”] itself. It is also easier to port than all of [No value for “GDBN”], so you may be able to get started more quickly on a new system by using gdbserver. Finally, if you develop code for real-time systems, you may find that the tradeoffs involved in real-time operation make it more convenient to do as much development work as possible on another system, for example by cross-compiling. You can use gdbserver to make a similar choice for debugging.

[No value for “GDBN”] and gdbserver communicate via either a serial line or a TCP connection, using the standard [No value for “GDBN”] remote serial protocol.

On the target machine, you need to have a copy of the program you want to debug. gdbserver does not need your program’s symbol table, so you can strip the program if necessary to save space. [No value for “GDBN”] on the host system does all the symbol handling. To use the server, you must tell it how to communicate with [No value for “GDBN”]; the name of your program; and the arguments for your program. The syntax is:

     target> gdbserver comm program [ args ... ]

comm is either a device name (to use a serial line) or a TCP hostname and portnumber. For example, to debug Emacs with the argument ‘foo.txt’ and communicate with [No value for “GDBN”] over the serial port ‘/dev/com1’:

     target> gdbserver /dev/com1 emacs foo.txt

gdbserver waits passively for the host [No value for “GDBN”] to communicate with it. To use a TCP connection instead of a serial line:

     target> gdbserver host:2345 emacs foo.txt

The only difference from the previous example is the first argument, specifying that you are communicating with the host [No value for “GDBN”] via TCP. The ‘host:2345’ argument means that gdbserver is to expect a TCP connection from machine ‘host’ to local TCP port 2345. (Currently, the ‘host’ part is ignored.) You can choose any number you want for the port number as long as it does not conflict with any TCP ports already in use on the target system (for example, 23 is reserved for telnet).^{1} You must use the same port number with the host [No value for “GDBN”] target remote command.

On the [No value for “GDBN”] host machine, you need an unstripped copy of your program, since [No value for “GDBN”] needs symbols and debugging information. Start up [No value for “GDBN”] as usual, using the name of the local copy of your program as the first argument. (You may also need the ‘--baud’ option if the serial line is running at anything other than 9600 bps.) After that, use target remote to establish communications with gdbserver. Its argument is either a device name (usually a serial device, like ‘/dev/ttyb’), or a TCP port descriptor in the form host:PORT. For example:

     ([No value for ``GDBP'']) target remote /dev/ttyb

communicates with the server via serial line ‘/dev/ttyb’, and

     ([No value for ``GDBP'']) target remote the-target:2345

communicates via a TCP connection to port 2345 on host ‘the-target’. For TCP connections, you must start up gdbserver prior to using the target remote command. Otherwise you may get an error whose text depends on the host system, but which usually looks something like ‘Connection refused’.

14.4.1.6 Using the gdbserve.nlm program

gdbserve.nlm is a control program for NetWare systems, which allows you to connect your program with a remote [No value for “GDBN”] via target remote.

[No value for “GDBN”] and gdbserve.nlm communicate via a serial line, using the standard [No value for “GDBN”] remote serial protocol.

On the target machine, you need to have a copy of the program you want to debug. gdbserve.nlm does not need your program’s symbol table, so you can strip the program if necessary to save space. [No value for “GDBN”] on the host system does all the symbol handling. To use the server, you must tell it how to communicate with [No value for “GDBN”]; the name of your program; and the arguments for your program. The syntax is:

     load gdbserve [ BOARD=board ] [ PORT=port ]

[ BAUD=baud ] program [ args ... ]

board and port specify the serial line; baud specifies the baud rate used by the connection. port and node default to 0, baud defaults to 9600 bps.

1 If you choose a port number that conflicts with another service, gdbserver prints an error message and

exits.

For example, to debug Emacs with the argument ‘foo.txt’and communicate with [No value for “GDBN”] over serial port number 2 or board 1 using a 19200 bps connection:

     load gdbserve BOARD=1 PORT=2 BAUD=19200 emacs foo.txt

On the [No value for “GDBN”] host machine, you need an unstripped copy of your program, since [No value for “GDBN”] needs symbols and debugging information. Start up [No value for “GDBN”] as usual, using the name of the local copy of your program as the first argument. (You may also need the ‘--baud’ option if the serial line is running at anything other than 9600 bps. After that, use target remote to establish communications with gdbserve.nlm. Its argument is a device name (usually a serial device, like ‘/dev/ttyb’). For example:

     ([No value for ``GDBP'']) target remote /dev/ttyb

communications with the server via serial line ‘/dev/ttyb’.

14.5 Kernel Object Display

Some targets support kernel object display. Using this facility, [No value for “GDBN”] communicates specially with the underlying operating system and can display information about operating system-level objects such as mutexes and other synchronization objects. Exactly which objects can be displayed is determined on a per-OS basis.

Use the set os command to set the operating system. This tells [No value for “GDBN”] which kernel object display module to initialize:

([No value for ``GDBP'']) set os cisco

If set os succeeds, [No value for “GDBN”] will display some information about the operating system, and will create a new info command which can be used to query the target. The info command is named after the operating system:

([No value for ``GDBP'']) info cisco
List of Cisco Kernel Objects
Object     Description
any        Any and all objects

Further subcommands can be used to query about particular objects known by the kernel.

There is currently no way to determine whether a given operating system is supported other than to try it.

15 Configuration-Specific Information

While nearly all [No value for “GDBN”] commands are available for all native and cross versions of the debugger, there are some exceptions. This chapter describes things that are only available in certain configurations.

There are three major categories of configurations: native configurations, where the host and target are the same, embedded operating system configurations, which are usually the same for several different processor architectures, and bare embedded processors, which are quite different from each other.

15.1 Native

This section describes details specific to particular native configurations.

15.1.1 HP-UX

On HP-UX systems, if you refer to a function or variable name that begins with a dollar sign, [No value for “GDBN”] searches for a user or system name first, before it searches for a convenience variable.

15.1.2 SVR4 process information

Many versions of SVR4 provide a facility called ‘/proc’ that can be used to examine the image of a running process using file-system subroutines. If [No value for “GDBN”] is configured for an operating system with this facility, the command info proc is available to report on several kinds of information about the process running your program. info proc works only on SVR4 systems that include the procfs code. This includes OSF/1 (Digital Unix), Solaris, Irix, and Unixware, but not HP-UX or Linux, for example.

info proc Summarize available information about the process.

info proc mappings

Report on the address ranges accessible in the program, with information on whether your program may read, write, or execute each range.

info proc times

Starting time, user CPU time, and system CPU time for your program and its children.

info proc id

Report on the process IDs related to your program: its own process ID, the ID of its parent, the process group ID, and the session ID.

info proc status

General information on the state of the process. If the process is stopped, this report includes the reason for stopping, and any signal received.

info proc all

Show all the above information about the process.

15.1.3 Features for Debugging djgpp Programs

djgpp is the port of gnu development tools to MS-DOS and MS-Windows. djgpp programs are 32-bit protected-mode programs that use the DPMI (DOS Protected-Mode Interface) API to run on top of real-mode DOS systems and their emulations.

[No value for “GDBN”] supports native debugging of djgpp programs, and defines a few commands specific to the djgpp port. This subsection describes those commands.

info dos This is a prefix of djgpp-specific commands which print information about the target system and important OS structures.

info dos sysinfo

This command displays assorted information about the underlying platform: the CPU type and features, the OS version and flavor, the DPMI version, and the available conventional and DPMI memory.

info dos gdt
info dos ldt
info dos idt

These 3 commands display entries from, respectively, Global, Local, and Interrupt Descriptor Tables (GDT, LDT, and IDT). The descriptor tables are data structures which store a descriptor for each segment that is currently in use. The segment’s selector is an index into a descriptor table; the table entry for that index holds the descriptor’s base address and limit, and its attributes and access rights. A typical djgpp program uses 3 segments: a code segment, a data segment (used for both data and the stack), and a DOS segment (which allows access to DOS/BIOS data structures and absolute addresses in conventional memory). However, the DPMI host will usually define additional segments in order to support the DPMI environment. These commands allow to display entries from the descriptor tables. Without an argument, all entries from the specified table are displayed. An argument, which should be an integer expression, means display a single entry whose index is given by the argument. For example, here’s a convenient way to display information about the debugged program’s data segment:

     ([No value for ``GDBP'']) info dos ldt $ds
     0x13f: base=0x11970000 limit=0x0009ffff 32-Bit Data (Read/Write, Exp-up)

This comes in handy when you want to see whether a pointer is outside the data segment’s limit (i.e. garbled).

info dos pde
info dos pte

These two commands display entries from, respectively, the Page Directory and the Page Tables. Page Directories and Page Tables are data structures which control how virtual memory addresses are mapped into physical addresses. A Page Table includes an entry for every page of memory that is mapped into the program’s address space; there may be several Page Tables, each one holding up to 4096 entries. A Page Directory has up to 4096 entries, one each for every Page Table that is currently in use. Without an argument, info dos pde displays the entire Page Directory, and info dos pte displays all the entries in all of the Page Tables. An argument, an integer expression, given to the info dos pde command means display only that entry from the Page Directory table. An argument given to the info dos pte command means display entries from a single Page Table, the one pointed to by the specified entry in the Page Directory. These commands are useful when your program uses DMA (Direct Memory Access), which needs physical addresses to program the DMA controller. These commands are supported only with some DPMI servers.

info dos address-pte

This command displays the Page Table entry for a specified linear address. The argument linear address should already have the appropriate segment’s base address added to it, because this command accepts addresses which may belong to any segment. For example, here’s how to display the Page Table entry for the page where the variable i is stored:

     ([No value for ``GDBP'']) info dos address-pte __djgpp_base_address + (char *)&i
     Page Table entry for address 0x11a00d30:
     Base=0x02698000 Dirty Acc. Not-Cached Write-Back Usr Read-Write +0xd30

This says that i is stored at offset 0xd30 from the page whose physical base address is 0x02698000, and prints all the attributes of that page. Note that you must cast the addresses of variables to a char *, since otherwise the value of __djgpp_base_address, the base address of all variables and functions in a djgpp program, will be added using the rules of C pointer arithmetics: if i is declared an int, [No value for “GDBN”] will add 4 times the value of __djgpp_base_address to the address of i. Here’s another example, it displays the Page Table entry for the transfer buffer:

     ([No value for ``GDBP'']) info dos address-pte *((unsigned *)&_go32_info_block + 3)
     Page Table entry for address 0x29110:
     Base=0x00029000 Dirty Acc. Not-Cached Write-Back Usr Read-Write +0x110

(The + 3 offset is because the transfer buffer’s address is the 3rd member of the _go32_info_block structure.) The output of this command clearly shows that addresses in conventional memory are mapped 1:1, i.e. the physical and linear addresses are identical. This command is supported only with some DPMI servers.

15.2 Embedded Operating Systems

This section describes configurations involving the debugging of embedded operating systems that are available for several different architectures.

[No value for “GDBN”] includes the ability to debug programs running on various realtime operating systems.

15.2.1 Using [No value for “GDBN”] with VxWorks

target vxworks machinename

A VxWorks system, attached via TCP/IP. The argument machinename is the target system’s machine name or IP address.

On VxWorks, load links filename dynamically on the current target system as well as adding its symbols in [No value for “GDBN”].

[No value for “GDBN”] enables developers to spawn and debug tasks running on networked VxWorks targets from a Unix host. Already-running tasks spawned from the Vx-Works shell can also be debugged. [No value for “GDBN”] uses code that runs on both the Unix host and on the VxWorks target. The program [No value for ``GDBP''] is installed and executed on the Unix host. (It may be installed with the name vxgdb, to distinguish it from a [No value for “GDBN”] for debugging programs on the host itself.)

VxWorks-timeout args

All VxWorks-based targets now support the option vxworks-timeout. This option is set by the user, and args represents the number of seconds [No value for “GDBN”] waits for responses to rpc’s. You might use this if your VxWorks target is a slow software simulator or is on the far side of a thin network line.

The following information on connecting to VxWorks was current when this manual was produced; newer releases of VxWorks may use revised procedures.

To use [No value for “GDBN”] with VxWorks, you must rebuild your VxWorks kernel to include the remote debugging interface routines in the VxWorks library ‘rdb.a’. To do this, define INCLUDE_RDB in the VxWorks configuration file ‘configAll.h’ and rebuild your VxWorks kernel. The resulting kernel contains ‘rdb.a’, and spawns the source debugging task tRdbTask when VxWorks is booted. For more information on configuring and remaking VxWorks, see the manufacturer’s manual.

Once you have included ‘rdb.a’ in your VxWorks system image and set your Unix execution search path to find [No value for “GDBN”], you are ready to run [No value for “GDBN”]. From your Unix host, run [No value for ``GDBP''] (or vxgdb, depending on your installation).

[No value for “GDBN”] comes up showing the prompt:

     (vxgdb)

15.2.1.1 Connecting to VxWorks

The [No value for “GDBN”] command target lets you connect to a VxWorks target on the network. To connect to a target whose host name is “tt”, type:

     (vxgdb) target vxworks tt

[No value for “GDBN”] displays messages like these:

     Attaching remote machine across net...
     Connected to tt.

[No value for “GDBN”] then attempts to read the symbol tables of any object modules loaded into the VxWorks target since it was last booted. [No value for “GDBN”] locates these files by searching the directories listed in the command search path (see Section 4.4 [Your program’s environment], page 25); if it fails to find an object file, it displays a message such as:

     prog.o: No such file or directory.

When this happens, add the appropriate directory to the search path with the [No value for “GDBN”] command path, and execute the target command again.

15.2.1.2 VxWorks download

If you have connected to the VxWorks target and you want to debug an object that has not yet been loaded, you can use the [No value for “GDBN”] load command to download a file from Unix to VxWorks incrementally. The object file given as an argument to the load command is actually opened twice: first by the VxWorks target in order to download the code, then by [No value for “GDBN”] in order to read the symbol table. This can lead to problems if the current working directories on the two systems differ. If both systems have NFS mounted the same filesystems, you can avoid these problems by using absolute paths. Otherwise, it is simplest to set the working directory on both systems to the directory in which the object file resides, and then to reference the file by its name, without any path. For instance, a program ‘prog.o’ may reside in ‘vxpath/vw/demo/rdb’ in VxWorks and in ‘hostpath/vw/demo/rdb’ on the host. To load this program, type this on VxWorks:

     -> cd "vxpath/vw/demo/rdb"

Then, in [No value for “GDBN”], type:

     (vxgdb) cd hostpath/vw/demo/rdb
     (vxgdb) load prog.o

[No value for “GDBN”] displays a response similar to this:

     Reading symbol data from wherever/vw/demo/rdb/prog.o... done.

You can also use the load command to reload an object module after editing and recompiling the corresponding source file. Note that this makes [No value for “GDBN”] delete all currently-defined breakpoints, auto-displays, and convenience variables, and to clear the value history. (This is necessary in order to preserve the integrity of debugger’s data structures that reference the target system’s symbol table.)

15.2.1.3 Running tasks

You can also attach to an existing task using the attach command as follows:

     (vxgdb) attach task

where task is the VxWorks hexadecimal task ID. The task can be running or suspended when you attach to it. Running tasks are suspended at the time of attachment.

15.3 Embedded Processors

This section goes into details specific to particular embedded configurations.

15.3.1 AMD A29K Embedded

target adapt dev

Adapt monitor for A29K.

target amd-eb dev speed PROG

Remote PC-resident AMD EB29K board, attached over serial lines. dev is the serial device, as for target remote; speed allows you to specify the linespeed; and PROG is the name of the program to be debugged, as it appears to DOS on the PC. See Section 15.3.1.2 [EBMON protocol for AMD29K], page 156.

15.3.1.1 A29K UDI

[No value for “GDBN”] supports AMD’s UDI (“Universal Debugger Interface”) protocol for debugging the a29k processor family. To use this configuration with AMD targets running the MiniMON monitor, you need the program MONTIP, available from AMD at no charge. You can also use [No value for “GDBN”] with the UDI-conformant a29k simulator program ISSTIP, also available from AMD.

target udi keyword

Select the UDI interface to a remote a29k board or simulator, where keyword is an entry in the AMD configuration file ‘udi_soc’. This file contains keyword entries which specify parameters used to connect to a29k targets. If the ‘udi_soc’ file is not in your working directory, you must set the environment variable ‘UDICONF’ to its pathname.

15.3.1.2 EBMON protocol for AMD29K

AMD distributes a 29K development board meant to fit in a PC, together with a DOShosted monitor program called EBMON. As a shorthand term, this development system is called the “EB29K”. To use [No value for “GDBN”] from a Unix system to run programs on the EB29K board, you must first connect a serial cable between the PC (which hosts the EB29K board) and a serial port on the Unix system. In the following, we assume you’ve hooked the cable between the PC’s ‘COM1’ port and ‘/dev/ttya’ on the Unix system.

15.3.1.3 Communications setup

The next step is to set up the PC’s port, by doing something like this in DOS on the PC:

     C:\> MODE com1:9600,n,8,1,none

This example—run on an MS DOS 4.0 system—sets the PC port to 9600 bps, no parity, eight data bits, one stop bit, and no “retry” action; you must match the communications parameters when establishing the Unix end of the connection as well.

To give control of the PC to the Unix side of the serial line, type the following at the DOS console:

     C:\> CTTY com1

(Later, if you wish to return control to the DOS console, you can use the command CTTY con—but you must send it over the device that had control, in our example over the ‘COM1’ serial line.)

From the Unix host, use a communications program such as tip or cu to communicate with the PC; for example,

cu -s 9600 -l /dev/ttya

The cu options shown specify, respectively, the linespeed and the serial port to use. If you use tip instead, your command line may look something like the following:

tip -9600 /dev/ttya

Your system may require a different name where we show ‘/dev/ttya’ as the argument to tip. The communications parameters, including which port to use, are associated with the tip argument in the “remote” descriptions file—normally the system table ‘/etc/remote’.

Using the tip or cu connection, change the DOS working directory to the directory containing a copy of your 29K program, then start the PC program EBMON (an EB29K control program supplied with your board by AMD). You should see an initial display from EBMON similar to the one that follows, ending with the EBMON prompt ‘#’—

C:\> G:
G:\> CD \usr\joe\work29k
G:\USR\JOE\WORK29K> EBMON
Am29000 PC Coprocessor Board Monitor, version 3.0-18
Copyright 1990 Advanced Micro Devices, Inc.
Written by Gibbons and Associates, Inc.
Enter '?' or 'H' for help
PC Coprocessor Type   = EB29K
I/O Base              = 0x208
Memory Base           = 0xd0000
Data Memory Size      = 2048KB
Available I-RAM Range = 0x8000 to 0x1fffff
Available D-RAM Range = 0x80002000 to 0x801fffff
PageSize              = 0x400
Register Stack Size   = 0x800
Memory Stack Size     = 0x1800
CPU PRL               = 0x3
Am29027 Available     = No
Byte Write Available  = Yes
# ~.

Then exit the cu or tip program (done in the example by typing ~. at the EBMON prompt). EBMON keeps running, ready for [No value for “GDBN”] to take over.

For this example, we’ve assumed what is probably the most convenient way to make sure the same 29K program is on both the PC and the Unix system: a PC/NFS connection that establishes “drive ‘G:’” on the PC as a file system on the Unix host. If you do not have PC/NFS or something similar connecting the two systems, you must arrange some other way—perhaps floppy-disk transfer—of getting the 29K program from the Unix system to the PC; [No value for “GDBN”] does not download it over the serial line.

15.3.1.4 EB29K cross-debugging

Finally, cd to the directory containing an image of your 29K program on the Unix system, and start [No value for “GDBN”]—specifying as argument the name of your 29K program:

     cd /usr/joe/work29k
     [No value for ``GDBP''] myfoo

Now you can use the target command:

     target amd-eb /dev/ttya 9600 MYFOO

In this example, we’ve assumed your program is in a file called ‘myfoo’. Note that the filename given as the last argument to target amd-eb should be the name of the program as it appears to DOS. In our example this is simply MYFOO, but in general it can include a DOS path, and depending on your transfer mechanism may not resemble the name on the Unix side.

At this point, you can set any breakpoints you wish; when you are ready to see your program run on the 29K board, use the [No value for “GDBN”] command run.

To stop debugging the remote program, use the [No value for “GDBN”] detach command.

To return control of the PC to its console, use tip or cu once again, after your [No value for “GDBN”] session has concluded, to attach to EBMON. You can then type the command q to shut down EBMON, returning control to the DOS command-line interpreter. Type CTTY con to return command input to the main DOS console, and type ~. to leave tip or cu.

15.3.1.5 Remote log

The target amd-eb command creates a file ‘eb.log’ in the current working directory, to help debug problems with the connection. ‘eb.log’ records all the output from EBMON, including echoes of the commands sent to it. Running ‘tail -f’ on this file in another window often helps to understand trouble with EBMON, or unexpected events on the PC side of the connection.

15.3.2 ARM

target rdi dev

ARM Angel monitor, via RDI library interface to ADP protocol. You may use this target to communicate with both boards running the Angel monitor, or with the EmbeddedICE JTAG debug device.

target rdp dev

ARM Demon monitor.

15.3.3 Hitachi H8/300

target hms dev

A Hitachi SH, H8/300, or H8/500 board, attached via serial line to your host. Use special commands device and speed to control the serial line and the communications speed used.

target e7000 dev

E7000 emulator for Hitachi H8 and SH.

target sh3 dev
target sh3e dev

Hitachi SH-3 and SH-3E target systems.

When you select remote debugging to a Hitachi SH, H8/300, or H8/500 board, the load command downloads your program to the Hitachi board and also opens it as the current executable target for [No value for “GDBN”] on your host (like the file command).

[No value for “GDBN”] needs to know these things to talk to your Hitachi SH, H8/300, or H8/500:

1. that you want to use ‘target hms’, the remote debugging interface for Hitachi microprocessors, or ‘target e7000’, the in-circuit emulator for the Hitachi SH and the Hitachi 300H. (‘target hms’ is the default when [No value for “GDBN”] is configured specifically for the Hitachi SH, H8/300, or H8/500.)

2. what serial device connects your host to your Hitachi board (the first serial device available on your host is the default).

3. what speed to use over the serial device.

15.3.3.1 Connecting to Hitachi boards

Use the special [No value for ``GDBN''] command ‘device port’ if you need to explicitly set the serial device. The default port is the first available port on your host. This is only necessary on Unix hosts, where it is typically something like ‘/dev/ttya’.

[No value for ``GDBN''] has another special command to set the communications speed: ‘speed bps’. This command also is only used from Unix hosts; on DOS hosts, set the line speed as usual from outside [No value for “GDBN”] with the DOS mode command (for instance, mode com2:9600,n,8,1,p for a 9600 bps connection).

The ‘device’ and ‘speed’ commands are available only when you use a Unix host to debug your Hitachi microprocessor programs. If you use a DOS host, [No value for “GDBN”] depends on an auxiliary terminate-and-stay-resident program called asynctsr to communicate with the development board through a PC serial port. You must also use the DOS mode command to set up the serial port on the DOS side.

The following sample session illustrates the steps needed to start a program under [No value for “GDBN”] control on an H8/300. The example uses a sample H8/300 program called ‘t.x’. The procedure is the same for the Hitachi SH and the H8/500.

First hook up your development board. In this example, we use a board attached to serial port COM2; if you use a different serial port, substitute its name in the argument of the mode command. When you call asynctsr, the auxiliary comms program used by the debugger, you give it just the numeric part of the serial port’s name; for example, ‘asyncstr

2’ below runs asyncstr on COM2.
     C:\H8300\TEST> asynctsr 2
     C:\H8300\TEST> mode com2:9600,n,8,1,p
     Resident portion of MODE loaded
     COM2: 9600, n, 8, 1, p

Warning: We have noticed a bug in PC-NFS that conflicts with asynctsr. If you also run PC-NFS on your DOS host, you may need to disable it, or even boot without it, to use asynctsr to control your development board.

Now that serial communications are set up, and the development board is connected, you can start up [No value for “GDBN”]. Call [No value for ``GDBP''] with the name of your program as the argument. [No value for ``GDBN''] prompts you, as usual, with the prompt ‘([No value for ``GDBP''])’. Use two special commands to begin your debugging session: ‘target hms’ to specify cross-debugging to the Hitachi board, and the load command to download your program to the board. load displays the names of the program’s sections, and a ‘*’ for each 2K of data downloaded. (If you want to refresh [No value for “GDBN”] data on symbols or on the executable file without downloading, use the [No value for “GDBN”] commands file or symbol-file. These commands, and load itself, are described in Section 13.1 [Commands to specify files], page 123.)

     (eg-C:\H8300\TEST) [No value for ``GDBP''] t.x
     [No value for ``GDBN''] is free software and you are welcome to distribute copies
      of it under certain conditions; type "show copying" to see
      the conditions.
     There is absolutely no warranty for [No value for ``GDBN'']; type "show warranty"
     for details.
     [No value for ``GDBN''] [No value for ``GDBVN''], Copyright 1992 Free Software Foundation, Inc...
     ([No value for ``GDBP'']) target hms
     Connected to remote H8/300 HMS system.
     ([No value for ``GDBP'']) load t.x
     .text   : 0x8000 .. 0xabde ***********
     .data   : 0xabde .. 0xad30 *
     .stack  : 0xf000 .. 0xf014 *

At this point, you’re ready to run or debug your program. From here on, you can use all the usual [No value for “GDBN”] commands. The break command sets breakpoints; the run command starts your program; print or x display data; the continue command resumes execution after stopping at a breakpoint. You can use the help command at any time to find out more about [No value for “GDBN”] commands.

Remember, however, that operating system facilities aren’t available on your development board; for example, if your program hangs, you can’t send an interrupt—but you can press the reset switch!

Use the reset button on the development board

• to interrupt your program (don’t use ctl-C on the DOS host—it has no way to pass an interrupt signal to the development board); and • to return to the [No value for “GDBN”] command prompt after your program fin-

ishes normally. The communications protocol provides no other way for [No value for “GDBN”] to detect program completion.

In either case, [No value for “GDBN”] sees the effect of a reset on the development board as a “normal exit” of your program.

15.3.3.2 Using the E7000 in-circuit emulator

You can use the E7000 in-circuit emulator to develop code for either the Hitachi SH or the H8/300H. Use one of these forms of the ‘target e7000’ command to connect [No value for “GDBN”] to your E7000:

target e7000 port speed

Use this form if your E7000 is connected to a serial port. The port argument identifies what serial port to use (for example, ‘com2’). The third argument is the line speed in bits per second (for example, ‘9600’).

target e7000 hostname

If your E7000 is installed as a host on a TCP/IP network, you can just specify its hostname; [No value for “GDBN”] uses telnet to connect.

15.3.3.3 Special [No value for “GDBN”] commands for Hitachi micros

Some [No value for “GDBN”] commands are available only for the H8/300:

set machine h8300
set machine h8300h

Condition [No value for “GDBN”] for one of the two variants of the H8/300 architecture with ‘set machine’. You can use ‘show machine’ to check which variant is currently in effect.

15.3.4 H8/500

set memory mod
show memory

Specify which H8/500 memory model (mod) you are using with ‘set memory’; check which memory model is in effect with ‘show memory’. The accepted values

for mod are small, big, medium, and compact.

15.3.5 Intel i960

target mon960 dev

MON960 monitor for Intel i960.

target nindy devicename

An Intel 960 board controlled by a Nindy Monitor. devicename is the name of the serial device to use for the connection, e.g. ‘/dev/ttya’.

Nindy is a ROM Monitor program for Intel 960 target systems. When [No value for “GDBN”] is configured to control a remote Intel 960 using Nindy, you can tell [No value for “GDBN”] how to connect to the 960 in several ways:

• Through command line options specifying serial port, version of the Nindy protocol, and communications speed;

• By responding to a prompt on startup;

• By using the target command at any point during your [No value for “GDBN”] session. See Section 14.2 [Commands for managing targets], page 129.

With the Nindy interface to an Intel 960 board, load downloads filename to the 960 as well as adding its symbols in [No value for “GDBN”].

15.3.5.1 Startup with Nindy

If you simply start [No value for ``GDBP''] without using any command-line options, you are prompted for what serial port to use, before you reach the ordinary [No value for “GDBN”] prompt:

     Attach /dev/ttyNN -- specify NN, or "quit" to quit:

Respond to the prompt with whatever suffix (after ‘/dev/tty’) identifies the serial port you want to use. You can, if you choose, simply start up with no Nindy connection by responding to the prompt with an empty line. If you do this and later wish to attach to Nindy, use target (see Section 14.2 [Commands for managing targets], page 129).

15.3.5.2 Options for Nindy

These are the startup options for beginning your [No value for “GDBN”] session with a Nindy-960 board attached:

-r port Specify the serial port name of a serial interface to be used to connect to the target system. This option is only available when [No value for “GDBN”] is configured for the Intel 960 target architecture. You may specify port as any of: a full pathname (e.g. ‘-r /dev/ttya’), a device name in ‘/dev’ (e.g. ‘-r ttya’), or simply the unique suffix for a specific tty (e.g. ‘-r a’).

-O (An uppercase letter “O”, not a zero.) Specify that [No value for “GDBN”] should use the “old” Nindy monitor protocol to connect to the target system. This option is only available when [No value for “GDBN”] is configured for the Intel 960 target architecture.

Warning: if you specify ‘-O’, but are actually trying to connect to a target system that expects the newer protocol, the connection fails, appearing to be a speed mismatch. [No value for “GDBN”] repeatedly attempts to reconnect at several different line speeds. You can abort this process with an interrupt.

-brk Specify that [No value for “GDBN”] should first send a BREAK signal to the

target system, in an attempt to reset it, before connecting to a Nindy target. Warning: Many target systems do not have the hardware that this requires; it only works with a few boards.

The standard ‘-b’ option controls the line speed used on the serial port.

15.3.5.3 Nindy reset command

reset For a Nindy target, this command sends a “break” to the remote target system; this is only useful if the target has been equipped with a circuit to perform a hard reset (or some other interesting action) when a break is detected.

15.3.6 Mitsubishi M32R/D

target m32r dev

Mitsubishi M32R/D ROM monitor.

15.3.7 M68k

The Motorola m68k configuration includes ColdFire support, and target command for the following ROM monitors.

target abug dev

ABug ROM monitor for M68K.

target cpu32bug dev

CPU32BUG monitor, running on a CPU32 (M68K) board.

target dbug dev

dBUG ROM monitor for Motorola ColdFire.

target est dev

EST-300 ICE monitor, running on a CPU32 (M68K) board.

target rom68k dev

ROM 68K monitor, running on an M68K IDP board.

If [No value for “GDBN”] is configured with m68*-ericsson-*, it will instead have only a single special target command:

target es1800 dev

ES-1800 emulator for M68K.

[context?]

target rombug dev

ROMBUG ROM monitor for OS/9000.

15.3.8 M88K

target bug dev

BUG monitor, running on a MVME187 (m88k) board.

15.3.9 MIPS Embedded

[No value for “GDBN”] can use the MIPS remote debugging protocol to talk to a MIPS board attached to a serial line. This is available when you configure [No value for “GDBN”]

with ‘--target=mips-idt-ecoff’.

Use these [No value for “GDBN”] commands to specify the connection to your target board:

target mips port

To run a program on the board, start up [No value for ``GDBP''] with the name of your program as the argument. To connect to the board, use the command ‘target mips port’, where port is the name of the serial port connected to the board. If the program has not already been downloaded to the board, you may use the load command to download it. You can then use all the usual [No value for “GDBN”] commands. For example, this sequence connects to the target board through a serial port, and loads and runs a program called prog through the debugger:

     host$ [No value for ``GDBP''] prog
     [No value for ``GDBN''] is free software and ...
     ([No value for ``GDBP'']) target mips /dev/ttyb
     ([No value for ``GDBP'']) load prog
     ([No value for ``GDBP'']) run

target mips hostname:portnumber On some [No value for “GDBN”] host configurations, you can specify a TCP connection (for instance, to a serial line managed by a terminal concentrator) instead of a serial port, using the syntax ‘hostname:portnumber’.

target pmon port

PMON ROM monitor.

target ddb port

NEC’s DDB variant of PMON for Vr4300.

target lsi port

LSI variant of PMON.

target r3900 dev

Densan DVE-R3900 ROM monitor for Toshiba R3900 Mips.

target array dev

Array Tech LSI33K RAID controller board.

[No value for “GDBN”] also supports these special commands for MIPS targets:

set processor args
show processor

Use the set processor command to set the type of MIPS processor when you want to access processor-type-specific registers. For example, set processor r3041 tells [No value for “GDBN”] to use the CPU registers appropriate for the 3041 chip. Use the show processor command to see what MIPS processor [No value for “GDBN”] is using. Use the info reg command to see what registers [No value for “GDBN”] is using.

set mipsfpu double
set mipsfpu single
set mipsfpu none
show mipsfpu

If your target board does not support the MIPS floating point coprocessor, you should use the command ‘set mipsfpu none’ (if you need this, you may wish to put the command in your [No value for “GDBN”] init file). This tells [No value for “GDBN”] how to find the return value of functions which return floating point values. It also allows [No value for “GDBN”] to avoid saving the floating point registers when calling functions on the board. If you are using a floating point coprocessor with only single precision floating point support, as on the r4650 processor, use the command ‘set mipsfpu single’. The default double precision floating point coprocessor may be selected using ‘set mipsfpu

double’.

In previous versions the only choices were double precision or no floating point, so ‘set mipsfpu on’ will select double precision and ‘set mipsfpu off’ will select no floating point. As usual, you can inquire about the mipsfpu variable with ‘show mipsfpu’.

set remotedebug n
show remotedebug

You can see some debugging information about communications with the board by setting the remotedebug variable. If you set it to 1 using ‘set remotedebug 1’, every packet is displayed. If you set it to 2, every character is displayed. You can check the current value at any time with the command ‘show remotedebug’.

set timeout seconds
set retransmit-timeout seconds
show timeout
show retransmit-timeout

You can control the timeout used while waiting for a packet, in the MIPS remote protocol, with the set timeout seconds command. The default is 5 seconds. Similarly, you can control the timeout used while waiting for an acknowledgement of a packet with the set retransmit-timeout seconds command. The default is 3 seconds. You can inspect both values with show timeout and show retransmit-timeout. (These commands are only available when [No value for “GDBN”] is configured for ‘--target=mips-idt-ecoff’.) The timeout set by set timeout does not apply when [No value for “GDBN”] is waiting for your program to stop. In that case, [No value for “GDBN”] waits forever because it has no way of knowing how long the program is going to run before stopping.

15.3.10 PowerPC

target dink32 dev

DINK32 ROM monitor.

target ppcbug dev

target ppcbug1 dev

PPCBUG ROM monitor for PowerPC.

target sds dev

SDS monitor, running on a PowerPC board (such as Motorola’s ADS).

15.3.11 HP PA Embedded

target op50n dev

OP50N monitor, running on an OKI HPPA board.

target w89k dev

W89K monitor, running on a Winbond HPPA board.

15.3.12 Hitachi SH

target hms dev

A Hitachi SH board attached via serial line to your host. Use special commands device and speed to control the serial line and the communications speed used.

target e7000 dev

E7000 emulator for Hitachi SH.

target sh3 dev
target sh3e dev

Hitachi SH-3 and SH-3E target systems.

15.3.13 Tsqware Sparclet

[No value for “GDBN”] enables developers to debug tasks running on Sparclet targets from a Unix host. [No value for “GDBN”] uses code that runs on both the Unix host and on the Sparclet target. The program [No value for ``GDBP''] is installed and executed on the Unix host.

remotetimeout args

[No value for “GDBN”] supports the option remotetimeout. This option is set by the user, and args represents the number of seconds [No value for “GDBN”] waits for responses.

When compiling for debugging, include the options ‘-g’ to get debug information and ‘-Ttext’ to relocate the program to where you wish to load it on the target. You may also want to add the options ‘-n’ or ‘-N’ in order to reduce the size of the sections. Example:

     sparclet-aout-gcc prog.c -Ttext 0x12010000 -g -o prog -N

You can use objdump to verify that the addresses are what you intended:

     sparclet-aout-objdump --headers --syms prog

Once you have set your Unix execution search path to find [No value for “GDBN”], you are ready to run [No value for “GDBN”]. From your Unix host, run [No value for ``GDBP''] (or sparclet-aout-gdb, depending on your installation).

[No value for “GDBN”] comes up showing the prompt:

     (gdbslet)

15.3.13.1 Setting file to debug

The [No value for “GDBN”] command file lets you choose with program to debug.

     (gdbslet) file prog

[No value for “GDBN”] then attempts to read the symbol table of ‘prog’. [No value for “GDBN”] locates the file by searching the directories listed in the command search path. If the file was compiled with debug information (option "-g"), source files will be searched as well. [No value for “GDBN”] locates the source files by searching the directories listed in the directory search path (see Section 4.4 [Your program’s environment], page 25). If it fails to find a file, it displays a message such as:

     prog: No such file or directory.

When this happens, add the appropriate directories to the search paths with the [No value for “GDBN”] commands path and dir, and execute the target command again.

15.3.13.2 Connecting to Sparclet

The [No value for “GDBN”] command target lets you connect to a Sparclet target. To connect to a target on serial port “ttya”, type:

     (gdbslet) target sparclet /dev/ttya
     Remote target sparclet connected to /dev/ttya
     main () at ../prog.c:3

[No value for “GDBN”] displays messages like these:

     Connected to ttya.

15.3.13.3 Sparclet download

Once connected to the Sparclet target, you can use the [No value for “GDBN”] load command to download the file from the host to the target. The file name and load offset should be given as arguments to the load command. Since the file format is aout, the program must be loaded to the starting address. You can use objdump to find out what this value is. The load offset is an offset which is added to the VMA (virtual memory address) of each of the file’s sections. For instance, if the program ‘prog’ was linked to text address 0x1201000, with data at 0x12010160 and bss at 0x12010170, in [No value for “GDBN”], type:

     (gdbslet) load prog 0x12010000
     Loading section .text, size 0xdb0 vma 0x12010000

If the code is loaded at a different address then what the program was linked to, you may need to use the section and add-symbol-file commands to tell [No value for “GDBN”] where to map the symbol table.

15.3.13.4 Running and debugging

You can now begin debugging the task using [No value for “GDBN”]’s execution control commands, b, step, run, etc. See the [No value for “GDBN”] manual for the list of commands.

     (gdbslet) b main
     Breakpoint 1 at 0x12010000: file prog.c, line 3.
     (gdbslet) run
     Starting program: prog
     Breakpoint 1, main (argc=1, argv=0xeffff21c) at prog.c:3
     3        char *symarg = 0;
     (gdbslet) step
     4        char *execarg = "hello!";
     (gdbslet)

15.3.14 Fujitsu Sparclite

target sparclite dev

Fujitsu sparclite boards, used only for the purpose of loading. You must use an additional command to debug the program. For example: target remote dev using [No value for “GDBN”] standard remote protocol.

15.3.15 Tandem ST2000

[No value for “GDBN”] may be used with a Tandem ST2000 phone switch, running Tandem’s STDBUG protocol.

To connect your ST2000 to the host system, see the manufacturer’s manual. Once the ST2000 is physically attached, you can run:

     target st2000 dev speed

to establish it as your debugging environment. dev is normally the name of a serial device, such as ‘/dev/ttya’, connected to the ST2000 via a serial line. You can instead specify dev as a TCP connection (for example, to a serial line attached via a terminal concentrator) using the syntax hostname:portnumber.

The load and attach commands are not defined for this target; you must load your program into the ST2000 as you normally would for standalone operation. [No value for “GDBN”] reads debugging information (such as symbols) from a separate, debugging version of the program available on your host computer.

These auxiliary [No value for “GDBN”] commands are available to help you with the ST2000 environment:

st2000 command Send a command to the STDBUG monitor. See the manufacturer’s manual for available commands.

connect Connect the controlling terminal to the STDBUG command monitor. When you are done interacting with STDBUG, typing either of two character sequences gets you back to the [No value for “GDBN”] command prompt: ^{⟨}\underline{\overline{RET}}^{⟩}~. (Return, followed by tilde and period) or ^{⟨}\underline{\overline{RET}}^{⟩}~^{⟨}\underline{\overline{C-d}}^{⟩} (Return, followed by tilde and control-D).

15.3.16 Zilog Z8000

When configured for debugging Zilog Z8000 targets, [No value for “GDBN”] includes a Z8000 simulator.

For the Z8000 family, ‘target sim’ simulates either the Z8002 (the unsegmented variant of the Z8000 architecture) or the Z8001 (the segmented variant). The simulator recognizes which architecture is appropriate by inspecting the object code.

target sim args

Debug programs on a simulated CPU. If the simulator supports setup options, specify them via args.

After specifying this target, you can debug programs for the simulated CPU in the same style as programs for your host computer; use the file command to load a new program image, the run command to run your program, and so on.

As well as making available all the usual machine registers (see Section 8.10 [Registers], page 79), the Z8000 simulator provides three additional items of information as specially named registers:

cycles Counts clock-ticks in the simulator.

insts Counts instructions run in the simulator.

time Execution time in 60ths of a second.

You can refer to these values in [No value for “GDBN”] expressions with the usual conventions; for example, ‘b fputc if $cycles>5000’ sets a conditional breakpoint that suspends only after at least 5000 simulated clock ticks.

15.4 Architectures

This section describes characteristics of architectures that affect all uses of [No value for “GDBN”] with the architecture, both native and cross.

15.4.1 A29K

set rstack_high_address address

On AMD 29000 family processors, registers are saved in a separate register stack. There is no way for [No value for “GDBN”] to determine the extent of this stack. Normally, [No value for “GDBN”] just assumes that the stack is “large enough”. This may result in [No value for “GDBN”] referencing memory locations that do not exist. If necessary, you can get around this problem by specifying the ending address of the register stack with the set rstack_high_ address command. The argument should be an address, which you probably want to precede with ‘0x’ to specify in hexadecimal.

show rstack_high_address

Display the current limit of the register stack, on AMD 29000 family processors.

15.4.2 Alpha

See the following section.

15.4.3 MIPS

Alpha- and MIPS-based computers use an unusual stack frame, which sometimes requires [No value for “GDBN”] to search backward in the object code to find the beginning of a function.

To improve response time (especially for embedded applications, where [No value for “GDBN”] may be restricted to a slow serial line for this search) you may want to limit the size of this search, using one of these commands:

set heuristic-fence-post limit

Restrict [No value for “GDBN”] to examining at most limit bytes in its search for the beginning of a function. A value of 0 (the default) means there is no limit. However, except for 0, the larger the limit the more bytes heuristicfence-post must search and therefore the longer it takes to run.

show heuristic-fence-post

Display the current limit.

These commands are available only when [No value for “GDBN”] is configured for debugging programs on Alpha or MIPS processors.

16 Controlling [No value for “GDBN”]

You can alter the way [No value for “GDBN”] interacts with you by using the set command. For commands controlling how [No value for “GDBN”] displays data, see Section 8.7 [Print settings], page 72. Other settings are described here.

16.1 Prompt

[No value for “GDBN”] indicates its readiness to read a command by printing a string called the prompt. This string is normally ‘([No value for ``GDBP''])’. You can change the prompt string with the set prompt command. For instance, when debugging [No value for “GDBN”] with [No value for “GDBN”], it is useful to change the prompt in one of the [No value for “GDBN”] sessions so that you can always tell which one you are talking to.

Note: set prompt does not add a space for you after the prompt you set. This allows you to set a prompt which ends in a space or a prompt that does not.

set prompt newprompt

Directs [No value for “GDBN”] to use newprompt as its prompt string henceforth.

show prompt

Prints a line of the form: ‘Gdb's prompt is: your-prompt’

16.2 Command editing

[No value for “GDBN”] reads its input commands via the readline interface. This gnu library provides consistent behavior for programs which provide a command line interface to the user. Advantages are gnu Emacs-style or vi-style inline editing of commands, csh-like history substitution, and a storage and recall of command history across debugging sessions.

You may control the behavior of command line editing in [No value for “GDBN”] with the command set.

set editing
set editing on

Enable command line editing (enabled by default).

set editing off

Disable command line editing.

show editing

Show whether command line editing is enabled.

16.3 Command history

[No value for “GDBN”] can keep track of the commands you type during your debugging sessions, so that you can be certain of precisely what happened. Use these commands to manage the [No value for “GDBN”] command history facility.

set history filename fname

Set the name of the [No value for “GDBN”] command history file to fname. This is the file where [No value for “GDBN”] reads an initial command history list, and where it writes the command history from this session when it exits. You can access this list through history expansion or through the history command editing characters listed below. This file defaults to the value of the

environment variable GDBHISTFILE, or to ‘./.gdb_history’ (‘./_gdb_history’

on MS-DOS) if this variable is not set.

set history save
set history save on

Record command history in a file, whose name may be specified with the set history filename command. By default, this option is disabled.

set history save off

Stop recording command history in a file.

set history size size

Set the number of commands which [No value for “GDBN”] keeps in its history list. This defaults to the value of the environment variable HISTSIZE, or to 256 if this variable is not set.

History expansion assigns special meaning to the character !. Since ! is also the logical not operator in C, history expansion is off by default. If you

decide to enable history expansion with the set history expansion on command, you may sometimes need to follow ! (when it is used as logical not, in an expression) with a space or a tab to prevent it from being expanded. The readline history facilities do not attempt substitution on the strings != and !(, even when history expansion is enabled.

The commands to control history expansion are:

set history expansion on
set history expansion

Enable history expansion. History expansion is off by default.

set history expansion off

Disable history expansion. The readline code comes with more complete documentation of editing and history expansion features. Users unfamiliar with gnu Emacs or vi may wish to read it.

show history
show history filename
show history save
show history size
show history expansion

These commands display the state of the [No value for “GDBN”] history parameters. show history by itself displays all four states.

show commands

Display the last ten commands in the command history.

show commands n

Print ten commands centered on command number n.

show commands +

Print ten commands just after the commands last printed.

16.4 Screen size

Certain commands to [No value for “GDBN”] may produce large amounts of information output to the screen. To help you read all of it, [No value for “GDBN”] pauses and asks you for input at the end of each page of output. Type ^{⟨}\underline{\overline{RET}}^{⟩} when you want to continue the output, or q to discard the remaining output. Also, the screen width setting determines when to wrap lines of output. Depending on what is being printed, [No value for “GDBN”] tries to break the line at a readable place, rather than simply letting it overflow onto the following line.

Normally [No value for “GDBN”] knows the size of the screen from the terminal driver software. For example, on Unix [No value for “GDBN”] uses the termcap data base together with the value of the TERM environment variable and the stty rows and stty cols settings. If this is not correct, you can override it with the set height and set width commands:

set height lpp
show height
set width cpl
show width

These set commands specify a screen height of lpp lines and a screen width of cpl characters. The associated show commands display the current settings. If you specify a height of zero lines, [No value for “GDBN”] does not pause during output no matter how long the output is. This is useful if output is to a file or to an editor buffer. Likewise, you can specify ‘set width 0’ to prevent [No value for “GDBN”] from wrapping its output.

16.5 Numbers

You can always enter numbers in octal, decimal, or hexadecimal in [No value for “GDBN”] by the usual conventions: octal numbers begin with ‘0’, decimal numbers end with ‘.’, and hexadecimal numbers begin with ‘0x’. Numbers that begin with none of these are, by default, entered in base 10; likewise, the default display for numbers—when no particular format is specified—is base 10. You can change the default base for both input and output with the set radix command.

set input-radix base

Set the default base for numeric input. Supported choices for base are decimal 8, 10, or 16. base must itself be specified either unambiguously or using the current default radix; for example, any of

               set radix 012
               set radix 10.
               set radix 0xa

sets the base to decimal. On the other hand, ‘set radix 10’ leaves the radix unchanged no matter what it was.

set output-radix base

Set the default base for numeric display. Supported choices for base are decimal 8, 10, or 16. base must itself be specified either unambiguously or using the current default radix.

show input-radix

Display the current default base for numeric input.

show output-radix

Display the current default base for numeric display.

16.6 Optional warnings and messages

By default, [No value for “GDBN”] is silent about its inner workings. If you are running on a slow machine, you may want to use the set verbose command. This makes [No value for “GDBN”] tell you when it does a lengthy internal operation, so you will not think it has crashed.

Currently, the messages controlled by set verbose are those which announce that the symbol table for a source file is being read; see symbol-file in Section 13.1 [Commands to specify files], page 123.

set verbose on

Enables [No value for “GDBN”] output of certain informational messages.

set verbose off

Disables [No value for “GDBN”] output of certain informational messages.

show verbose

Displays whether set verbose is on or off.

By default, if [No value for “GDBN”] encounters bugs in the symbol table of an object file, it is silent; but if you are debugging a compiler, you may find this information useful (see Section 13.2 [Errors reading symbol files], page 127).

set complaints limit

Permits [No value for “GDBN”] to output limit complaints about each type of unusual symbols before becoming silent about the problem. Set limit to zero to suppress all complaints; set it to a large number to prevent complaints from being suppressed.

show complaints

Displays how many symbol complaints [No value for “GDBN”] is permitted to produce.

By default, [No value for “GDBN”] is cautious, and asks what sometimes seems to be a lot of stupid questions to confirm certain commands. For example, if you try to run a program which is already running:

([No value for ``GDBP'']) run
The program being debugged has been started already.
Start it from the beginning? (y or n)

If you are willing to unflinchingly face the consequences of your own commands, you can disable this “feature”:

set confirm off

Disables confirmation requests.

set confirm on

Enables confirmation requests (the default).

show confirm

Displays state of confirmation requests.

16.7 Optional messages about internal happenings

set debug arch

Turns on or off display of gdbarch debugging info. The default is off

show debug arch

Displays the current state of displaying gdbarch debugging info.

set debug event

Turns on or off display of [No value for “GDBN”] event debugging info. The default is off.

show debug event

Displays the current state of displaying [No value for “GDBN”] event debugging info.

set debug expression

Turns on or off display of [No value for “GDBN”] expression debugging info. The default is off.

show debug expression

Displays the current state of displaying [No value for “GDBN”] expression debugging info.

set debug overload

Turns on or off display of [No value for “GDBN”] C++ overload debugging info. This includes info such as ranking of functions, etc. The default is off.

show debug overload

Displays the current state of displaying [No value for “GDBN”] C++ overload debugging info.

set debug remote

Turns on or off display of reports on all packets sent back and forth across the serial line to the remote machine. The info is printed on the [No value for “GDBN”] standard output stream. The default is off.

show debug remote

Displays the state of display of remote packets.

set debug serial

Turns on or off display of [No value for “GDBN”] serial debugging info. The default is off.

show debug serial

Displays the current state of displaying [No value for “GDBN”] serial debugging info.

set debug target

Turns on or off display of [No value for “GDBN”] target debugging info. This info includes what is going on at the target level of GDB, as it happens. The default is off.

show debug target

Displays the current state of displaying [No value for “GDBN”] target debugging info.

set debug varobj

Turns on or off display of [No value for “GDBN”] variable object debugging info. The default is off.

show debug varobj

Displays the current state of displaying [No value for “GDBN”] variable object debugging info.

17 Canned Sequences of Commands

Aside from breakpoint commands (see Section 5.1.7 [Breakpoint command lists], page 43), [No value for “GDBN”] provides two ways to store sequences of commands for execution as a unit: user-defined commands and command files.

17.1 User-defined commands

A user-defined command is a sequence of [No value for “GDBN”] commands to which you assign a new name as a command. This is done with the define command. User commands may accept up to 10 arguments separated by whitespace. Arguments are accessed within the user command via $arg0…$arg9. A trivial example:

define adder
print $arg0 + $arg1 + $arg2

To execute the command use:

adder 1 2 3

This defines the command adder, which prints the sum of its three arguments. Note the arguments are text substitutions, so they may reference variables, use complex expressions, or even perform inferior functions calls.

define commandname Define a command named commandname. If there is already a command by that name, you are asked to confirm that you want to redefine it. The definition of the command is made up of other [No value for “GDBN”] command lines, which are given following the define command. The end of these commands is marked by a line containing end.

if Takes a single argument, which is an expression to evaluate. It is followed by a series of commands that are executed only if the expression is true (nonzero). There can then optionally be a line else, followed by a series of commands that are only executed if the expression was false. The end of the list is marked by a line containing end.

while The syntax is similar to if: the command takes a single argument, which is an expression to evaluate, and must be followed by the commands to execute, one per line, terminated by an end. The commands are executed repeatedly as long as the expression evaluates to true.

document commandname Document the user-defined command commandname, so that it can be accessed by help. The command commandname must already be defined. This command reads lines of documentation just as define reads the lines of the command definition, ending with end. After the document command is finished, help on command commandname displays the documentation you have written. You may use the document command again to change the documentation of a command. Redefining the command with define does not change the documentation.

help user-defined

List all user-defined commands, with the first line of the documentation (if any) for each.

show user

show user commandname Display the [No value for “GDBN”] commands used to define commandname (but not its documentation). If no commandname is given, display the definitions for all user-defined commands.

When user-defined commands are executed, the commands of the definition are not printed. An error in any command stops execution of the user-defined command.

If used interactively, commands that would ask for confirmation proceed without asking when used inside a user-defined command. Many [No value for “GDBN”] commands that normally print messages to say what they are doing omit the messages when used in a user-defined command.

17.2 User-defined command hooks

You may define hooks, which are a special kind of user-defined command. Whenever you run the command ‘foo’, if the user-defined command ‘hook-foo’ exists, it is executed (with no arguments) before that command.

A hook may also be defined which is run after the command you executed. Whenever you run the command ‘foo’, if the user-defined command ‘hookpost-foo’ exists, it is executed (with no arguments) after that command. Post-execution hooks may exist simultaneously with pre-execution hooks, for the same command.

It is valid for a hook to call the command which it hooks. If this occurs, the hook is not re-executed, thereby avoiding infinte recursion.

In addition, a pseudo-command, ‘stop’ exists. Defining (‘hook-stop’) makes the associated commands execute every time execution stops in your program: before breakpoint commands are run, displays are printed, or the stack frame is printed.

For example, to ignore SIGALRM signals while single-stepping, but treat them normally during normal execution, you could define:

     define hook-stop
     handle SIGALRM nopass
     end
     define hook-run
     handle SIGALRM pass
     end
     define hook-continue
     handle SIGLARM pass
     end

As a further example, to hook at the begining and end of the echo command, and to add extra text to the beginning and end of the message, you could define:

     define hook-echo
     echo <<<---
     end
     define hookpost-echo
     echo --->>>\n
     end
     ([No value for ``GDBP'']) echo Hello World
     <<<---Hello World--->>>
     ([No value for ``GDBP''])

You can define a hook for any single-word command in [No value for “GDBN”], but not for command aliases; you should define a hook for the basic command name, e.g. backtrace rather than bt. If an error occurs during the execution of your hook, execution of [No value for “GDBN”] commands stops and [No value for “GDBN”] issues a prompt (before the command that you actually typed had a chance to run).

If you try to define a hook which does not match any known command, you get a warning from the define command.

17.3 Command files

A command file for [No value for “GDBN”] is a file of lines that are [No value for “GDBN”] commands. Comments (lines starting with #) may also be included. An empty line in a command file does nothing; it does not mean to repeat the last command, as it would from the terminal.

When you start [No value for “GDBN”], it automatically executes commands from its init files. These are files named ‘.gdbinit’ on Unix and ‘gdb.ini’ on DOS/Windows. During startup, [No value for “GDBN”] does the following:

1. Reads the init file (if any) in your home directory^{1}.

2. Processes command line options and operands.

3. Reads the init file (if any) in the current working directory.

4. Reads command files specified by the ‘-x’ option.

The init file in your home directory can set options (such as ‘set complaints’) that affect subsequent processing of command line options and operands. Init files are not executed if you use the ‘-nx’ option (see Section 2.1.2 [Choosing modes], page 13).

On some configurations of [No value for “GDBN”], the init file is known by a different name (these are typically environments where a specialized form of [No value for “GDBN”] may need to coexist with other forms, hence a different name for the specialized version’s init file). These are the environments with special init file names:

• VxWorks (Wind River Systems real-time OS): ‘.vxgdbinit’

• OS68K (Enea Data Systems real-time OS): ‘.os68gdbinit’

• ES-1800 (Ericsson Telecom AB M68000 emulator): ‘.esgdbinit’

1 On DOS/Windows systems, the home directory is the one pointed to by the HOME environment variable.

You can also request the execution of a command file with the source command:

source filename Execute the command file filename.

The lines in a command file are executed sequentially. They are not printed as they are executed. An error in any command terminates execution of the command file.

Commands that would ask for confirmation if used interactively proceed without asking when used in a command file. Many [No value for “GDBN”] commands that normally print messages to say what they are doing omit the messages when called from command files.

17.4 Commands for controlled output

During the execution of a command file or a user-defined command, normal [No value for “GDBN”] output is suppressed; the only output that appears is what is explicitly printed by the commands in the definition. This section describes three commands useful for generating exactly the output you want.

echo text Print text. Nonprinting characters can be included in text using C escape sequences, such as ‘\n’ to print a newline. No newline is printed unless you specify one. In addition to the standard C escape sequences, a backslash followed by a space stands for a space. This is useful for displaying a string with spaces at the beginning or the end, since leading and trailing spaces are otherwise trimmed from all arguments. To print ‘ and foo = ’, use the command ‘echo \ and foo

= \ ’.

A backslash at the end of text can be used, as in C, to continue the command onto subsequent lines. For example,

     echo This is some text\n\
     which is continued\n\
     onto several lines.\n

produces the same output as

     echo This is some text\n
     echo which is continued\n
     echo onto several lines.\n

output expression Print the value of expression and nothing but that value: no newlines, no ‘$nn = ’. The value is not entered in the value history either. See Section 8.1 [Expressions], page 65, for more information on expressions.

output/fmt expression Print the value of expression in format fmt. You can use the same formats as for print. See Section 8.4 [Output formats], page 68, for more information.

printf string, expressions...

Print the values of the expressions under the control of string. The expressions are separated by commas and may be either numbers or pointers. Their values are printed as specified by string, exactly as if your program were to execute the C subroutine

printf (string, expressions...); For example, you can print two values in hex like this:

     printf "foo, bar-foo = 0x%x, 0x%x\n", foo, bar-foo

The only backslash-escape sequences that you can use in the format string are the simple ones that consist of backslash followed by a letter.

182 Debugging with [No value for “GDBN”]

18 [No value for “GDBN”] Text User Interface

The [No value for “GDBN”] Text User Interface, TUI in short, is a terminal interface which uses the curses library to show the source file, the assembly output, the program registers and [No value for “GDBN”] commands in separate text windows. The TUI is available only when [No value for “GDBN”] is configured with the --enable-tui configure option (see Section B.3 [Configure Options], page 210).

18.1 TUI overview

The TUI has two display modes that can be switched while [No value for “GDBN”] runs:

• A curses (or TUI) mode in which it displays several text windows on the terminal.

• A standard mode which corresponds to the [No value for “GDBN”] configured without the TUI.

In the TUI mode, [No value for “GDBN”] can display several text window on the terminal:

command This window is the [No value for “GDBN”] command window with the [No value for “GDBN”] prompt and the [No value for “GDBN”] outputs. The [No value for “GDBN”] input is still managed using readline but through the TUI. The command window is always visible.

source The source window shows the source file of the program. The current line as well as active breakpoints are displayed in this window. The current program position is shown with the ‘>’ marker and active breakpoints are shown with ‘*’ markers.

assembly The assembly window shows the disassembly output of the program.

register This window shows the processor registers. It detects when a register is changed and when this is the case, registers that have changed are highlighted.

The source, assembly and register windows are attached to the thread and the frame position. They are updated when the current thread changes, when the frame changes or when the program counter changes. These three windows are arranged by the TUI according to several layouts. The layout defines which of these three windows are visible. The following layouts are available:

• source

• assembly

• source and assembly

• source and registers

• assembly and registers

18.2 TUI Key Bindings

The TUI installs several key bindings in the readline keymaps (see ⟨undefined⟩ [Command Line Editing], page ⟨undefined⟩). They allow to leave or enter in the TUI mode or they operate directly on the TUI layout and windows. The following key bindings are installed for both TUI mode and the [No value for “GDBN”] standard mode.

C-x C-a
C-x a

C-x A Enter or leave the TUI mode. When the TUI mode is left, the curses window management is left and [No value for “GDBN”] operates using its standard mode writing on the terminal directly. When the TUI mode is entered, the control is given back to the curses windows. The screen is then refreshed.

C-x 1 Use a TUI layout with only one window. The layout will either be ‘source’ or ‘assembly’. When the TUI mode is not active, it will switch to the TUI mode. Think of this key binding as the Emacs C-x 1 binding.

C-x 2 Use a TUI layout with at least two windows. When the current layout shows already two windows, a next layout with two windows is used. When a new layout is chosen, one window will always be common to the previous layout and the new one. Think of it as the Emacs C-x 2 binding.

The following key bindings are handled only by the TUI mode:

⟨\underline{\overline{PgUp}}^{⟩} Scroll the active window one page up.

⟨\underline{\overline{PgDn}}^{⟩} Scroll the active window one page down.

⟨\underline{\overline{Up}}^{⟩} Scroll the active window one line up.

⟨\underline{\overline{Down}}^{⟩} Scroll the active window one line down.

⟨\underline{\overline{Left}}^{⟩} Scroll the active window one column left.

⟨\underline{\overline{Right}}^{⟩} Scroll the active window one column right.

⟨\underline{\overline{C-L}}^{⟩} Refresh the screen.

In the TUI mode, the arrow keys are used by the active window for scrolling. This means they are not available for readline. It is necessary to use other readline key bindings such as ^{⟨}\underline{\overline{C-p}}^{⟩}, ^{⟨}\underline{\overline{C-n}}^{⟩}, ^{⟨}\underline{\overline{C-b}}^{⟩} and ^{⟨}\underline{\overline{C-f}}^{⟩}.

18.3 TUI specific commands

The TUI has specific commands to control the text windows. These commands are always available, that is they do not depend on the current terminal mode in which [No value for “GDBN”] runs. When [No value for “GDBN”] is in the standard mode, using these commands will automatically switch in the TUI mode.

layout next

Display the next layout.

layout prev

Display the previous layout.

layout src

Display the source window only.

layout asm

Display the assembly window only.

layout split

Display the source and assembly window.

layout regs

Display the register window together with the source or assembly window.

focus next | prev | src | asm | regs | split

Set the focus to the named window. This command allows to change the active window so that scrolling keys can be affected to another window.

refresh Refresh the screen. This is similar to using ^{⟨}\underline{\overline{C-L}}^{⟩} key.

update Update the source window and the current execution point.

winheight name +count
winheight name -count

Change the height of the window name by count lines. Positive counts increase the height, while negative counts decrease it.

18.4 TUI configuration variables

The TUI has several configuration variables that control the appearance of windows on the terminal.

set tui border-kind kind

Select the border appearance for the source, assembly and register windows. The possible values are the following:

space Use a space character to draw the border.

ascii Use ascii characters + - and | to draw the border.

acs Use the Alternate Character Set to draw the border. The border is drawn using character line graphics if the terminal supports them.

set tui active-border-mode mode

Select the attributes to display the border of the active window. The possible

values are normal, standout, reverse, half, half-standout, bold and bold-
standout.
set tui border-mode mode

Select the attributes to display the border of other windows. The mode can be one of the following:

normal Use normal attributes to display the border.

standout Use standout mode. reverse Use reverse video mode. half Use half bright mode.

half-standout

Use half bright and standout mode.

bold Use extra bright or bold mode.

bold-standout

Use extra bright or bold and standout mode.

19 Using [No value for “GDBN”] under gnu Emacs

A special interface allows you to use gnu Emacs to view (and edit) the source files for the program you are debugging with [No value for “GDBN”].

To use this interface, use the command M-x gdb in Emacs. Give the executable file you want to debug as an argument. This command starts [No value for “GDBN”] as a subprocess of Emacs, with input and output through a newly created Emacs buffer.

Using [No value for “GDBN”] under Emacs is just like using [No value for “GDBN”] normally except for two things:

• All “terminal” input and output goes through the Emacs buffer.

This applies both to [No value for “GDBN”] commands and their output, and to the input and output done by the program you are debugging.

This is useful because it means that you can copy the text of previous commands and input them again; you can even use parts of the output in this way.

All the facilities of Emacs’ Shell mode are available for interacting with your program. In particular, you can send signals the usual way—for example, C-c C-c for an interrupt, C-c C-z for a stop.

• [No value for “GDBN”] displays source code through Emacs.

Each time [No value for “GDBN”] displays a stack frame, Emacs automatically finds the source file for that frame and puts an arrow (‘=>’) at the left margin of the current line. Emacs uses a separate buffer for source display, and splits the screen to show both your [No value for “GDBN”] session and the source.

Explicit [No value for “GDBN”] list or search commands still produce output as usual, but you probably have no reason to use them from Emacs.

Warning: If the directory where your program resides is not your current directory, it can be easy to confuse Emacs about the location of the source files, in which case the auxiliary display buffer does not appear to show your source. [No value for “GDBN”] can find programs by searching your environment’s PATH variable, so the [No value for “GDBN”] input and output session proceeds normally; but Emacs does not get enough information back from [No value for “GDBN”] to locate the source files in this situation. To avoid this problem, either start [No value for “GDBN”] mode from the directory where your program resides, or specify an absolute file name when prompted for the M-x gdb argument. A similar confusion can result if you use the [No value for “GDBN”] file command to switch to debugging a program in some other location, from an existing [No value for “GDBN”] buffer in Emacs.

By default, M-x gdb calls the program called ‘gdb’. If you need to call [No value for “GDBN”] by a different name (for example, if you keep several configurations around, with different names) you can set the Emacs variable gdb-command-name; for example,

     (setq gdb-command-name "mygdb")

(preceded by M-: or ESC :, or typed in the *scratch* buffer, or in your ‘.emacs’ file) makes Emacs call the program named “mygdb” instead.

In the [No value for “GDBN”] I/O buffer, you can use these special Emacs commands in addition to the standard Shell mode commands:

C-h m Describe the features of Emacs’ [No value for “GDBN”] Mode.

M-s Execute to another source line, like the [No value for “GDBN”] step command; also update the display window to show the current file and location.

M-n Execute to next source line in this function, skipping all function calls, like the [No value for “GDBN”] next command. Then update the display window to show the current file and location.

M-i Execute one instruction, like the [No value for “GDBN”] stepi command; update display window accordingly.

M-x gdb-nexti

Execute to next instruction, using the [No value for “GDBN”] nexti command; update display window accordingly.

C-c C-f Execute until exit from the selected stack frame, like the [No value for “GDBN”] finish command.

M-c Continue execution of your program, like the [No value for “GDBN”] continue command. Warning: In Emacs v19, this command is C-c C-p.

M-u Go up the number of frames indicated by the numeric argument (see section “Numeric Arguments” in The gnu Emacs Manual), like the [No value for “GDBN”] up command. Warning: In Emacs v19, this command is C-c C-u.

M-d Go down the number of frames indicated by the numeric argument, like the [No value for “GDBN”] down command. Warning: In Emacs v19, this command is C-c C-d.

C-x & Read the number where the cursor is positioned, and insert it at the end of the [No value for “GDBN”] I/O buffer. For example, if you wish to disassemble code around an address that was displayed earlier, type disassemble; then move the cursor to the address display, and pick up the argument for disassemble by typing C-x &. You can customize this further by defining elements of the list gdb-print-command; once it is defined, you can format or otherwise process numbers picked up by C-x & before they are inserted. A numeric argument to C-x & indicates that you wish special formatting, and also acts as an index to pick an element of the list. If the list element is a string, the number to be inserted is formatted using the Emacs function format; otherwise the number is passed as an argument to the corresponding list element.

In any source file, the Emacs command C-x SPC (gdb-break) tells [No value for “GDBN”] to set a breakpoint on the source line point is on.

If you accidentally delete the source-display buffer, an easy way to get it back is to type the command f in the [No value for “GDBN”] buffer, to request a frame display; when you run under Emacs, this recreates the source buffer if necessary to show you the context of the current frame.

The source files displayed in Emacs are in ordinary Emacs buffers which are visiting the source files in the usual way. You can edit the files with these buffers if you wish; but keep in mind that [No value for “GDBN”] communicates with Emacs in terms of line numbers. If you add or delete lines from the text, the line numbers that [No value for “GDBN”] knows cease to correspond properly with the code.

190 Debugging with [No value for “GDBN”]

20 [No value for “GDBN”] Annotations

This chapter describes annotations in [No value for “GDBN”]. Annotations are designed to interface [No value for “GDBN”] to graphical user interfaces or other similar programs which want to interact with [No value for “GDBN”] at a relatively high level.

20.1 What is an Annotation?

To produce annotations, start [No value for “GDBN”] with the --annotate=2 option. Annotations start with a newline character, two ‘control-z’ characters, and the name

of the annotation. If there is no additional information associated with this annotation, the name of the annotation is followed immediately by a newline. If there is additional information, the name of the annotation is followed by a space, the additional information, and a newline. The additional information cannot contain newline characters.

Any output not beginning with a newline and two ‘control-z’ characters denotes literal output from [No value for “GDBN”]. Currently there is no need for [No value for “GDBN”] to output a newline followed by two ‘control-z’ characters, but if there was such a need, the annotations could be extended with an ‘escape’ annotation which means those three characters as output.

A simple example of starting up [No value for “GDBN”] with annotations is:

$ gdb --annotate=2
GNU GDB 5.0
Copyright 2000 Free Software Foundation, Inc.
GDB is free software, covered by the GNU General Public License,
and you are welcome to change it and/or distribute copies of it
under certain conditions.
Type "show copying" to see the conditions.
There is absolutely no warranty for GDB.  Type "show warranty"
for details.
This GDB was configured as "sparc-sun-sunos4.1.3"
^Z^Zpre-prompt
(gdb)
^Z^Zprompt
quit
^Z^Zpost-prompt
$

Here ‘quit’ is input to [No value for “GDBN”]; the rest is output from [No value for “GDBN”]. The three lines beginning ‘^Z^Z’ (where ‘^Z’ denotes a ‘control-z’ character) are annotations; the rest is output from [No value for “GDBN”].

20.2 The Server Prefix

To issue a command to [No value for “GDBN”] without affecting certain aspects of the state which is seen by users, prefix it with ‘server ’. This means that this command will not affect the command history, nor will it affect [No value for “GDBN”]’s notion of which command to repeat if ^{⟨}\underline{\overline{RET}}^{⟩} is pressed on a line by itself.

The server prefix does not affect the recording of values into the value history; to print a value without recording it into the value history, use the output command instead of the print command.

20.3 Values

When a value is printed in various contexts, [No value for “GDBN”] uses annotations to delimit the value from the surrounding text.

If a value is printed using print and added to the value history, the annotation looks like

^Z^Zvalue-history-begin history-number value-flags history-string

^Z^Zvalue-history-value

the-value

^Z^Zvalue-history-end

where history-number is the number it is getting in the value history, history-string is a string, such as ‘$5 = ’, which introduces the value to the user, the-value is the output corresponding to the value itself, and value-flags is ‘*’ for a value which can be dereferenced and ‘-’ for a value which cannot.

If the value is not added to the value history (it is an invalid float or it is printed with the output command), the annotation is similar:

^Z^Zvalue-begin value-flags

the-value

^Z^Zvalue-end

When [No value for “GDBN”] prints an argument to a function (for example, in the output from the backtrace command), it annotates it as follows:

^Z^Zarg-begin

argument-name

^Z^Zarg-name-end

separator-string

^Z^Zarg-value value-flags

the-value

^Z^Zarg-end

where argument-name is the name of the argument, separator-string is text which separates the name from the value for the user’s benefit (such as ‘=’), and value-flags and the-value have the same meanings as in a value-history-begin annotation.

When printing a structure, [No value for “GDBN”] annotates it as follows:

^Z^Zfield-begin value-flags

field-name

^Z^Zfield-name-end

separator-string

^Z^Zfield-value

the-value

     ^Z^Zfield-end

where field-name is the name of the field, separator-string is text which separates the name from the value for the user’s benefit (such as ‘=’), and value-flags and the-value have the same meanings as in a value-history-begin annotation.

When printing an array, [No value for “GDBN”] annotates it as follows:

     ^Z^Zarray-section-begin array-index value-flags

where array-index is the index of the first element being annotated and value-flags has the same meaning as in a value-history-begin annotation. This is followed by any number of elements, where is element can be either a single element:

`,' whitespace ; omitted for the first element the-value

     ^Z^Zelt

or a repeated element `,' whitespace ; omitted for the first element the-value ^Z^Zelt-rep number-of-repititions repetition-string

     ^Z^Zelt-rep-end

In both cases, the-value is the output for the value of the element and whitespace can contain spaces, tabs, and newlines. In the repeated case, number-of-repititons is the number of consecutive array elements which contain that value, and repetition-string is a string which is designed to convey to the user that repitition is being depicted.

Once all the array elements have been output, the array annotation is ended with

     ^Z^Zarray-section-end

20.4 Frames

Whenever [No value for “GDBN”] prints a frame, it annotates it. For example, this applies to frames printed when [No value for “GDBN”] stops, output from commands such

as backtrace or up, etc.

The frame annotation begins with

     ^Z^Zframe-begin level address

level-string where level is the number of the frame (0 is the innermost frame, and other frames have positive numbers), address is the address of the code executing in that frame, and levelstring is a string designed to convey the level to the user. address is in the form ‘0x’ followed by one or more lowercase hex digits (note that this does not depend on the language). The frame ends with

     ^Z^Zframe-end

Between these annotations is the main body of the frame, which can consist of •

         ^Z^Zfunction-call

function-call-string where function-call-string is text designed to convey to the user that this frame is associated with a function call made by [No value for “GDBN”] to a function in the program being debugged.

•

     ^Z^Zsignal-handler-caller

signal-handler-caller-string

where signal-handler-caller-string is text designed to convey to the user that this frame is associated with whatever mechanism is used by this operating system to call a signal handler (it is the frame which calls the signal handler, not the frame for the signal handler itself).

• A normal frame.

This can optionally (depending on whether this is thought of as interesting information for the user to see) begin with

     ^Z^Zframe-address

address

     ^Z^Zframe-address-end

separator-string

where address is the address executing in the frame (the same address as in the framebegin annotation, but printed in a form which is intended for user consumption—in particular, the syntax varies depending on the language), and separator-string is a string intended to separate this address from what follows for the user’s benefit.

Then comes

     ^Z^Zframe-function-name

function-name

     ^Z^Zframe-args

arguments

where function-name is the name of the function executing in the frame, or ‘??’ if not known, and arguments are the arguments to the frame, with parentheses around them (each argument is annotated individually as well, see Section 20.3 [Value Annotations], page 192).

If source information is available, a reference to it is then printed:

     ^Z^Zframe-source-begin

source-intro-string

     ^Z^Zframe-source-file

filename

     ^Z^Zframe-source-file-end
     :
     ^Z^Zframe-source-line

line-number

     ^Z^Zframe-source-end

where source-intro-string separates for the user’s benefit the reference from the text which precedes it, filename is the name of the source file, and line-number is the line number within that file (the first line is line 1).

If [No value for “GDBN”] prints some information about where the frame is from (which library, which load segment, etc.; currently only done on the RS/6000), it is annotated with

    ^Z^Zframe-where

information Then, if source is to actually be displayed for this frame (for example, this is not true for output from the backtrace command), then a source annotation (see Section 20.11 [Source Annotations], page 198) is displayed. Unlike most annotations, this is output instead of the normal text which would be output, not in addition.

20.5 Displays

When [No value for “GDBN”] is told to display something using the display command, the results of the display are annotated:

^Z^Zdisplay-begin

number

^Z^Zdisplay-number-end

number-separator

^Z^Zdisplay-format

format

^Z^Zdisplay-expression

expression

^Z^Zdisplay-expression-end

expression-separator

^Z^Zdisplay-value

value

^Z^Zdisplay-end

where number is the number of the display, number-separator is intended to separate the number from what follows for the user, format includes information such as the size, format, or other information about how the value is being displayed, expression is the expression being displayed, expression-separator is intended to separate the expression from the text that follows for the user, and value is the actual value being displayed.

20.6 Annotation for [No value for “GDBN”] Input

When [No value for “GDBN”] prompts for input, it annotates this fact so it is possible to know when to send output, when the output from a given command is over, etc.

Different kinds of input each have a different input type. Each input type has three annotations: a pre- annotation, which denotes the beginning of any prompt which is being output, a plain annotation, which denotes the end of the prompt, and then a post- annotation which denotes the end of any echo which may (or may not) be associated with the input. For example, the prompt input type features the following annotations:

^Z^Zpre-prompt
^Z^Zprompt
^Z^Zpost-prompt

The input types are prompt When [No value for “GDBN”] is prompting for a command (the main [No value

for “GDBN”] prompt).

commands When [No value for “GDBN”] prompts for a set of commands, like in the commands command. The annotations are repeated for each command which is input.

overload-choice

When [No value for “GDBN”] wants the user to select between various overloaded functions.

query When [No value for “GDBN”] wants the user to confirm a potentially dangerous operation.

prompt-for-continue

When [No value for “GDBN”] is asking the user to press return to continue. Note: Don’t expect this to work well; instead use set height 0 to disable prompting. This is because the counting of lines is buggy in the presence of annotations.

20.7 Errors

     ^Z^Zquit

This annotation occurs right before [No value for “GDBN”] responds to an interrupt.

     ^Z^Zerror

This annotation occurs right before [No value for “GDBN”] responds to an error.

Quit and error annotations indicate that any annotations which [No value for “GDBN”] was in the middle of may end abruptly. For example, if a value-history-begin annotation is followed by a error, one cannot expect to receive the matching value-history-end. One cannot expect not to receive it either, however; an error annotation does not necessarily mean that [No value for “GDBN”] is immediately returning all the way to the top level.

A quit or error annotation may be preceded by

     ^Z^Zerror-begin

Any output between that and the quit or error annotation is the error message.

Warning messages are not yet annotated.

20.8 Information on Breakpoints

The output from the info breakpoints command is annotated as follows:

     ^Z^Zbreakpoints-headers

header-entry

     ^Z^Zbreakpoints-table

where header-entry has the same syntax as an entry (see below) but instead of containing data, it contains strings which are intended to convey the meaning of each field to the user. This is followed by any number of entries. If a field does not apply for this entry, it is omitted. Fields may contain trailing whitespace. Each entry consists of:

^Z^Zrecord
^Z^Zfield 0

number

^Z^Zfield 1

type

^Z^Zfield 2

disposition

^Z^Zfield 3

enable

^Z^Zfield 4

address

^Z^Zfield 5

what

^Z^Zfield 6

frame

^Z^Zfield 7

condition

^Z^Zfield 8

ignore-count

^Z^Zfield 9

commands Note that address is intended for user consumption—the syntax varies depending on the

language.

The output ends with

^Z^Zbreakpoints-table-end

20.9 Invalidation Notices

The following annotations say that certain pieces of state may have changed.

^Z^Zframes-invalid

The frames (for example, output from the backtrace command) may have changed.

^Z^Zbreakpoints-invalid

The breakpoints may have changed. For example, the user just added or deleted a breakpoint.

20.10 Running the Program

When the program starts executing due to a [No value for “GDBN”] command such as

step or continue,
^Z^Zstarting

is output. When the program stops,

^Z^Zstopped

is output. Before the stopped annotation, a variety of annotations describe how the program stopped.

^Z^Zexited exit-status The program exited, and exit-status is the exit status (zero for successful exit, otherwise nonzero).

^Z^Zsignalled

The program exited with a signal. After the ^Z^Zsignalled, the annotation continues:

intro-text

               ^Z^Zsignal-name

name

               ^Z^Zsignal-name-end

middle-text

               ^Z^Zsignal-string

string

               ^Z^Zsignal-string-end

end-text where name is the name of the signal, such as SIGILL or SIGSEGV, and string is the explanation of the signal, such as Illegal Instruction or Segmentation fault. intro-text, middle-text, and end-text are for the user’s benefit and have no particular format.

^Z^Zsignal

The syntax of this annotation is just like signalled, but [No value for “GDBN”] is just saying that the program received the signal, not that it was terminated with it.

^Z^Zbreakpoint number

The program hit breakpoint number number.

^Z^Zwatchpoint number

The program hit watchpoint number number.

20.11 Displaying Source

The following annotation is used instead of displaying source code:

^Z^Zsource filename:line:character:middle:addr where filename is an absolute file name indicating which source file, line is the line

number within that file (where 1 is the first line in the file), character is the character position within the file (where 0 is the first character in the file) (for most debug formats this will necessarily point to the beginning of a line), middle is ‘middle’ if addr is in the middle of the line, or ‘beg’ if addr is at the beginning of the line, and addr is the address in the target program associated with the source which is being displayed. addr is in the form ‘0x’ followed by one or more lowercase hex digits (note that this does not depend on the language).

20.12 Annotations We Might Want in the Future

- target-invalid the target might have changed (registers, heap contents, or execution status). For performance, we might eventually want to hit ‘registers-invalid’ and ‘all-registers-invalid’ with greater precision

- systematic annotation for set/show parameters (including invalidation notices).

- similarly, ‘info’ returns a list of candidates for invalidation notices. %% %% This is file ‘.tex’, %% generated with the docstrip utility. %% %% The original

source files were: %% %% fileerr.dtx (with options: ‘return’) %% %% This is a generated file. %% %% Copyright 1993 1994 1995 1996 1997 1998 1999 %% The LaTeX3 Project and any individual authors listed elsewhere %% in this file. %% %% This file is part of the Standard LaTeX ‘Tools Bundle’. %% ——————————————————- %% %% This file, and all the other files in this bundle as listed below, %% may be distributed under the terms of the LaTeX Project Public %% License, as described in lppl.txt in the base LaTeX distribution. %% Either version 1.0 or, at your option, any later version. %% \messageFile ignored \endinput %% %% End of file ‘.tex’.

200 Debugging with [No value for “GDBN”]

21 Reporting Bugs in [No value for “GDBN”]

Your bug reports play an essential role in making [No value for “GDBN”] reliable.

Reporting a bug may help you by bringing a solution to your problem, or it may not. But in any case the principal function of a bug report is to help the entire community by making the next version of [No value for “GDBN”] work better. Bug reports are your contribution to the maintenance of [No value for “GDBN”].

In order for a bug report to serve its purpose, you must include the information that enables us to fix the bug.

21.1 Have you found a bug?

If you are not sure whether you have found a bug, here are some guidelines:

• If the debugger gets a fatal signal, for any input whatever, that is a [No value for “GDBN”] bug. Reliable debuggers never crash.

• If [No value for “GDBN”] produces an error message for valid input, that is a bug. (Note that if you’re cross debugging, the problem may also be somewhere in the connection to the target.)

• If [No value for “GDBN”] does not produce an error message for invalid input, that is a bug. However, you should note that your idea of “invalid input” might be our idea of “an extension” or “support for traditional practice”.

• If you are an experienced user of debugging tools, your suggestions for improvement of [No value for “GDBN”] are welcome in any case.

21.2 How to report bugs

A number of companies and individuals offer support for gnu products. If you obtained [No value for “GDBN”] from a support organization, we recommend you contact that organization first.

You can find contact information for many support companies and individuals in the file ‘etc/SERVICE’ in the gnu Emacs distribution.

In any event, we also recommend that you send bug reports for [No value for “GDBN”] to this addresses:

     bug-gdb@gnu.org

Do not send bug reports to ‘info-gdb’, or to ‘help-gdb’, or to any newsgroups. Most users of [No value for “GDBN”] do not want to receive bug reports. Those that do have arranged to receive ‘bug-gdb’.

The mailing list ‘bug-gdb’ has a newsgroup ‘gnu.gdb.bug’ which serves as a repeater. The mailing list and the newsgroup carry exactly the same messages. Often people think of posting bug reports to the newsgroup instead of mailing them. This appears to work, but it has one problem which can be crucial: a newsgroup posting often lacks a mail path back to the sender. Thus, if we need to ask for more information, we may be unable to reach you. For this reason, it is better to send bug reports to the mailing list.

As a last resort, send bug reports on paper to:

 gnu Debugger Bugs
 Free Software Foundation Inc.
 59 Temple Place - Suite 330
 Boston, MA 02111-1307
 USA

The fundamental principle of reporting bugs usefully is this: report all the facts. If you are not sure whether to state a fact or leave it out, state it!

Often people omit facts because they think they know what causes the problem and assume that some details do not matter. Thus, you might assume that the name of the variable you use in an example does not matter. Well, probably it does not, but one cannot be sure. Perhaps the bug is a stray memory reference which happens to fetch from the location where that name is stored in memory; perhaps, if the name were different, the contents of that location would fool the debugger into doing the right thing despite the bug. Play it safe and give a specific, complete example. That is the easiest thing for you to do, and the most helpful.

Keep in mind that the purpose of a bug report is to enable us to fix the bug. It may be that the bug has been reported previously, but neither you nor we can know that unless your bug report is complete and self-contained.

Sometimes people give a few sketchy facts and ask, “Does this ring a bell?” Those bug reports are useless, and we urge everyone to refuse to respond to them except to chide the sender to report bugs properly.

To enable us to fix the bug, you should include all these things:

• The version of [No value for “GDBN”]. [No value for “GDBN”] announces it if you start with no arguments; you can also print it at any time using show version. Without this, we will not know whether there is any point in looking for the bug in the current version of [No value for “GDBN”].

• The type of machine you are using, and the operating system name and version number.

• What compiler (and its version) was used to compile [No value for “GDBN”]—e.g. “[No value for “GCC”]–2.8.1”.

• What compiler (and its version) was used to compile the program you are debugging— e.g. “[No value for “GCC”]–2.8.1”, or “HP92453-01 A.10.32.03 HP C Compiler”. For GCC, you can say gcc --version to get this information; for other compilers, see the documentation for those compilers.

• The command arguments you gave the compiler to compile your example and observe the bug. For example, did you use ‘-O’? To guarantee you will not omit something important, list them all. A copy of the Makefile (or the output from make) is sufficient. If we were to try to guess the arguments, we would probably guess wrong and then we might not encounter the bug.

• A complete input script, and all necessary source files, that will reproduce the bug.

• A description of what behavior you observe that you believe is incorrect. For example, “It gets a fatal signal.” Of course, if the bug is that [No value for “GDBN”] gets a fatal signal, then we will certainly notice it. But if the bug is incorrect output, we might not notice unless it is glaringly wrong. You might as well not give us a chance to make a mistake.

Even if the problem you experience is a fatal signal, you should still say so explicitly. Suppose something strange is going on, such as, your copy of [No value for “GDBN”] is out of synch, or you have encountered a bug in the C library on your system. (This has happened!) Your copy might crash and ours would not. If you told us to expect a crash, then when ours fails to crash, we would know that the bug was not happening for us. If you had not told us to expect a crash, then we would not be able to draw any conclusion from our observations.

• If you wish to suggest changes to the [No value for “GDBN”] source, send us context diffs. If you even discuss something in the [No value for “GDBN”] source, refer to it by context, not by line number. The line numbers in our development sources will not match those in your sources. Your line numbers would convey no useful information to us.

Here are some things that are not necessary:

• A description of the envelope of the bug.

Often people who encounter a bug spend a lot of time investigating which changes to the input file will make the bug go away and which changes will not affect it. This is often time consuming and not very useful, because the way we will find the bug is by running a single example under the debugger with breakpoints, not by pure deduction from a series of examples. We recommend that you save your time for something else. Of course, if you can find a simpler example to report instead of the original one, that is a convenience for us. Errors in the output will be easier to spot, running under the debugger will take less time, and so on. However, simplification is not vital; if you do not want to do this, report the bug anyway and send us the entire test case you used.

• A patch for the bug.

A patch for the bug does help us if it is a good one. But do not omit the necessary information, such as the test case, on the assumption that a patch is all we need. We might see problems with your patch and decide to fix the problem another way, or we might not understand it at all. Sometimes with a program as complicated as [No value for “GDBN”] it is very hard to construct an example that will make the program follow a certain path through the code. If you do not send us the example, we will not be able to construct one, so we will not be able to verify that the bug is fixed. And if we cannot understand what bug you are trying to fix, or why your patch should be an improvement, we will not install it. A test case will help us to understand.

• A guess about what the bug is or what it depends on.

Such guesses are usually wrong. Even we cannot guess right about such things without first using the debugger to find the facts.

%% %% This is file ‘.tex’, %% generated with the docstrip utility. %% %% The original source files were: %% %% fileerr.dtx (with options: ‘return’) %% %% This is a generated file. %% %% Copyright 1993 1994 1995 1996 1997 1998 1999 %% The LaTeX3 Project and any individual authors listed elsewhere %% in this file. %% %% This file is part of the Standard LaTeX ‘Tools Bundle’. %% ——————————————————- %% %% This file, and all the other files in this bundle as listed below, %% may be distributed under the terms of the LaTeX Project Public %% License, as described in lppl.txt in the base LaTeX distribution. %% Either version 1.0 or, at your option, any later version. %% \messageFile ignored \endinput %% %% End of file ‘.tex’. %% %% This is file ‘.tex’, %% generated with the docstrip utility. %% %% The original source files were: %% %% fileerr.dtx (with options: ‘return’) %% %% This is a generated file. %% %% Copyright 1993 1994 1995 1996 1997 1998 1999 %% The LaTeX3 Project and any individual authors listed elsewhere %% in this file. %% %% This file is part of the Standard LaTeX ‘Tools Bundle’. %% ——————————————————- %% %% This file, and all the other files in this bundle as listed below, %% may be distributed under the terms of the LaTeX Project Public %% License, as described in lppl.txt in the base LaTeX distribution. %% Either version 1.0 or, at your option, any later version. %% \messageFile ignored \endinput %% %% End of file ‘.tex’.

Appendix A Formatting Documentation

The [No value for “GDBN”] 4 release includes an already-formatted reference card, ready for printing with PostScript or Ghostscript, in the ‘gdb’ subdirectory of the main source directory^{1}. If you can use PostScript or Ghostscript with your printer, you can print the reference card immediately with ‘refcard.ps’.

The release also includes the source for the reference card. You can format it, using TEX, by typing:

     make refcard.dvi

The [No value for “GDBN”] reference card is designed to print in landscape mode on US “letter” size paper; that is, on a sheet 11 inches wide by 8.5 inches high. You will need to specify this form of printing as an option to your dvi output program.

All the documentation for [No value for “GDBN”] comes as part of the machine-readable distribution. The documentation is written in Texinfo format, which is a documentation system that uses a single source file to produce both on-line information and a printed manual. You can use one of the Info formatting commands to create the on-line version of the documentation and TEX (or texi2roff) to typeset the printed version.

[No value for “GDBN”] includes an already formatted copy of the on-line Info version of this manual in the ‘gdb’ subdirectory. The main Info file is ‘gdb-[No value for ``GDBVN'']/gdb/gdb.info’, and it refers to subordinate files matching ‘gdb.info*’ in the same directory. If necessary, you can print out these files, or read them with any editor; but they are easier to read using the info subsystem in gnu Emacs or the standalone info program, available as part of the gnu Texinfo distribution.

If you want to format these Info files yourself, you need one of the Info formatting

programs, such as texinfo-format-buffer or makeinfo.

If you have makeinfo installed, and are in the top level [No value for “GDBN”] source directory (‘gdb-[No value for ``GDBVN'']’, in the case of version [No value for “GDBVN”]), you can make the Info file by typing:

     cd gdb
     make gdb.info

If you want to typeset and print copies of this manual, you need TEX, a program to print its dvi output files, and ‘texinfo.tex’, the Texinfo definitions file.

TEX is a typesetting program; it does not print files directly, but produces output files called dvi files. To print a typeset document, you need a program to print dvi files. If your system has TEX installed, chances are it has such a program. The precise command to use depends on your system; lpr -d is common; another (for PostScript devices) is dvips. The dvi print command may require a file name without any extension or a ‘.dvi’ extension.

TEX also requires a macro definitions file called ‘texinfo.tex’. This file tells TEX how to typeset a document written in Texinfo format. On its own, TEX cannot either read or typeset a Texinfo file. ‘texinfo.tex’ is distributed with GDB and is located in the ‘gdb-version-number/texinfo’ directory.

1 In ‘gdb-[No value for ``GDBVN'']/gdb/refcard.ps’ of the version [No value for “GDBVN”] release.

If you have TEX and a dvi printer program installed, you can typeset and print this manual. First switch to the the ‘gdb’ subdirectory of the main source directory (for example,

to ‘gdb-[No value for ``GDBVN'']/gdb’) and type:
     make gdb.dvi

Then give ‘gdb.dvi’ to your dvi printing program.

Appendix B Installing [No value for “GDBN”]

[No value for “GDBN”] comes with a configure script that automates the process of preparing [No value for “GDBN”] for installation; you can then use make to build the gdb program.^{1}

The [No value for “GDBN”] distribution includes all the source code you need for [No value for “GDBN”] in a single directory, whose name is usually composed by appending the version number to ‘gdb’.

For example, the [No value for “GDBN”] version [No value for “GDBVN”] distribution is in the ‘gdb-[No value for ``GDBVN'']’ directory. That directory contains:

gdb-[No value for ``GDBVN'']/configure (and supporting files)

script for configuring [No value for “GDBN”] and all its supporting libraries

gdb-[No value for ``GDBVN'']/gdb

the source specific to [No value for “GDBN”] itself

gdb-[No value for ``GDBVN'']/bfd

source for the Binary File Descriptor library

gdb-[No value for ``GDBVN'']/include

gnu include files

gdb-[No value for ``GDBVN'']/libiberty

source for the ‘-liberty’ free software library

gdb-[No value for ``GDBVN'']/opcodes

source for the library of opcode tables and disassemblers

gdb-[No value for ``GDBVN'']/readline

source for the gnu command-line interface

gdb-[No value for ``GDBVN'']/glob

source for the gnu filename pattern-matching subroutine

gdb-[No value for ``GDBVN'']/mmalloc

source for the gnu memory-mapped malloc package

The simplest way to configure and build [No value for “GDBN”] is to run configure from the ‘gdb-version-number’ source directory, which in this example is the ‘gdb-[No

value for ``GDBVN'']’ directory.

First switch to the ‘gdb-version-number’ source directory if you are not already in it; then run configure. Pass the identifier for the platform on which [No value for “GDBN”] will run as an argument.

For example:

     cd gdb-[No value for ``GDBVN'']
     ./configure host
     make

1 If you have a more recent version of [No value for “GDBN”] than [No value for “GDBVN”], look at

the ‘README’ file in the sources; we may have improved the installation procedures since publishing this

manual.

where host is an identifier such as ‘sun4’ or ‘decstation’, that identifies the platform where [No value for “GDBN”] will run. (You can often leave off host; configure tries to guess the correct value by examining your system.)

Running ‘configure host’ and then running make builds the ‘bfd’, ‘readline’, ‘mmalloc’, and ‘libiberty’ libraries, then gdb itself. The configured source files, and the binaries, are left in the corresponding source directories.

configure is a Bourne-shell (/bin/sh) script; if your system does not recognize this automatically when you run a different shell, you may need to run sh on it explicitly:

     sh configure host

If you run configure from a directory that contains source directories for multiple libraries or programs, such as the ‘gdb-[No value for ``GDBVN'']’ source directory for version [No value for “GDBVN”], configure creates configuration files for every directory level underneath (unless you tell it not to, with the ‘--norecursion’ option).

You can run the configure script from any of the subordinate directories in the [No value for “GDBN”] distribution if you only want to configure that subdirectory, but be sure to specify a path to it.

For example, with version [No value for “GDBVN”], type the following to configure only the bfd subdirectory:

     cd gdb-[No value for ``GDBVN'']/bfd
     ../configure host

You can install [No value for ``GDBP''] anywhere; it has no hardwired paths. However, you should make sure that the shell on your path (named by the ‘SHELL’ environment variable) is publicly readable. Remember that [No value for “GDBN”] uses the shell to start your program—some systems refuse to let [No value for “GDBN”] debug child processes whose programs are not readable.

B.1 Compiling [No value for “GDBN”] in another directory

If you want to run [No value for “GDBN”] versions for several host or target machines, you need a different gdb compiled for each combination of host and target. configure is designed to make this easy by allowing you to generate each configuration in a separate subdirectory, rather than in the source directory. If your make program handles the ‘VPATH’ feature (gnu make does), running make in each of these directories builds the gdb program specified there.

To build gdb in a separate directory, run configure with the ‘--srcdir’ option to specify where to find the source. (You also need to specify a path to find configure itself from your working directory. If the path to configure would be the same as the argument to ‘--srcdir’, you can leave out the ‘--srcdir’ option; it is assumed.)

For example, with version [No value for “GDBVN”], you can build [No value for “GDBN”] in a separate directory for a Sun 4 like this:

     cd gdb-[No value for ``GDBVN'']
     mkdir ../gdb-sun4
     cd ../gdb-sun4
     ../gdb-[No value for ``GDBVN'']/configure sun4
     make

When configure builds a configuration using a remote source directory, it creates a tree for the binaries with the same structure (and using the same names) as the tree under the source directory. In the example, you’d find the Sun 4 library ‘libiberty.a’ in the directory ‘gdb-sun4/libiberty’, and [No value for “GDBN”] itself in ‘gdb-sun4/gdb’.

One popular reason to build several [No value for “GDBN”] configurations in separate directories is to configure [No value for “GDBN”] for cross-compiling (where [No value for “GDBN”] runs on one machine—the host—while debugging programs that run on another machine—the target). You specify a cross-debugging target by giving the ‘--target=target’

option to configure.

When you run make to build a program or library, you must run it in a configured directory—whatever directory you were in when you called configure (or one of its subdirectories).

The Makefile that configure generates in each source directory also runs recursively. If you type make in a source directory such as ‘gdb-[No value for ``GDBVN'']’ (or in a separate configured directory configured with ‘--srcdir=dirname/gdb-[No value for ``GDBVN'']’), you will build all the required libraries, and then build GDB.

When you have multiple hosts or targets configured in separate directories, you can run make on them in parallel (for example, if they are NFS-mounted on each of the hosts); they will not interfere with each other.

B.2 Specifying names for hosts and targets

The specifications used for hosts and targets in the configure script are based on a three-part naming scheme, but some short predefined aliases are also supported. The full naming scheme encodes three pieces of information in the following pattern:

architecture-vendor-os For example, you can use the alias sun4 as a host argument, or as the value for target

in a --target=target option. The equivalent full name is ‘sparc-sun-sunos4’.

The configure script accompanying [No value for “GDBN”] does not provide any query facility to list all supported host and target names or aliases. configure calls the Bourne shell script config.sub to map abbreviations to full names; you can read the script, if you wish, or you can use it to test your guesses on abbreviations—for example:

     % sh config.sub i386-linux
     i386-pc-linux-gnu
     % sh config.sub alpha-linux
     alpha-unknown-linux-gnu
     % sh config.sub hp9k700
     hppa1.1-hp-hpux
     % sh config.sub sun4
     sparc-sun-sunos4.1.1
     % sh config.sub sun3
     m68k-sun-sunos4.1.1
     % sh config.sub i986v
     Invalid configuration `i986v': machine `i986v' not recognized

config.sub is also distributed in the [No value for “GDBN”] source directory (‘gdb-[No value for ``GDBVN'']’, for version [No value for “GDBVN”]).

B.3 configure options

Here is a summary of the configure options and arguments that are most often useful for building [No value for “GDBN”]. configure also has several other options not listed here. See Info file ‘configure.info’, node ‘What Configure Does’, for a full explanation

of configure.
configure [--help]
     [--prefix=dir]
     [--exec-prefix=dir]
     [--srcdir=dirname]
     [--norecursion] [--rm]
     [--target=target]

host

You may introduce options with a single ‘-’ rather than ‘--’ if you prefer; but you may abbreviate option names if you use ‘--’.

--help Display a quick summary of how to invoke configure.

--prefix=dir

Configure the source to install programs and files under directory ‘dir’.

--exec-prefix=dir

Configure the source to install programs under directory ‘dir’.

--srcdir=dirname

Warning: using this option requires gnu make, or another make that implements the VPATH feature. Use this option to make configurations in directories separate from the [No value for “GDBN”] source directories. Among other things, you can use this to build (or maintain) several configurations simultaneously, in separate directories. configure writes configuration specific files in the current directory, but arranges for them to use the source in the directory dirname. configure creates directories under the working directory in parallel to the source directories below dirname.

--norecursion

Configure only the directory level where configure is executed; do not propagate configuration to subdirectories.

--target=target

Configure [No value for “GDBN”] for cross-debugging programs running on the specified target. Without this option, [No value for “GDBN”] is configured to debug programs that run on the same machine (host) as [No value for “GDBN”] itself. There is no convenient way to generate a list of all available targets.

host ... Configure [No value for “GDBN”] to run on the specified host.

There is no convenient way to generate a list of all available hosts.

There are many other options available as well, but they are generally needed for special purposes only.

Index 211

Index

#

# (a comment). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 # in Modula-2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

$

$ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 $$ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 $_and info breakpoints . . . . . . . . . . . . . . . . . . . . 36 $_and info line . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 $_, $__, and value history . . . . . . . . . . . . . . . . . . . . 70 $_, convenience variable . . . . . . . . . . . . . . . . . . . . . . 78 $__, convenience variable . . . . . . . . . . . . . . . . . . . . . 78 $_exitcode, convenience variable . . . . . . . . . . . . . 79 $bpnum, convenience variable. . . . . . . . . . . . . . . . . . 34 $cdir, convenience variable . . . . . . . . . . . . . . . . . . . 61 $cwdr, convenience variable . . . . . . . . . . . . . . . . . . . 61 $tpnum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 $trace_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 $trace_frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 $trace_func . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 $trace_line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 $tracepoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

-

--annotate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --async. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --batch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --baud . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --cd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --core . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --epoch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --fullname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --interpreter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --mapped. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --noasync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --nowindows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --nx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --quiet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --readnow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --se . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --silent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --tty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --tui . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --write. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -nw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

.

., Modula-2 scope operator . . . . . . . . . . . . . . . . . . 107 ‘.esgdbinit’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 ‘.gdbinit’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 ‘.os68gdbinit’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 ‘.vxgdbinit’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179

/

/proc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

:

::, context for variables/functions . . . . . . . . . . . . 66 ::, in Modula-2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

@

@, referencing memory as an array. . . . . . . . . . . . . 67

‘

“No symbol "foo" in current context” . . . . . . . . . 67

[No value for “GDBN”]

[No value for “GDBN”] bugs, reporting. . . . . . . 201 [No value for “GDBN”] reference card . . . . . . . . 205

                              {

{type} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

212 Debugging with [No value for “GDBN”]

A

a.out and C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 abbreviation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 acknowledgment, for [No value for “GDBN”]

remote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 actions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 active targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

add-shared-symbol-file . . . . . . . . . . . . . . . . . . . 125

add-symbol-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 address of a symbol . . . . . . . . . . . . . . . . . . . . . . . . . 115 Alpha stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 AMD 29K register stack. . . . . . . . . . . . . . . . . . . . . 169 AMD EB29K. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 AMD29K via UDI . . . . . . . . . . . . . . . . . . . . . . . . . . 156 annotations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 annotations for breakpoints. . . . . . . . . . . . . . . . . . 196 annotations for display . . . . . . . . . . . . . . . . . . . . . . 195 annotations for errors, warnings and interrupts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 annotations for frames . . . . . . . . . . . . . . . . . . . . . . 193 annotations for invalidation messages. . . . . . . . . 197 annotations for prompts . . . . . . . . . . . . . . . . . . . . . 195 annotations for running programs . . . . . . . . . . . . 197 annotations for source display . . . . . . . . . . . . . . . 198 annotations for values . . . . . . . . . . . . . . . . . . . . . . . 192 apropos. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 arg-begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 arg-end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 arg-name-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 arg-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 arguments (to your program) . . . . . . . . . . . . . . . . . 24 array-section-end . . . . . . . . . . . . . . . . . . . . . . . . . 193 artificial array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 assembly instructions . . . . . . . . . . . . . . . . . . . . . 62, 63 assignment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 AT&T disassembly flavor . . . . . . . . . . . . . . . . . . . . . 63 attach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 automatic display . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 automatic thread selection. . . . . . . . . . . . . . . . . . . . 30 awatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

B

b (break) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 backtrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 backtraces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 break ... thread threadno . . . . . . . . . . . . . . . . . . . 50 break in overloaded functions . . . . . . . . . . . . . . . . 102 breakpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 breakpoint commands . . . . . . . . . . . . . . . . . . . . . . . . 43 breakpoint conditions . . . . . . . . . . . . . . . . . . . . . . . . 42 breakpoint numbers . . . . . . . . . . . . . . . . . . . . . . . . . . 33 breakpoint on events . . . . . . . . . . . . . . . . . . . . . . . . . 33 breakpoint on memory address. . . . . . . . . . . . . . . . 33 breakpoint on variable modification . . . . . . . . . . . 33 breakpoint ranges. . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 breakpoint subroutine, remote . . . . . . . . . . . . . . 133 breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 breakpoints and threads . . . . . . . . . . . . . . . . . . . . . . 50 breakpoints-headers. . . . . . . . . . . . . . . . . . . . . . . 196 breakpoints-invalid. . . . . . . . . . . . . . . . . . . . . . . 197 breakpoints-table . . . . . . . . . . . . . . . . . . . . . . . . . 196 breakpoints-table-end . . . . . . . . . . . . . . . . . . . . 197 bt (backtrace) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 bug criteria . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 bug reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 bugs in [No value for “GDBN”] . . . . . . . . . . . . . . 201

C

c (continue) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 C and C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 C and C++ checks. . . . . . . . . . . . . . . . . . . . . . . . . . . 102 C and C++ constants . . . . . . . . . . . . . . . . . . . . . . . . 100 C and C++ defaults . . . . . . . . . . . . . . . . . . . . . . . . . 101 C and C++ operators . . . . . . . . . . . . . . . . . . . . . . . . . 98 C-L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 C-x 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 C-x 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 C-x a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 C-x A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 C-x C-a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 C++ and object formats . . . . . . . . . . . . . . . . . . . . . 100 C++ exception handling . . . . . . . . . . . . . . . . . . . . . 102 C++ scope resolution . . . . . . . . . . . . . . . . . . . . . . . . . 66 C++ support, not in coff . . . . . . . . . . . . . . . . . . . 100 C++ symbol decoding style . . . . . . . . . . . . . . . . . . . 75 C++ symbol display . . . . . . . . . . . . . . . . . . . . . . . . . 102 call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 call overloaded functions . . . . . . . . . . . . . . . . . . . . 101 call stack. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 calling functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 calling make. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 casts, to view memory . . . . . . . . . . . . . . . . . . . . . . . 66 catch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catch catch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catch exceptions, list active handlers . . . . . . . . . . 56 catch exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catch fork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catch load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catch throw. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catch unload. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catch vfork. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 catchpoints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 catchpoints, setting . . . . . . . . . . . . . . . . . . . . . . . . . . 39 cd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 cdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 checks, range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 checks, type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 checksum, for [No value for “GDBN”] remote. . 136 Chill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 choosing target byte order . . . . . . . . . . . . . . . . . . . 131 clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 clearing breakpoints, watchpoints, catchpoints. . 40 coff versus C++. . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 collect (tracepoints) . . . . . . . . . . . . . . . . . . . . . . . . 85 collected data discarded . . . . . . . . . . . . . . . . . . . . . . 87 colon, doubled as scope operator . . . . . . . . . . . . . 107 colon-colon, context for variables/functions …. 66 command files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 command hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 command line editing . . . . . . . . . . . . . . . . . . . . . . . 171 commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43, 196 commands for C++ . . . . . . . . . . . . . . . . . . . . . . . . . . 102 commands to STDBUG (ST2000) . . . . . . . . . . . . 168 comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 compilation directory . . . . . . . . . . . . . . . . . . . . . . . . 61 compiling, on Sparclet . . . . . . . . . . . . . . . . . . . . . . 166 complete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 completion of quoted strings . . . . . . . . . . . . . . . . . . 18 condition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 conditional breakpoints . . . . . . . . . . . . . . . . . . . . . . 42 configuring [No value for “GDBN”]. . . . . . . . . . . 207 confirmation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 connect (to STDBUG) . . . . . . . . . . . . . . . . . . . . . . 168 continue. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 continuing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 continuing threads . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 control C, and remote debugging. . . . . . . . . . . . . 134 controlling terminal . . . . . . . . . . . . . . . . . . . . . . . . . . 27 convenience variables. . . . . . . . . . . . . . . . . . . . . . . . . 78 convenience variables for tracepoints . . . . . . . . . . 90 core . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 core dump file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 core-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 crash of debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 current directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 current stack frame . . . . . . . . . . . . . . . . . . . . . . . . . . 54 current thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 cwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

D

d (delete) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 debugger crash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 debugging optimized code . . . . . . . . . . . . . . . . . . . . 23 debugging stub, example . . . . . . . . . . . . . . . . . . . . 136 debugging target . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 define. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 delete breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 delete display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 delete mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 delete tracepoint . . . . . . . . . . . . . . . . . . . . . . . . . . 84 deleting breakpoints, watchpoints, catchpoints

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 demangling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 descriptor tables display. . . . . . . . . . . . . . . . . . . . . 152 detach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 directories for source files . . . . . . . . . . . . . . . . . . . . . 61 directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 directory, compilation . . . . . . . . . . . . . . . . . . . . . . . . 61 directory, current . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 dis (disable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 disable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 disable breakpoints . . . . . . . . . . . . . . . . . . . . . . . . 41 disable display . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 disable mem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 disable tracepoint . . . . . . . . . . . . . . . . . . . . . . . . . 84 disassemble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 display of expressions . . . . . . . . . . . . . . . . . . . . . . . . 70 display-begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 display-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 display-expression. . . . . . . . . . . . . . . . . . . . . . . . 195

display-expression-end . . . . . . . . . . . . . . . . . . . 195

display-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 display-number-end. . . . . . . . . . . . . . . . . . . . . . . . 195 display-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 djgpp debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 do (down) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 down-silently . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 download to H8/300 or H8/500 . . . . . . . . . . . . . . 159 download to Hitachi SH . . . . . . . . . . . . . . . . . . . . . 159 download to Nindy-960. . . . . . . . . . . . . . . . . . . . . . 162 download to Sparclet. . . . . . . . . . . . . . . . . . . . . . . . 167 download to VxWorks. . . . . . . . . . . . . . . . . . . . . . . 155 dump all data collected at tracepoint . . . . . . . . . . 89 dynamic linking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

E

‘eb.log’, a log file for EB29K. . . . . . . . . . . . . . . . 158 EB29K board . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 EBMON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 ecoff and C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 elf/dwarf and C++ . . . . . . . . . . . . . . . . . . . . . . . . 100 elf/stabs and C++. . . . . . . . . . . . . . . . . . . . . . . . . . 100 else . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 elt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 elt-rep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 elt-rep-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 enable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 enable breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . 41 enable display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 enable mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 enable tracepoint . . . . . . . . . . . . . . . . . . . . . . . . . . 84 end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 entering numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 environment (of your program) . . . . . . . . . . . . . . . 25 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 error on valid input . . . . . . . . . . . . . . . . . . . . . . . . . 201 error-begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 event handling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 examining data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 examining memory. . . . . . . . . . . . . . . . . . . . . . . . . . . 69 exception handlers . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 exception handlers, how to list . . . . . . . . . . . . . . . . 56 exceptionHandler . . . . . . . . . . . . . . . . . . . . . . . . . . 134 exec-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 executable file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 exited. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 exiting [No value for “GDBN”] . . . . . . . . . . . . . . . 15 expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 expressions in C or C++ . . . . . . . . . . . . . . . . . . . . . . 98 expressions in C++ . . . . . . . . . . . . . . . . . . . . . . . . . . 100 expressions in Modula-2 . . . . . . . . . . . . . . . . . . . . . 103

F

f (frame) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 fatal signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 fatal signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 fg (resume foreground execution) . . . . . . . . . . . . . 46 field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 field-begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 field-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 field-name-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 field-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 find trace snapshot. . . . . . . . . . . . . . . . . . . . . . . . . . . 87 finish . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 flinching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 floating point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 floating point registers . . . . . . . . . . . . . . . . . . . . . . . 79 floating point, MIPS remote . . . . . . . . . . . . . . . . . 165 flush_i_cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 focus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 focus of debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 foo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 fork, debugging programs which call. . . . . . . . . . . 31 format options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 formatted output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Fortran . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 forward-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 frame number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 frame pointer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 frame, command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 frame, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 frame, selecting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 frame-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 frame-address-end . . . . . . . . . . . . . . . . . . . . . . . . . 194 frame-args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 frame-begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 frame-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 frame-function-name. . . . . . . . . . . . . . . . . . . . . . . 194 frame-source-begin. . . . . . . . . . . . . . . . . . . . . . . . 194 frame-source-end . . . . . . . . . . . . . . . . . . . . . . . . . . 194 frame-source-file . . . . . . . . . . . . . . . . . . . . . . . . . 194 frame-source-file-end . . . . . . . . . . . . . . . . . . . . 194 frame-source-line . . . . . . . . . . . . . . . . . . . . . . . . . 194 frame-where . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 frameless execution . . . . . . . . . . . . . . . . . . . . . . . . . . 53 frames-invalid . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 free memory information (MS-DOS). . . . . . . . . . 152 Fujitsu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 function-call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 functions without line info, and stepping. . . . . . . 47

G

g++, gnu C++ compiler . . . . . . . . . . . . . . . . . . . . . . . 98 garbled pointers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 ‘gdb.ini’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 GDBHISTFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 gdbserve.nlm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 gdbserver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 GDT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 getDebugChar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 gnu C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 gnu Emacs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

H

h (help) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 H8/300 or H8/500 download. . . . . . . . . . . . . . . . . 159 handle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 handle_exception . . . . . . . . . . . . . . . . . . . . . . . . . . 133 handling signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 hardware watchpoints . . . . . . . . . . . . . . . . . . . . . . . . 37 hbreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 help target. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 help user-defined . . . . . . . . . . . . . . . . . . . . . . . . . 177 heuristic-fence-post (Alpha, MIPS) . . . . . . . 170 history expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 history file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 history number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 history save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 history size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 history substitution . . . . . . . . . . . . . . . . . . . . . . . . . 171 Hitachi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 Hitachi SH download. . . . . . . . . . . . . . . . . . . . . . . . 159 hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 hook- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 hookpost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 hookpost- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 hooks, for commands. . . . . . . . . . . . . . . . . . . . . . . . 178 hooks, post-command . . . . . . . . . . . . . . . . . . . . . . . 178 hooks, pre-command . . . . . . . . . . . . . . . . . . . . . . . . 178

I

i (info) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 i/o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ‘i386-stub.c’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 IDT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 ignore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 ignore count (of breakpoint) . . . . . . . . . . . . . . . . . . 43 INCLUDE_RDB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 info address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 info all-registers . . . . . . . . . . . . . . . . . . . . . . . . . 79 info args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 info breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 info catch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 info display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 info dos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 info extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 info f (info frame) . . . . . . . . . . . . . . . . . . . . . . . . . 56 info files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 info float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 info frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 info frame, show the source language . . . . . . . . . 95 info functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 info line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 info locals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 info mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 info proc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 info proc id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 info proc mappings . . . . . . . . . . . . . . . . . . . . . . . . 151 info proc status. . . . . . . . . . . . . . . . . . . . . . . . . . . 151 info proc times. . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 info program. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 info registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 info s (info stack) . . . . . . . . . . . . . . . . . . . . . . . . . 54 info scope . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 info set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 info share . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 info sharedlibrary . . . . . . . . . . . . . . . . . . . . . . . . 126 info signals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 info source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 info source, show the source language . . . . . . . . 95 info sources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 info stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 info symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 info target. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 info terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 info threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29, 30 info tracepoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 info types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 info variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 info watchpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 information about tracepoints . . . . . . . . . . . . . . . . 86 inheritance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 init file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 init file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 initial frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 innermost frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 inspect. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 instructions, assembly . . . . . . . . . . . . . . . . . . . . 62, 63 Intel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 Intel disassembly flavor. . . . . . . . . . . . . . . . . . . . . . . 63 internal [No value for “GDBN”] breakpoints … 36 interrupt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 interrupting remote programs. . . . . . . . . . . . . . . . 136 interrupting remote targets . . . . . . . . . . . . . . . . . . 134 invalid input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

J

jump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

K

kernel object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 kernel object display . . . . . . . . . . . . . . . . . . . . . . . . 150 kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 KOD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150

L

l (list) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 last tracepoint number . . . . . . . . . . . . . . . . . . . . . . . 84 latest breakpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 layout asm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 layout next. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 layout prev. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 layout regs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 layout split. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 layout src . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 LDT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 leaving [No value for “GDBN”] . . . . . . . . . . . . . . . 15 Left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 linespec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 listing machine instructions . . . . . . . . . . . . . . . 62, 63 load filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 local variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 locate address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 log file for EB29K . . . . . . . . . . . . . . . . . . . . . . . . . . 158

M

m680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ‘m68k-stub.c’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 machine instructions. . . . . . . . . . . . . . . . . . . . . . 62, 63 maint info breakpoints . . . . . . . . . . . . . . . . . . . . . 37 maint print psymbols . . . . . . . . . . . . . . . . . . . . . . 118 maint print symbols . . . . . . . . . . . . . . . . . . . . . . . 118 make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 mapped. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 member functions. . . . . . . . . . . . . . . . . . . . . . . . . . . 100 memory models, H8/500 . . . . . . . . . . . . . . . . . . . . 161 memory region attributes. . . . . . . . . . . . . . . . . . . . . 80 memory tracing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 memory, viewing as typed object. . . . . . . . . . . . . . 66 memory-mapped symbol file . . . . . . . . . . . . . . . . . 124 memset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 MIPS boards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 MIPS remote floating point. . . . . . . . . . . . . . . . . . 165 MIPS remotedebug protocol . . . . . . . . . . . . . . . . . 165 MIPS stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 Modula-2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Modula-2 built-ins . . . . . . . . . . . . . . . . . . . . . . . . . . 105 Modula-2 checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 Modula-2 constants . . . . . . . . . . . . . . . . . . . . . . . . . 106 Modula-2 defaults . . . . . . . . . . . . . . . . . . . . . . . . . . 107 Modula-2 operators . . . . . . . . . . . . . . . . . . . . . . . . . 104 Modula-2, [No value for “GDBN”] support. … 103 Modula-2, deviations from. . . . . . . . . . . . . . . . . . . 107 Motorola 680x0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 MS-DOS system info . . . . . . . . . . . . . . . . . . . . . . . . 152 MS-DOS-specific commands . . . . . . . . . . . . . . . . . 152 multiple processes . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 multiple targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 multiple threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

N

n (next) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 names of symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 namespace in C++ . . . . . . . . . . . . . . . . . . . . . . . . . . 100 native djgpp debugging . . . . . . . . . . . . . . . . . . . . . 152 negative breakpoint numbers . . . . . . . . . . . . . . . . . 36 New systag message . . . . . . . . . . . . . . . . . . . . . . . . . . 29 New systag message, on HP-UX . . . . . . . . . . . . . . . 29 next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 nexti . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ni (nexti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Nindy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 number representation . . . . . . . . . . . . . . . . . . . . . . 173 numbers for breakpoints. . . . . . . . . . . . . . . . . . . . . . 33

O

object formats and C++ . . . . . . . . . . . . . . . . . . . . . 100 online documentation . . . . . . . . . . . . . . . . . . . . . . . . 19 optimized code, debugging. . . . . . . . . . . . . . . . . . . . 23 outermost frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 output formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 overload-choice . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 overloaded functions, calling . . . . . . . . . . . . . . . . . 101 overloaded functions, overload resolution . . . . . 103 overloading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 overloading in C++ . . . . . . . . . . . . . . . . . . . . . . . . . . 102

P

packets, reporting on stdout . . . . . . . . . . . . . . . . . 175 page tables display (MS-DOS) . . . . . . . . . . . . . . . 152 partial symbol dump . . . . . . . . . . . . . . . . . . . . . . . . 118 Pascal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 passcount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 patching binaries . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 pauses in output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 PgDn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 PgUp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 physical address from linear address . . . . . . . . . . 153 pipes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 pointer, finding referent . . . . . . . . . . . . . . . . . . . . . . 73 post-commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 post-overload-choice. . . . . . . . . . . . . . . . . . . . . . 196 post-prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

post-prompt-for-continue . . . . . . . . . . . . . . . . . 196

post-query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 pre-commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 pre-overload-choice. . . . . . . . . . . . . . . . . . . . . . . 196 pre-prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

pre-prompt-for-continue . . . . . . . . . . . . . . . . . . 196

pre-query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 print settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 printf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 printing data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 process image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 processes, multiple . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 prompt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 prompt-for-continue. . . . . . . . . . . . . . . . . . . . . . . 196 protocol, [No value for “GDBN”] remote serial

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 ptype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 putDebugChar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 pwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

Q

q (quit) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 quit [expression] . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 quotes in commands . . . . . . . . . . . . . . . . . . . . . . . . . 18 quoting names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

R

r (run). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 raise exceptions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 range checking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ranges of breakpoints . . . . . . . . . . . . . . . . . . . . . . . . 33 rbreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 reading symbols immediately . . . . . . . . . . . . . . . . 124 readline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 readnow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 recent tracepoint number . . . . . . . . . . . . . . . . . . . . . 84 record. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 reference card . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 reference declarations . . . . . . . . . . . . . . . . . . . . . . . 101 refresh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 register stack, AMD29K . . . . . . . . . . . . . . . . . . . . . 169 registers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 regular expression. . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 reloading symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . 117 remote connection without stubs . . . . . . . . . . . . . 148 remote debugging. . . . . . . . . . . . . . . . . . . . . . . . . . . 132 remote programs, interrupting . . . . . . . . . . . . . . . 136 remote serial debugging summary . . . . . . . . . . . . 135 remote serial debugging, overview . . . . . . . . . . . . 132 remote serial protocol . . . . . . . . . . . . . . . . . . . . . . . 136 remote serial stub. . . . . . . . . . . . . . . . . . . . . . . . . . . 133 remote serial stub list . . . . . . . . . . . . . . . . . . . . . . . 132 remote serial stub, initialization. . . . . . . . . . . . . . 133 remote serial stub, main routine . . . . . . . . . . . . . 133 remote stub, example . . . . . . . . . . . . . . . . . . . . . . . 136 remote stub, support routines. . . . . . . . . . . . . . . . 134 remotedebug, MIPS protocol . . . . . . . . . . . . . . . . 165 remotetimeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 remove actions from a tracepoint. . . . . . . . . . . . . . 85 repeating commands . . . . . . . . . . . . . . . . . . . . . . . . . 17 reporting bugs in [No value for “GDBN”] . . . . . 201 reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 response time, MIPS debugging . . . . . . . . . . . . . . 170 resuming execution . . . . . . . . . . . . . . . . . . . . . . . . . . 46 RET (repeat last command) . . . . . . . . . . . . . . . . . . . 17 retransmit-timeout, MIPS protocol . . . . . . . . . 165 return. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 returning from a function . . . . . . . . . . . . . . . . . . . 121 reverse-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 run . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 running . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 running 29K programs . . . . . . . . . . . . . . . . . . . . . . 156 running and debugging Sparclet programs …. 167 running VxWorks tasks . . . . . . . . . . . . . . . . . . . . . 155 running, on Sparclet . . . . . . . . . . . . . . . . . . . . . . . . 166 rwatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

S

s (step) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 save tracepoints for future sessions . . . . . . . . . . . . 90 save-tracepoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 saving symbol table . . . . . . . . . . . . . . . . . . . . . . . . . 124 scope . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 searching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 section. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 segment descriptor tables. . . . . . . . . . . . . . . . . . . . 152 select trace snapshot . . . . . . . . . . . . . . . . . . . . . . . . . 87 select-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 selected frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 selecting frame silently . . . . . . . . . . . . . . . . . . . . . . . 54 sequence-id, for [No value for “GDBN”] remote

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 serial connections, debugging . . . . . . . . . . . . . . . . 175 serial device, Hitachi micros . . . . . . . . . . . . . . . . . 159 serial line speed, Hitachi micros . . . . . . . . . . . . . . 159 serial line, target remote . . . . . . . . . . . . . . . . . . . 135 serial protocol, [No value for “GDBN”] remote

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 server prefix for annotations . . . . . . . . . . . . . . . . . 191 set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 set args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 set auto-solib-add . . . . . . . . . . . . . . . . . . . . . . . . 126 set check range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 set check type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 set check, range . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 set check, type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 set complaints . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 set confirm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 set debug arch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 set debug event. . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 set debug expression . . . . . . . . . . . . . . . . . . . . . . 175 set debug overload . . . . . . . . . . . . . . . . . . . . . . . . 175 set debug remote. . . . . . . . . . . . . . . . . . . . . . . . . . . 175 set debug serial. . . . . . . . . . . . . . . . . . . . . . . . . . . 176 set debug target. . . . . . . . . . . . . . . . . . . . . . . . . . . 176 set debug varobj. . . . . . . . . . . . . . . . . . . . . . . . . . . 176 set demangle-style . . . . . . . . . . . . . . . . . . . . . . . . . 75 set disassembly-flavor . . . . . . . . . . . . . . . . . . . . . 63 set editing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 set endian auto. . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 set endian big . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 set endian little. . . . . . . . . . . . . . . . . . . . . . . . . . 131 set environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 set extension-language . . . . . . . . . . . . . . . . . . . . . 95 set follow-fork-mode . . . . . . . . . . . . . . . . . . . . . . . 31 set gnutarget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 set height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 set history expansion . . . . . . . . . . . . . . . . . . . . . 172 set history filename . . . . . . . . . . . . . . . . . . . . . . 171 set history save. . . . . . . . . . . . . . . . . . . . . . . . . . . 172 set history size. . . . . . . . . . . . . . . . . . . . . . . . . . . 172 set input-radix . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 set language. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 set listsize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 set machine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 set memory mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 set mipsfpu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

set opaque-type-resolution . . . . . . . . . . . . . . . 117

set output-radix . . . . . . . . . . . . . . . . . . . . . . . . . . 174

set overload-resolution. . . . . . . . . . . . . . . . . . . 103

set print address. . . . . . . . . . . . . . . . . . . . . . . . . . . 72 set print array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 set print asm-demangle . . . . . . . . . . . . . . . . . . . . . 75 set print demangle . . . . . . . . . . . . . . . . . . . . . . . . . 75 set print elements . . . . . . . . . . . . . . . . . . . . . . . . . 73

set print max-symbolic-offset. . . . . . . . . . . . . . 73

set print null-stop . . . . . . . . . . . . . . . . . . . . . . . . 74 set print object. . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 set print pretty. . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

set print sevenbit-strings . . . . . . . . . . . . . . . . . 74
set print static-members . . . . . . . . . . . . . . . . . . . 76
set print symbol-filename . . . . . . . . . . . . . . . . . . 73

set print union . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 set print vtbl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 set processor args . . . . . . . . . . . . . . . . . . . . . . . . . 164 set prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 set remotedebug, MIPS protocol . . . . . . . . . . . . 165 set retransmit-timeout . . . . . . . . . . . . . . . . . . . . 165

set rstack_high_address. . . . . . . . . . . . . . . . . . . 169

set step-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 set symbol-reloading . . . . . . . . . . . . . . . . . . . . . . 117 set timeout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 set tracepoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

set tui active-border-mode . . . . . . . . . . . . . . . . 185

set tui border-kind . . . . . . . . . . . . . . . . . . . . . . . 185 set tui border-mode . . . . . . . . . . . . . . . . . . . . . . . 185 set variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 set verbose. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 set width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 set write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 set_debug_traps . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 setting variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 setting watchpoints . . . . . . . . . . . . . . . . . . . . . . . . . . 37 SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ‘sh-stub.c’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 share . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 shared libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 sharedlibrary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 shell escape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 show . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 show args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 show auto-solib-add . . . . . . . . . . . . . . . . . . . . . . . 126 show check range. . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 show check type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 show complaints . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 show confirm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 show convenience . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 show copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 show debug arch. . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 show debug event. . . . . . . . . . . . . . . . . . . . . . . . . . . 175 show debug expression . . . . . . . . . . . . . . . . . . . . . 175 show debug overload . . . . . . . . . . . . . . . . . . . . . . . 175 show debug remote. . . . . . . . . . . . . . . . . . . . . . . . . . 176 show debug serial. . . . . . . . . . . . . . . . . . . . . . . . . . 176 show debug target. . . . . . . . . . . . . . . . . . . . . . . . . . 176 show debug varobj. . . . . . . . . . . . . . . . . . . . . . . . . . 176 show demangle-style . . . . . . . . . . . . . . . . . . . . . . . . 76 show directories . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 show editing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 show environment . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 show gnutarget . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 show height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 show history. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 show input-radix . . . . . . . . . . . . . . . . . . . . . . . . . . 174 show language. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 show listsize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 show machine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 show mipsfpu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

show opaque-type-resolution . . . . . . . . . . . . . . 118

show output-radix . . . . . . . . . . . . . . . . . . . . . . . . . 174 show paths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 show print address . . . . . . . . . . . . . . . . . . . . . . . . . 72 show print array. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 show print asm-demangle . . . . . . . . . . . . . . . . . . . . 75 show print demangle . . . . . . . . . . . . . . . . . . . . . . . . 75 show print elements . . . . . . . . . . . . . . . . . . . . . . . . 74

show print max-symbolic-offset. . . . . . . . . . . . . 73

show print object. . . . . . . . . . . . . . . . . . . . . . . . . . . 76 show print pretty. . . . . . . . . . . . . . . . . . . . . . . . . . . 74

show print sevenbit-strings . . . . . . . . . . . . . . . . 74
show print static-members . . . . . . . . . . . . . . . . . . 76
show print symbol-filename . . . . . . . . . . . . . . . . . 73

show print union. . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 show print vtbl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 show processor . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 show prompt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 show remotedebug, MIPS protocol . . . . . . . . . . . 165

show retransmit-timeout. . . . . . . . . . . . . . . . . . . 165
show rstack_high_address. . . . . . . . . . . . . . . . . . 169

show symbol-reloading . . . . . . . . . . . . . . . . . . . . . 117 show timeout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 show user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 show values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 show verbose. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 show version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 show warranty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 show width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 show write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 shows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 si (stepi) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121, 198 signal-handler-caller . . . . . . . . . . . . . . . . . . . . 194 signal-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 signal-name-end . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 signal-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 signal-string-end . . . . . . . . . . . . . . . . . . . . . . . . . 198 signalled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 silent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 sim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 simulator, Z8000. . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 size of screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 software watchpoints . . . . . . . . . . . . . . . . . . . . . . . . . 37 source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180, 198 source path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Sparc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ‘sparc-stub.c’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ‘sparcl-stub.c’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 Sparclet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 SparcLite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 ST2000 auxiliary commands . . . . . . . . . . . . . . . . . 168 st2000 cmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 stack frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 stack on Alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 stack on MIPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 stack traces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 stacking targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 start a new trace experiment . . . . . . . . . . . . . . . . . 87 starting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 starting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 status of trace data collection . . . . . . . . . . . . . . . . . 87 STDBUG commands (ST2000) . . . . . . . . . . . . . . 168 step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 stepi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 stepping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 stepping into functions with no line info . . . . . . . 47 stop a running trace experiment . . . . . . . . . . . . . . 87 stop, a pseudo-command . . . . . . . . . . . . . . . . . . . . 178 stopped threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 stopping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 stub example, remote debugging . . . . . . . . . . . . . 136 stupid questions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 switching threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 switching threads automatically. . . . . . . . . . . . . . . 30 symbol decoding style, C++ . . . . . . . . . . . . . . . . . . . 75 symbol dump. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 symbol from address . . . . . . . . . . . . . . . . . . . . . . . . 115 symbol names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 symbol overloading . . . . . . . . . . . . . . . . . . . . . . . . . . 45 symbol table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 symbol-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 symbols, reading immediately. . . . . . . . . . . . . . . . 124 sysinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

T

target. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 target abug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target adapt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 target amd-eb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 target array. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 target bug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target byte order . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 target core. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 target cpu32bug . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target dbug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target ddb port . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 target dink32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 target e7000, with H8/300. . . . . . . . . . . . . . . . . . 159 target e7000, with Hitachi ICE . . . . . . . . . . . . . 161 target e7000, with Hitachi SH . . . . . . . . . . . . . . 166 target es1800 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target est . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target exec. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 target hms, and serial protocol . . . . . . . . . . . . . . 160 target hms, with H8/300 . . . . . . . . . . . . . . . . . . . . 159 target hms, with Hitachi SH . . . . . . . . . . . . . . . . 166 target lsi port . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 target m32r. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target mips port . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 target mon960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 target nindy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 target nrom. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 target op50n. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 target pmon port . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 target ppcbug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 target ppcbug1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 target r3900. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 target rdi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 target rdp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 target remote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 target rom68k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target rombug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 target sds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 target sh3, with H8/300 . . . . . . . . . . . . . . . . . . . . 159 target sh3, with SH . . . . . . . . . . . . . . . . . . . . . . . . 166 target sh3e, with H8/300. . . . . . . . . . . . . . . . . . . 159 target sh3e, with SH . . . . . . . . . . . . . . . . . . . . . . . 166 target sim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 target sim, with Z8000 . . . . . . . . . . . . . . . . . . . . . 169 target sparclite . . . . . . . . . . . . . . . . . . . . . . . . . . 168 target vxworks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 target w89k. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 tbreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 TCP port, target remote . . . . . . . . . . . . . . . . . . . 136 tdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 tfind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 thbreak. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 this, inside C++ member functions. . . . . . . . . . . 100 thread apply. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 thread breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 thread identifier (GDB) . . . . . . . . . . . . . . . . . . . . . . 29 thread identifier (system). . . . . . . . . . . . . . . . . . . . . 29 thread identifier (system), on HP-UX. . . . . . . . . . 29 thread number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 thread threadno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 threads and watchpoints . . . . . . . . . . . . . . . . . . . . . 39 threads of execution . . . . . . . . . . . . . . . . . . . . . . . . . 28 threads, automatic switching . . . . . . . . . . . . . . . . . 30 threads, continuing . . . . . . . . . . . . . . . . . . . . . . . . . . 51 threads, stopped . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 timeout, MIPS protocol. . . . . . . . . . . . . . . . . . . . . 165 trace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 trace experiment, status of . . . . . . . . . . . . . . . . . . . 87 tracebacks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 tracepoint actions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 tracepoint data, display . . . . . . . . . . . . . . . . . . . . . . 89 tracepoint deletion . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 tracepoint number . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 tracepoint pass count . . . . . . . . . . . . . . . . . . . . . . . . 84 tracepoint variables . . . . . . . . . . . . . . . . . . . . . . . . . . 90 tracepoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 tstart . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 tstatus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 tstop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 tty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 TUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 TUI commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 TUI configuration variables. . . . . . . . . . . . . . . . . . 185 TUI key bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 type casting memory . . . . . . . . . . . . . . . . . . . . . . . . . 66 type checking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 type conversions in C++ . . . . . . . . . . . . . . . . . . . . . 101

U

u (until) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 udi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 UDI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 undisplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 unknown address, locating. . . . . . . . . . . . . . . . . . . . 68 unset environment . . . . . . . . . . . . . . . . . . . . . . . . . . 26 until . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 up-silently . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 update. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 user-defined command . . . . . . . . . . . . . . . . . . . . . . 177

V

value history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 value-begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 value-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 value-history-begin. . . . . . . . . . . . . . . . . . . . . . . 192 value-history-end . . . . . . . . . . . . . . . . . . . . . . . . . 192 value-history-value. . . . . . . . . . . . . . . . . . . . . . . 192 variable name conflict . . . . . . . . . . . . . . . . . . . . . . . . 66 variable values, wrong . . . . . . . . . . . . . . . . . . . . . . . . 66 variables, setting. . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 version number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 VxWorks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 vxworks-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

W

watch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 watchpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 watchpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 watchpoints and threads . . . . . . . . . . . . . . . . . . . . . 39 whatis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 where . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 while-stepping (tracepoints) . . . . . . . . . . . . . . . . 86 wild pointer, interpreting . . . . . . . . . . . . . . . . . . . . . 73 winheight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 word completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 working directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 working directory (of your program). . . . . . . . . . . 26 working language . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 writing into corefiles . . . . . . . . . . . . . . . . . . . . . . . . 122 writing into executables . . . . . . . . . . . . . . . . . . . . . 122 wrong values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

X

x (examine memory) . . . . . . . . . . . . . . . . . . . . . . . . . 69 x(examine), and info line . . . . . . . . . . . . . . . . . . . 62 xcoff and C++. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

Z

Z8000. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Zilog Z8000 simulator . . . . . . . . . . . . . . . . . . . . . . . 169

The body of this manual is set in

cmr10 at 10.95pt, with headings in cmb10 at 10.95pt

and examples in cmtt10 at 10.95pt.

cmti10 at 10.95pt, cmb10 at 10.95pt, and

cmsl10 at 10.95pt are used for emphasis.